### v1

# Landauer Principle and Thermodynamics of Computation

Abstract
According to the Landauer principle, any logically irreversible process accompanies entropy production, which results in heat dissipation in the environment. Erasing of information, one of the primary logically irreversible processes, has a lower bound on heat dissipated into the environment, called the Landauer bound (LB). However, the practical erasure processes dissipate much more heat than the LB. Recently, there have been a few experimental investigations to reach this bound both in the classical and quantum domains. There has also been a spate of activities to enquire about this LB in finite time, with finite-size heat baths, non-Markovian and nonequilibrium environment in the quantum regime where the effects of fluctuations and correlation of the systems with the bath can no longer be ignored. This article provides a comprehensive review of the recent progress on the Landauer bound, which serves as a fundamental principle in the thermodynamics of computation. We also provide a perspective for future endeavors in these directions.
Furthermore, we review the recent exploration toward establishing energetic bounds of a computational process. We also review the thermodynamic aspects of error correction, which is an indispensable part of information processing and computations. In doing so, we briefly discuss the basics of these fields to provide a complete picture.
Notations:
$\bullet$ $\mathbb{S}\,\,\,\epsilon\,\,\{0,1\}$
$\bullet$ $\mathbb{S}^{\star}\,\rightarrow$ the set of all binary strings.
$\bullet$ $\delta_{ij}=\left\{\begin{array}[]{lc}1,&\ i=j,\\ 0,&\ \text{otherwise}.\end{array}\right.$
$\bullet$ $\rho\rightarrow$ density matrix
$\bullet$ $H\rightarrow$ Hamiltonian
$\bullet$ $S\rightarrow$ Entropy
$\bullet$ $U\rightarrow$ Unitary
[…]
$\bullet$ $k_{B}\rightarrow$ Boltzmann constant
$\bullet$ $\beta\rightarrow\frac{1}{k_{B}T}$
$\bullet$ Tr (or  tr) $\rightarrow$ trace
$\bullet$ $\ln\rightarrow$ loge
$\bullet$ $eV\rightarrow$ Electron volt
$\bullet$ J $\rightarrow$ Joule
Acronym:
$\bullet$ RNA $\rightarrow$ Ribonucleic Acid
$\bullet$ TTL $\rightarrow$ Transistor Transistor Logic
$\bullet$ MD $\rightarrow$ Maxwell’s Demon
$\bullet$ LB $\rightarrow$ Landauer Bound
$\bullet$ LL $\rightarrow$ Landauer Limit
$\bullet$ LP $\rightarrow$ Landauer Principle
$\bullet$ LE $\rightarrow$ Landauer Erasure
$\bullet$ OQS $\rightarrow$ Open Quantum System
$\bullet$ CGF $\rightarrow$ Cumulant Generating Functions
$\bullet$ TLS $\rightarrow$ Two-level System
$\bullet$ NM $\rightarrow$ Non Markovian
$\bullet$ CPTP $\rightarrow$ Completely Positive Trace Preserving
$\bullet$ ABEL $\rightarrow$ Anti-Brownian Electrokinetic
$\bullet$ MOKE $\rightarrow$ Magneto-Optical Kerr Effect
$\bullet$ NMR $\rightarrow$ Nuclear Magnetic Resonance
$\bullet$ BLC $\rightarrow$ Ballistic Computer
$\bullet$ BWC $\rightarrow$ Brownian Computer
$\bullet$ TM $\rightarrow$ Turing Machine
$\bullet$ FA $\rightarrow$ Finite Automata
$\bullet$ FSM $\rightarrow$ Finite State Machine
$\bullet$ KC $\rightarrow$ Kolmogorov complexity
$\bullet$ DFA $\rightarrow$ Deterministic Finite Automata
$\bullet$ NFA $\rightarrow$ Non-deterministic Finite Automata
$\bullet$ PFA $\rightarrow$ Probabilistic Finite Automata
$\bullet$ ID $\rightarrow$ Instantaneous Description
$\bullet$ UTM $\rightarrow$ Universal Turing Machine
$\bullet$ EC $\rightarrow$ Error Correction
$\bullet$ ECC $\rightarrow$ Error Correcting code
$\bullet$ QEC $\rightarrow$ Quantum Error Correction
$\bullet$ LHS $\rightarrow$ Left Hand Side
$\bullet$ RHS $\rightarrow$ Right Hand Side
$\bullet$ MOS $\rightarrow$ Metal Oxide Semiconductor
$\bullet$ CMOS $\rightarrow$ Complementary Metal Oxide Semiconductor
$\bullet$ RTO $\rightarrow$ Restore To One
$\bullet$ BHT $\rightarrow$ Brassard Hoyer Tapp
$\bullet$ RAM $\rightarrow$ Random Access Memory
$\bullet$ QPT $\rightarrow$ Quantum Phase Transition
I Introduction
John von Neumann, in his 1948 lectures, posed a fundamental question regarding whether a computer operating at temperature $T$ must necessarily dissipate heat. The intricate details of this concept were later formalized in his book Theory of Self-Reproducing Automata, which was completed in 1966 by Arthur W. Burks Von Neumann et al. (1966). Building upon this idea, Landauer famously asserted that “real-world computation involves thermodynamic costs” and emphasized the substantial implications of this principle Landauer (1961, 1991). Further support for this notion came from Brillouin’s thought experiment Brillouin (1962), which validated the connection between computation and thermodynamics, albeit with some error probability. The man-made computers and even all naturally occurring processes like biological computers have thermodynamic costs. It is quite fascinating to analyze the difference in the thermodynamic cost of the naturally occurring process and the artificial ones that are created. Translation of ribonucleic acid (RNA) into amino acids is one such natural biological process where one encounters energy costs for the execution of the process. These biological processes are thermodynamically more efficient than the artificial ones Kempes et al. (2017).
Among the myriad artificial processes engineered by humans, digital computation stands out as one of the most significant. Modern digital computers can be viewed as engines that irreversibly dissipate energy to execute mathematical and logical operations. Early scientific thought postulated that there must exist a fundamental thermodynamic bound on the efficiency of such computational engines, independent of the hardware architecture employed. However, contemporary understanding has revealed a more refined reality: while the fundamental thermodynamic limit for the energy cost of erasing a single bit is set by $k_{B}T\ln 2$ (where $k_{B}$ is the Boltzmann constant and $T$ is the operational temperature), modern computers dissipate energy per logic operation that exceeds this bound by many orders of magnitude. Consequently, despite being capable of executing a vast number of reliable computations, practical devices remain highly inefficient relative to the theoretical minimum. A principal cause of this inefficiency lies in the reliance on volatile memory elements—such as Transistor-Transistor Logic (TTL) flip-flops Chandrakasan et al. (1992); Horowitz (2014)—which inherently waste energy. The macroscopic nature of the existing computers is one of the basic reasons for the inefficiency in the context of energy. One of the spectacular thermodynamically reversible computation models is the ballistic computer, proposed by Fredkin and Toffoli Fredkin and Toffoli (1982). However, other models Bennett (1973); Keyes and Landauer (1970); Likharev (1982) have been developed, which are more physically realistic than that of Fredkin’s version.
These examples of artificial and natural systems hinge on the deep connection between computation and thermodynamics. The connection between thermodynamics and a logically irreversible process gets most prominent in the context of Maxwell’s demon (MD) Maxwell (1871); Maruyama et al. (2009). Landauer (later Bennett) argued that one must pay an entropic cost that gets dumped in the environment as heat while performing a logically irreversible process that erases or throws away information. This argument played the pivotal role in exorcising the MD and thereby saving the second law of thermodynamics. It also established that information is physical. There have been several explorations to model MD in physical systems and study the Landauer principle (LP). However, they mostly belong to equilibrium statistical physics. But when we approach to the miniaturized domain, the systems are, in general, highly non-equilibrium in their form. The major breakthroughs in non-equilibrium statistical physics Zwanzig (2001); Kubo et al. (2012); Jarzynski (1997a, b) allow us to analyze the thermodynamic behavior of systems that are arbitrarily far from their equilibrium, and even for systems that undergo arbitrary external driving. With the advent of non-equilibrium statistical physics, the researchers have shown a keen interest to analyze the thermodynamic cost for the erasure process in this domain which includes finite time, finite size, open quantum systems, and so on Goold et al. (2015); Khoudiri et al. (2025); Lorenzo et al. (2015); Proesmans et al. (2020a); Miller et al. (2020); Esposito and Van den Broeck (2011); Mandal and Jarzynski (2012); Helms and Limmer (2022); Ray and Crutchfield (2023); Kuang et al. (2022); Proesmans and Bechhoefer (2021); Barnett and Vaccaro (2013); Berta et al. (2018); Henao and Uzdin (2023); Barato and Seifert (2013); Mandal et al. (2013); Deffner and Jarzynski (2013); Barato and Seifert (2014); Strasberg et al. (2014).
In nature, copying information is a fundamental mechanism within natural systems Watson and Crick (1953); Hopfield (1974); Ouldridge et al. (2017); Horowitz and England (2017a). Yet, replication is inherently prone to error. The accuracy of a copy thus relies on its accurate reproduction. It can be quantified by counting the wrongly copied bits while execution. Though the error can be reduced at the macro level, at the molecular level, perfect copying is not achievable due to thermal fluctuations. It is the primary source of error at the molecular level. The replication process is limited by the thermal noise, so it must be interpreted in terms of thermodynamics as proposed by von Neumann Von Neumann (1956). This raises an important question: whether one can develop a connection between the errors that occur when copying remains.
Generally, while executing a copying process, it has to undergo various intermediate steps to control the accuracy and speed of the process. This is true for artificial as well as natural scenarios Johnson (1993). If we try to explain the errors from thermodynamic laws (second law), one must account for the fact that the copying process is repeated cyclically rather than occurring as a single-shot operation Bennett (1982). The works Sartori and Pigolotti (2015); Korepin and Terilla (2002) forge a profound link between thermodynamics and informational errors, demonstrating that errors arising in copying protocols are intrinsically connected to thermodynamic observables, which characterize the errors.Another aspect to connect thermodynamics and information theory is built on the basis of using entanglement in quantum systems Horodecki et al. (2001); Popescu and Rohrlich (1997); Rohrlich (2001).
In this article, we review the recent developments in Landauer erasing in various contexts, including the most recent experimental demonstrations, thermodynamic aspects, and cost analysis of computation and error correction.
Scope of the review
Aiming to highlight recent advancements and fundamental insights into Landauer’s principle (LP) and the thermodynamics of computation, this review addresses several key aspects spanning both theoretical and experimental fronts. We begin in Sec. II with a discussion of the generalized LP. Sec. III reviews significant recent developments concerning LP, including scenarios involving finite time, finite-size systems, and non-Markovian reservoirs. In Sec. IV, we explore the role of Landauer’s limit (LL) in computational processes, followed by Sec. V, which presents an overview of recent experimental efforts aimed at realizing and validating LL in practical systems.
Consequently, we move towards exploring the thermodynamic ramifications and implications for computation and error correction. Sec. VI delves into reversible computational models, laying the groundwork for a broader discussion in Sec. VII on the thermodynamics of computational paradigms, including finite state automata and Turing machines. In Sec. VIII, we critically examine the thermodynamic consistency of error-correcting codes. Prior to concluding and outlining potential future research directions in Sec. X, we briefly address a range of complementary topics under the umbrella of miscellaneous discussions in Sec. IX, such as viewing the computer as a heat engine, the thermodynamics of algorithms, and Landauer bound in switching protocols.
Areas not covered in this review
Though the review covers an extensive amount of works in the context of LP and thermodynamics of computation, there are some related topics and works, which are not discussed, that are worth mentioning. For exmaple, while most of the aspects of the LP have been reviewed here, some of the left-out aspects in this domain are like the role of LP in gravity Bormashenko (2019a); Haranas et al. (2021); Daffertshofer and Plastino (2007), relativity Herrera (2020), quantum field theory Xu et al. (2022), many-body phenomena Bonança (2023); Parrondo (2001), material Zivieri (2022).
In this review, we have considered mainly two fundamental computational models: Turing machine and Finite state machine, namely, for detailed thermodynamic analysis. Some aspects, like LB in analog computers and algorithmic thermodynamics, have not been discussed here. Interested readers can go through Diamantini et al. (2016) for the LB analysis in the analog system and Baez and Stay (2012) for algorithmic thermodynamics, which in turn allows one to apply the laws and techniques of thermodynamics for studying algorithmic information theory; in contrast, we have focused on the thermodynamic cost of algorithms in Sec. IX.
In what follows, we briefly deal with computational complexity when we estimate the energetic cost of computation. However, a somewhat related direction, the fundamental limitation on the computability of the physical process, is not discussed in this review. Interested readers can have a look at the following seminal articles Pour-El and Richards (1982); Moore (1990); Lloyd (2000, 2017). Similarly, the thermodynamics of controlled systems is not covered. Please go through Touchette and Lloyd (2004, 2000); Barato and Seifert (2017); Sagawa and Ueda (2008, 2012); Wilming et al. (2016); Large and Large (2021); Gingrich et al. (2016); Horowitz and England (2017b) for details.
Another important aspect that has not been covered here is the thermodynamic analysis of biological and biochemical processes. For the same, one can go through Ouldridge and Ten Wolde (2017); Ouldridge (2018); Brittain et al. (2019); Sartori et al. (2014); Hasegawa (2018); Mehta and Schwab (2012); Mehta et al. (2016); Lan et al. (2012); Ouldridge et al. (2017); Govern and Ten Wolde (2014); Barato and Seifert (2015). Similarly, we do not cover the modeling of computational machines based on biochemical and biological systems. Interested readers can go through  Prohaska et al. (2010); Bryant (2012); Benenson (2012); Chen et al. (2014); Dong (2012); Soloveichik et al. (2008); Mougkogiannis and Adamatzky (2025) to have a clear idea in this direction.
II Landauer’s Principle
The search for the development of more impactful computing circuits leads to the question: What would be the physical limitation of this process? Rolf Landauer, in the year 1961, in his seminal work Landauer (1961) proposed an important limit to the conjecture surfaced by von Neumann Von Neumann et al. (1966), which is coined as “Landauer bound” (LB). This is a physical principle that provides the lower bound on the consumption of a computation process, i.e., an irreversible process in the computer dissipates a minimum amount of heat per bit Bennett (2003), which is expressed as
$\Delta Q\geq k_{B}T\ln 2,$ (1)
where $k_{B}$ is the Boltzman constant and $T$ is the temperature at which the computation process operates. The bound at room temperature is approximately 0.018 $eV$ (2.9 x $10^{-21}$ J), whereas the modern world computers use a million times more energy per operation Lambson et al. (2011a); Moore (2012).
Landauer Principle (LP) though widely accepted has faced various challenges and objections Earman and Norton (1998, 1999); Shenker (1998); Maroney (2005); Norton (2005, 2011); Bennett (2003). The prime objections to LP were put forth by Earman and Norton Earman and Norton (1999), where they have argued that LP is dependent on the second law of thermodynamics, it can be considered to be either unnecessary or insufficient for “exorcism of Maxwell’s Demon”. The other objections over LP are generally of three kinds Bennett (2003): (a) The principle fails, as thermodynamic quantities such as heat and work are fundamentally unrelated to mathematical constructs like logical reversibility. As a result, drawing direct parallels between them lacks any meaningful basis (as it was thought of to be a memory erasure process at that time). (b) In all cases of data-processing operations, there is a dissipation of at least $k_{B}T\ln 2$ amount of energy irrespective of the condition whether it is logically reversible or not. (c) Another reason for objection is that, in principle, logically irreversible operations can be engineered in a thermodynamically reversible way.
Attempts have been made to demonstrate LP with precision. Piechocinska Piechocinska (2000) offered proof of LP with the help of statistical mechanics. The proof had a particular physical model assumption. The notable advancement was provided in the work Turgut (2009); Ladyman et al. (2007, 2008); Cao and Feito (2009); Leff and Rex (2002), where it provides a generalized version of the LP without a particular physical model assumption. The generalization of the LP surfaced in Vaccaro and Barnett (2011), where it has been shown that the information erasure will cause an increase in the entropy of the environment with no energy cost but the cost can be attributed to angular momentum of a spin-reservoir. In fact, angular momentum is exploited rather than energy from a spin-reservoir to erase the memory and it is consistent with the second law of thermodynamics.
In the low-temperature limit, quantum effects play an important role in the analysis of the thermodynamic aspect of the erasure principle. The validity of the erasure principle in this limit has been challenged Allahverdyan and Nieuwenhuizen (2000); Nieuwenhuizen and Allahverdyan (2002); Hörhammer and Büttner (2005); Hilt and Lutz (2009). It has been claimed that due to the presence of entanglement between the system and the environment LP is broken, which implies information is erased while heat is absorbed Allahverdyan and Nieuwenhuizen (2001); Hörhammer and Büttner (2008); Cápek and Sheehan (2005); Maruyama et al. (2009) which counters the LP. Lutz et al. Hilt et al. (2011), tackled this quantum conundrum and demonstrated that LP remains valid regardless of the specific nature of the interaction between the system and its environment. Their findings provided significant evidence supporting the applicability of LP in quantum settings.
To address the need for a theory with minimal assumptions in both regimes, approaches are considered that derive the LP without relying explicitly on the second law of thermodynamics Shizume (1995); Piechocinska (2000); Sagawa and Ueda (2009, 2011). Reeb’s work Reeb and Wolf (2014) is on this line of thought. The assumptions that are considered for the analysis are minimal based on the benchmark works. The assumptions are: (a) In the process, the system and the reservoir belong to a Hilbert space, (b) the reservoir is initially in a thermal state, (c) the system and the reservoir will be initially uncorrelated, and (d) the evolution process is unitary.
With this minimum condition, a sharpened equality version of LP has been derived for a system $\mathcal{S}$ and a reservoir $R$ as
$\beta\Delta Q=\Delta S+I(\mathcal{S}^{\prime}:R^{\prime})+S(\rho_{R}^{\prime}||\rho_{R})\geq\Delta S.$ (2)
Here $\prime$ denotes the final states, $\Delta S$ is the change of von Neumann entropy of the system, $I(\mathcal{S}^{\prime}:R^{\prime})$ describes the mutual information and quantifies the correlation of the system with the bath, and $\rho_{R}$ denotes the reservoir state. $\beta=1/k_{B}T$ denotes the inverse temperature of the bath. $S(\rho_{R}^{\prime}||\rho_{R})$ quantifies the free energy increase in the bath, where $S(\rho||\rho^{\prime})=\text{Tr}\,\rho\,\ln\,\rho-\text{Tr}\,\rho\,\ln\,\rho^{\prime}$ is the quantum relative entropy between $\rho,\,\rho^{\prime}$.
The setup mentioned above can be generalized by considering an initial correlation in the process. So the assumption gets modified as: (a) In the process, the system, reservoir, and memory are in a joint quantum state initially, (b) the reservoir is initially in a thermal state, (c) the heat exchange and the entropy of the system are expounded on the marginal states, and (d) the evolution is processed by an unital positive trace-preserving map. In this generalized setup, where the system, reservoir, and memory are initially correlated, the standard Landauer bound can be modified significantly. Initial quantum or classical correlations can effectively reduce the minimal heat dissipation required for information erasure. Since entropy and heat are evaluated from marginal states, part of the entropy change may be absorbed by shared correlations rather than dissipated as heat. When the evolution is governed by a unital, trace-preserving map, which does not reduce entropy on its own, the role of initial correlations becomes even more prominent.
In the present-day protocols where the LP plays a vital role, Reeb’s version of the erasure process is considered for the analysis. Even with the minimal assumption of Reeb, it is clear that one requires dissipative dynamics to explore the LP. However, it is a difficult task in pure Hamiltonian dynamics where the evolution is for the system only. It has been conveyed in Holtzman et al. (2021) that it is possible to implement an erasure bit with no thermodynamic cost using Hamiltonian dynamics if one has the information of the system with infinite accuracy. In this case, it corresponds to the energy of the system being known with infinite accuracy.
So far, we have discussed LB for a non-zero finite temperature $T$ of the environment. What if the temperature of the environment tends to zero? In this limit (i.e., $T\rightarrow 0$) the bound in Eq. (1) is trivial. It states that $Q\geq 0$ when $T=0$ as the bath is in the ground state. A non-trivial bound for this condition has been proposed in Timpanaro et al. (2020). There, the authors have assumed that initially the environment is in a thermal state, as in the original case of LEP. The state of the system or the type of system-environment interaction is considered to be general. The improved bound is always greater than Eq. (1) and coincides with it when $T$ is high. It is derived based on two principles, the positivity of mutual information and the maximum entropy principle Guiasu and Shenitzer (1985); Wu (2012); Pressé et al. (2013), namely, and can be expressed in terms of the equilibrium heat capacity of the bath. It is interesting to explore what the modifications will be on the bound when the system and environment are initially correlated for the limit $T\rightarrow 0$.
III Landauer Principle considering the open-system dynamics
The study of the dynamics of a quantum system that interacts with other system or systems (environment) is described as an open quantum dynamics (OQD) Breuer et al. (2002); Rivas and Huelga (2012); Rotter and Bird (2015); Viola et al. (1999); Breuer et al. (2016); De Vega and Alonso (2017); Mukhopadhyay et al. (2017); Bhattacharya et al. (2017). It plays a crucial role in almost every aspects of quantum technology as the system which we are interested in is affected by the surrounding noise Breuer et al. (2002).
Till now, we have discussed the LB considering the entropic bounds sans the dynamics during erasing. The dynamics of erasing bear significant importance in practical scenarios. For example, the LB is reached in infinite time. But the practical uses demand finite-time erasure. Also, during the erasing procedure, the environment can get far from its equilibrium (due to finite size), and the fluctuations due to it are crucial for LB. What do the NM features add to the LP? These questions are not only practically important, but can also bring forth fundamental issues. In the following, we review the progress on the finite-time, non-equilibrium, and non-Markovian erasing processes, respectively, mainly in the quantum domain.
III.1 Finite Time Landauer’s Principle
The heat dissipated (LL) in the environment to erase one bit of information is achieved under the assumption of a quasistatic transformation. In practical scenarios, the information erasure occurs in finite time. Therefore, finite-time analysis Andresen (2011) bears much importance both in the classical and quantum regimes. There has been an upsurge of research on finite-time erasure of bits both in the quantum and stochastic thermodynamics Seifert (2012); Van den Broeck and Esposito (2015). In this finite-time regime, the erasing process takes place under non-equilibrium conditions, and the fluctuations in the dissipated heat become significant. This bears important consequences when designing minuscule logical devices that must be able to combat destructive fluctuations that lie well above the LB.
Research on minimizing the average dissipation of a mesoscopic thermodynamic system during finite-time transformations has primarily focused on optimizing a limited (and often small) number of control parameters that influence the system’s potential landscape Schmiedl and Seifert (2007); Bonança and Deffner (2014); Sivak and Crooks (2012); Tafoya et al. (2019); Plata et al. (2020); Bryant and Machta (2020); Boyd et al. (2018); Riechers et al. (2020); Rolandi and Perarnau-Llobet (2023). A significant advancement in this domain was made by Aurell et al. Aurell et al. (2011, 2012), who, using stochastic thermodynamics, developed protocols with full control over the potential landscape to minimize entropy production in both slow and fast limits, while constraining the final state to a fixed microscopic probability distribution. Building on this approach, the authors in their work Proesmans et al. (2020a) propose a framework that also provides full control over the potential landscape but removes the constraint that the final state must be a best-fitted state. Extensions Proesmans et al. (2020b) and alternative approaches Zhen et al. (2021); Dago et al. (2021); Dago and Bellon (2022) in this direction, without accounting for quantum effects, have been explored to establish an optimal bound on the cost of erasure.
Reeb and Wolf’s seminal work Reeb and Wolf (2014) provided a rigorous generalization of LP, demonstrating that quantum coherence fundamentally alters the thermodynamic cost of erasure. Their analysis showed that when the erased state exhibits coherence in the energy eigenbasis, the dissipation cost necessarily exceeds the classical LB. This correction arises because coherence prevents full thermalization through classical energy exchange alone, requiring additional dissipation mechanisms.
Expanding on this, Miller et al. Miller et al. (2020) investigated finite-time erasure under Markovian dynamics generated by the adiabatic Lindblad equation: $\dot{\rho}_{t}=\mathcal{L}\rho_{t}$. The generator $\mathcal{L}$ satisfies quantum details balance condition with respect to the instantaneous control Hamiltonian $H_{t}$ that guarantees an instantaneous fixed point of the dynamics, $\pi_{t}=e^{-\beta H_{t}}/\mbox{Tr}(e^{-\beta H_{t}})$ such that $\mathcal{L}(\pi_{t})=0$. Additionally, the driving of the control Hamiltonian is performed slowly in the time interval $t\in[0,\tau]$, relative to the relaxation timescale of the dynamics which implies that the system remains close to the equilibrium state at all times, i.e., $\rho_{t}=\pi_{t}+\tau^{-1}\delta\rho_{t}$ where corrections of higher orders $O(\tau^{-n})$ can be neglected. In this quasistatic limit, the dissipation approaches the LE bound.
The LP is performed by taking the initial Hamiltonian $H_{0}\simeq 0$ and then slowly increasing the energy gap of the Hamiltonian until it reaches far beyond $k_{B}T$. This is equivalent to ensuring the boundary conditions on the system’s state: $\rho_{0}=\pi_{0}\simeq I/d$ where $d$ denotes the dimensionality of the Hilbert space and $\rho_{\tau}=\pi_{\tau}\simeq|0\rangle\langle 0|$. To obtain the full statistics of the dissipated heat, the authors have used the cumulant generating functions (CGF) and quantified the excess stochastic heat in addition to the LB as follows
$Q=k_{B}T\ln d+Q_{d}+Q_{c}.$ (3)
In (3), the quantities are averaged over many trajectories, which represent each run of the experiment. Here, $Q_{d}$ is the classical (diagonal) contribution and $Q_{c}$ is the coherent contribution in the dissipated heat excess to LB, both being non-negative. The classical heat distribution $Q_{d}$ follows a Gaussian distribution like a classical process in the slow-driving limit Scandi et al. (2020), whereas for the coherent case, due to the presence of non-negative higher order cumulants, $Q_{c}$ can be non-Gaussian.
Further, the authors of Miller et al. (2020) demonstrate the fluctuations in dissipated heat with a two-level system (TLS) described by the following Hamiltonian
$H_{t}=\frac{\epsilon_{t}}{2}(\cos\theta_{t}\sigma_{z}+\sin\theta_{t}\sigma_{x}),$ (4)
which can well approximate the low-energy dynamics of a system in a double-well potential Leggett et al. (1987). Here $\epsilon_{t}$ is energy splitting, $\theta_{t}$ denotes the mixing angle, and $\sigma_{i}$ ($i=x,y,z$) represents the Pauli spin matrices. The thermal dissipation is realized by an adiabatic Lindblad master equation Albash et al. (2012) in the slow driving and weak coupling (to a bosonic heat bath) limit. The competition between the energetic ($\sigma_{z}$) and coherent tunneling is depicted by the mixing angle $\theta_{t}$. $\dot{\theta}_{t}=0$ describes the classical bit whereas $\dot{\theta}_{t}\neq 0$ describes the non-commuting quantum double-well case.
Building on the insights of Miller et al. (2020), recent work Taranto et al. (2023) explored the thermodynamic limits of quantum cooling, juxtaposing LP with Nernst’s unattainability principle. This study introduced a Carnot-Landauer limit, demonstrating that perfect cooling—achieving absolute zero temperature—demands either infinite energy or unbounded control complexity. The cooling bound is given by:
$\left|\Delta E_{H}-\frac{1}{\eta_{c}}\Delta F_{S}^{(\beta)}\right|<\epsilon_{E},$ (5)
where $\eta_{c}=1-\frac{\beta_{H}}{\beta}$ is the Carnot efficiency, $\beta_{h}$ and $\beta$ being the temperatures of the two thermal baths, $\Delta E_{H}$ quantifies the energy transferred to the subsystem from the hot bath, and $\Delta F_{S}^{(\beta)}$ represents the free energy difference between the initial and final states of the system, and $\epsilon_{E}\geq 0$ denotes the error in the cooling protocol.
The role of coherence in finite-time processes becomes even more pronounced when considering many-body effects and collective erasure. The collective effects of many-body systems in finite-time erasure Rolandi et al. (2023) demonstrated that many-body interactions can significantly reduce dissipated work in finite-time thermodynamic processes. Unlike independent qubit erasure, where excess work scales linearly with system size, many-body protocols exhibit sublinear scaling: $W_{\text{diss}}\propto N^{x},\quad x<1$, indicating an accelerated convergence to LB and an enhancement in the efficiency of information erasure.
Building on Miller’s framework, Van Vu and Saito Van Vu and Saito (2022) further examined the interplay between coherence and erasure speed, establishing that finite-time erasure introduces an additional distance cost (distance error to the ground state) beyond the LL. Their findings reinforced the idea that quantum coherence consistently amplifies dissipation, regardless of the control protocol or driving speed. These insights collectively deepen our understanding of the thermodynamic constraints on quantum information processing, emphasizing the inescapable energy costs imposed by coherence, speed, and many-body effects.
Importantly, one can implement an arbitrarily fast erasure process through a highly optimized algorithm that prescribes the exact microstate manipulations needed to transform any initial state into a desired pure final state. If algorithmic complexity is unconstrained, meaning we are allowed arbitrarily complex and precise instructions (programs), then it is in principle possible to design a perfect erasure protocol that acts optimally on any input, driving the system to a unique pure state in finite time. These algorithms would encode a complete understanding of the initial conditions and the exact transformations required, bypassing the usual thermodynamic cost that arises from ignorance or randomness in the input. However, this does not violate Landauer’s principle; rather, it circumvents the physical cost by assuming unbounded computational resources, shifting the “cost” from thermodynamic to computational complexity  Zurek (1989a).
In practice, limitations arise due to thermodynamic speed limits Deffner and Campbell (2017) and coherence-induced excess dissipation Faist et al. (2015), finite computational resources. Recent studies Van Vu and Saito (2022) further quantify the finite-time cost, revealing that even with optimal control, a residual energy cost remains due to quantum fluctuations and irreversibility. Thus, while the Landauer limit can be theoretically approached, the interplay of coherence, control complexity, and finite-time constraints fundamentally prevents its exact realization in realistic settings.
III.2 Non-equilibrium Landauer Process
The miniaturization in modern technology has led to the development of small systems that are out-of-equilibrium in classical Jarzynski (2011); Seifert (2012) as well as in the quantum regime Esposito et al. (2009); Campisi et al. (2011a); Goold et al. (2016). The fluctuation relation Jarzynski (2011, 1997c, 2004); Crooks (1999); Tasaki (2000); Kurchan (2000); Mukamel (2003); Campisi et al. (2011b) plays a promising role in understanding the thermodynamics of these small systems that operate under the non-equilibrium condition where the thermal and quantum fluctuations cannot be neglected. In this review article, we specifically focus on nonequilibrium erasure protocols within the quantum regime.
A key challenge in non-equilibrium thermodynamics is characterizing the dissipation associated with quantum erasure. Reeb and Wolf Reeb and Wolf (2014) established a fundamental lower bound for heat dissipation in equilibrium conditions, highlighting the role of quantum coherence in modifying LP. However, real-world erasure processes often operate in non-equilibrium regimes, where such equilibrium-based bounds may not be directly applicable. To bridge this gap, their framework has been recast from a fluctuation relation perspective, which provides a generalized thermodynamic bound for erasure beyond equilibrium.
In Goold et al. (2015), an erasure protocol that involves a finite-size environment that interacts with the system is proposed. The system is a single qubit system that is coupled to a finite-dimensional environment. A tantalizing fact is that the nonunitality of the open system dynamics leads to a tighter bound for the heat dissipation for the erasure process. This opened the door for the analysis of the cost of computation in non-equilibrium circumstances. Following the same methodology, a comparative analysis of the performance of LB with that of the bound proposed in Goold et al. (2015) for the non-equilibrium condition is addressed Campbell et al. (2017); Zhang et al. (2023). Reeb and Wolf Reeb and Wolf (2014) in their work provided the minimum framework that is required for the Landauer process in the equilibrium condition. The minimum framework for the execution of the Landauer process in the non-equilibrium condition has been addressed in Taranto et al. (2018), where the system and the environment are equivalent to the models in Campbell et al. (2017); Zhang et al. (2023). Although this approach offers valuable insights, it remains limited to particular models and does not yet provide a universally applicable methodology. To rigorously understand LP in the context of quantum non-equilibrium dynamics, further investigation is essential to develop a minimal yet broadly generalizable framework that captures the complexities unique to quantum systems.
The prerequisites for the erasure protocol are (i) A system $\mathcal{S}$ with $H_{\mathcal{S}}$ as the free Hamiltonian of the system. (ii) An environment $\mathbf{E}$ initially uncorrelated with the system, i.e., $\rho_{\mathcal{S}\mathbf{E}}=\rho_{\mathcal{S}}\otimes\rho_{\mathbf{E}}$. (iii) The initial state of the environment will be the Gibbs state. (iv) The interaction of the environment and the system is unitary. The heat fluctuation relation of the environment for the erasure protocol is:
[…]
which follows from the equality proposed by Jarzynski Jarzynski (1997c) using the work distribution. $P(Q)$ denotes the heat distribution of the environment. ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}_{l=jk}}=\sqrt{\lambda_{j}}\langle s_{k}|\hat{U}|s_{j}\rangle$, with $\lambda_{j}$ being the eigenvalue and $|s_{j}\rangle$ being the eigenstate of $\rho_{\mathcal{S}}$. The process is unital iff $\hat{\textbf{A}}=\mathbf{I}_{\mathbf{E}}$. If the operator $\hat{\textbf{A}}$ is expanded in terms of the initial state of both $\mathcal{S}$ and $\mathbf{E}$ under the action of the global unitary evolution, the heat fluctuation relation becomes
$\langle e^{-\beta Q}\rangle=\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}],$ (7)
where $\hat{\textbf{M}}=\text{tr}_{\mathbf{E}}[\hat{U}^{\dagger}\mathbf{I}_{\mathcal{S}}\otimes\rho_{\mathbf{E}}\hat{U}$]. The desired heat dissipation during the erasure process is
$\beta\langle Q\rangle\geq\mathbf{B}_{Q},$ (8)
where $\mathbf{B}_{Q}=-\,\ln(\text{tr}[\hat{\textbf{A}}\rho_{\mathbf{E}}])=-\,\ln(\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}])$ is termed the thermodynamic bound. The operator $\hat{\textbf{A}}$ is dependent on the choice of the state of the system. So, for the execution of the erasure process of a state of choice, one needs to compute $\hat{\textbf{A}}$ for every instance. On the other hand, $\hat{\textbf{M}}$ as defined in Eq. (7) can be easily evaluated just by implementing the unitary interaction. This proposed bound is tested on a physical system where the system is considered to be a single qubit and the environment comprises an interacting spin chain (as shown in Fig. 1). The Hamiltonian of the environment, given by the $XX$ model, is
$H_{\mathbf{E}}=J\sum_{j=1}^{N-1}\left(\sigma_{x}^{j}\sigma_{x}^{j+1}+\sigma_{y}^{j}\sigma_{y}^{j+1}\right)+B\sum_{j=1}^{N}\sigma_{z}^{j}.$ (9)
$J$ denotes the coupling strength of the interspin, $B$ is homogeneous external magnetic field and $\sigma$ describes the Pauli spin operators. The heat dissipation during the execution of the erasure protocol for this physical system is evaluated to be
$\beta\langle Q\rangle$ $=$ $B\sin^{2}(2Jt)(2\alpha^{2}+\tanh(\beta B)-1),$
$\mathbf{B}_{Q}$ $=$ $-\ln\left[2(1-\phi^{2})(\alpha^{2}+p+2p\alpha^{2})+\phi^{2}\right].$ (10)
Here $\phi=\cos(2Jt)$ at time $t$, $p=[1+\tanh(\beta B)]/2$ and $\alpha\in\mathbb{R}$. For $\alpha\approx 1$ and for a particular temperature it is observed that $\mathbf{B}_{Q}\geq\mathbf{B}_{rw}$, where $\mathbf{B}_{rw}$ denotes the bound proposed in Reeb and Wolf (2014).
Figure 1: The schematic of the process where the system $\mathcal{S}$ interacts with the environment $\mathbf{E}$.
A comparison of different forms of the LBs reveals a subtle dependence on the initial state of the system and the environment’s temperature. In the parameter space, sharp boundaries emerged when assessing the relative effectiveness of these bounds Campbell et al. (2017). There were scenarios where the bounds were negative in both frameworks, yet the process’s heat dissipation remained positive. In such cases, the bounds proved to be weaker than the Clausius statement of the second law.
The full counting statistic Esposito et al. (2009) formalism, when applied to analyze the bounds (lower and upper) on the average heat dissipation in an erasure process, a single-parameter bound is developed that can be made arbitrarily tight and is independent of the map being used for the execution of the process Guarnieri et al. (2017). This inherently marks the difference of the lower bound of this process over the previous bound Goold et al. (2015). For this formalism, the minimal set of assumptions is considered as follows in Reeb and Wolf (2014), which validates LP.
The full counting statistics of heat dissipation are defined as the change of energy in the environment that characterizes the mean value of heat dissipation. The lower bound for the heat dissipation using the CGF Rockafellar (1970) is
$\beta\langle Q\rangle_{t}\geq-\frac{\beta}{\eta}\Theta(\eta,\beta,t)\equiv\mathbf{B}_{Q}^{\eta}(t)\quad(\eta>0),$ (11)
with $\eta$ being the counting parameter. Using the large deviation function (useful for studying dynamical phase transitions), which is a powerful tool to study statistical properties for a long time scale Garrahan and Lesanovsky (2010); Lesanovsky et al. (2013); Pigeon et al. (2015), the upper bound for heat dissipation is proposed as
$\beta\langle Q\rangle_{t}\leq\frac{\beta}{|\eta|}\Theta(\eta,\beta,t)\equiv\mathbf{\tilde{B}}_{Q}^{\eta}(t)\quad(\eta<0).$ (12)
The Landauer-like bound here is proposed with one parameter on a two-time measurement protocol. For $\eta=\beta$ the derived form of the bound is equivalent to that of the bound proposed in Goold et al. (2015). This bound is tested on a physical system where the system is considered to be a three-level V-system and the environment is modeled by a two-level system (as in Fig. 2). The transition is pumped by a transition frequency $\Omega_{1}$.
Figure 2: The schematic of the process with a three-level system and a two-level environment.
This physical model highlights the tightness of the proposed bound for heat dissipation during the execution of the erasure protocol.
III.3 Landauer Bound in Non-Markovian Process
The Markovian approximation is the most convenient way to express the dynamics of OQS. In this process, the evolution timescale is considered to be larger than the correlation time of the environment. In other words, in this approximation, the memory effect (or information backflow) is neglected. However, it plays a non-minuscule role in the analysis of the dynamics of the system and thermodynamic cost. There comes the role of non-Markovian (NM) dynamics to explore the dynamics of the system De Vega and Alonso (2017); Breuer et al. (2016). Various methods and approaches have surfaced for defining as well as quantifying non-Makovianity Breuer et al. (2009); Rivas et al. (2010); Chruściński and Maniscalco (2014).
Out of the various methods, one is the applicability of the collision model for NM dynamics Rau (1963); Alicki and Lendi (2007); Scarani et al. (2002); Ziman and Bužek (2005); Gennaro et al. (2009). In the collision model, the NM dynamics for a system can be achieved when the interaction of the system and the environment is intervened by degrees of freedom of the ancilla. Under NM dynamics, where coherence endures and information retrogrades from the environment, a coherence-dependent correction to LB Reeb and Wolf (2014) becomes imperative. Unlike Markovian erasure, where dissipation follows a strict lower bound, NM effects can modify these limits, potentially reducing or amplifying heat dissipation. Here, we extend Landauer’s principle to NM dynamics, exploring how memory effects influence erasure costs and coherence-driven corrections. This provides a more comprehensive view of quantum erasure beyond the standard Markovian framework.
The formulation of a Ladauer-like principle for heat flux in an erasure process delineated by the collision model of OQS is first proposed in Lorenzo et al. (2015). The process of analyzing the information to energy conversion in the collision model provides the platform for the foundation of the LP in NM dynamics.
Figure 3: The system $\mathcal{S}$ interacts with the particle of the environment $\mathbf{E}$ which is termed as the subenvironment. Then the system undergoes pairwise collision with the subenvironment with the restriction that it will not interact with an ancilla of the environment twice.
For the execution of the protocol Lorenzo et al. (2015), the thermalization of the system $\mathcal{S}$ with the environment $\mathbf{E}$ is considered (as shown in Fig. 3). In this model, the environment has N identical noninteracting elements $\mathbf{E}_{N}$, which are conveyed as subenvironments (Fig. 3), and each of them is considered to be in a thermal state $\eta=\otimes_{n=1}^{N}\eta_{n}^{th}$, where $\eta_{n}^{th}=e^{-\beta H_{E_{n}}}/Tr[e^{-\beta H_{E_{n}}}]$. Here, $H_{E_{n}}$ denotes the free Hamiltonian of the n-th subenvironment and $H_{\mathcal{S}}$ denotes the free Hamiltonian of the system. The system will interact with the environment, and the interaction will be a sequence of pairwise collisions with the subenvironments. The environment is of a large size to curtail the situation that the system interacts with the same subenvironment twice. Each collision of this process is described by unitary evolution $U=e^{-igH_{int}\tau}$, where $g$ is the interaction strength, $\tau$ is the collision time, and $H_{int}$ denotes the interaction Hamiltonian of the system and the environment. The information that is stored in the system gets diluted while interacting with the environment. After (n+1) collisions, the state of the system and the environment are respectively described as
$\rho_{n+1}=\text{Tr}_{E}[U\rho_{n}\otimes\eta_{n+1}^{th}U^{\dagger}]=\Phi[\rho_{n}],$
$\eta_{n+1}=\text{Tr}_{S}[U\rho_{n}\otimes\eta_{n+1}^{th}U^{\dagger}]=\Lambda_{n}[\eta_{n+1}^{th}],$ (13)
where $\Phi$ and $\Lambda_{n}$ denote the completely positive trace-preserving (CPTP) map. The variation of energy of the system and the heat exchange with the environment are
$\Delta\mathcal{E}_{n+1}=\text{Tr}[H_{\mathcal{S}}(\Phi-\mathbf{I})[\rho_{n}]],$
$\delta Q_{n+1}=\text{Tr}[H_{\mathbf{E}}(\Lambda_{n}-\mathbf{I})[\eta^{th}]].$ (14)
[…]
$\beta\dot{Q}(t)\geq\dot{S^{\prime}}(\rho),$ (15)
where $S^{\prime}(\rho)=-S(\rho)$. Eq. (15) provides the formulation of the LP in NM dynamics of OQS. The formalism proposed in the work Lorenzo et al. (2015) has twofold power. One, it provides a time-resolved analysis of the erasure by thermalization, which enabled the formulation of the LP for the NM dynamics. Secondly, it was possible to elucidate the role of correlation in the information-erasure processes. This formalism provides the linkage between the erasure process with the intrasystem correlations that arise in the open quantum dynamics for multipartite systems.
The violation of the LB due to strong correlation (proposed in Lorenzo et al. (2015)) is encountered in the spin-1/2 particle system Pezzutto et al. (2016) when used as the framework for the analysis. The sequence of discrete-time collisions of the system with one of the environmental particles at a time occurs in this process, and the Hamiltonian is ruled by the Heisenberg Hamiltonian. In the long time limit, the environment (for non-interacting particles) undergoes homogenization dynamics, i.e., in the long time limit which is equivalent to a large number of collisions of the system with the environment, the state of the system approaches asymptotically to the initial preparation of the environment Ziman et al. (2001); Scarani et al. (2002). If the environment is considered to be composed of interacting particles, the system undergoes an NM process. Here, the elements of the environment are in a thermal state, and in the asymptotic limit homogenization dynamics are encountered. This behavior occurs when the state of the elements of the environment has weak fluctuation. Due to the inter-environment interactions, there is a memory effect in the dynamics of the system, which features the non-Markovianity. It was observed that there is an instantaneous violation of the LB for the system, the cause of which is the strong system-environment correlation.
Figure 4: The system $\mathcal{S}$ interacts with the auxiliary system $\mathcal{S}_{1}$, which then interacts with the environment $\mathbf{E}$ and is redirected to its initial value after each collision. The auxiliary system and the environment form the non-Markovian environment.
The explicit analysis of the cause and the condition for the violation of the LP in the NM environment is explicated in Man et al. (2019). A modified form of the collision model is considered in Man et al. (2019), where the system ($\mathcal{S}$) information is first transferred to the subenvironment via the system-subenvironment collision. A part of this information gets transferred to the next subenvironment via intra-collision. This intra-collision leads to the system-environment correlation, and then the system interacts with the next subenvironment. So, there is prior information of the system in the subenvironment before the next system-subenvironment collision. This causes the non-Markovianity in the process. The system-environment collision is governed by a unitary operation $U$. The change in entropy of the system due to the interaction with the subenvironment is
$\Delta S_{n}$ $=$ $S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})+\text{Tr}_{E_{n}}[(\rho_{\mathbf{E_{n}}}^{\prime}-\rho_{\mathbf{E_{n}}}^{th})\,\,\ln\,\rho_{\mathbf{E_{n}}}^{th}]$ (16)
-- $I(\rho_{\mathcal{S}\mathbf{E_{n}}})-S(\rho_{\mathbf{E_{n}}}||\rho_{\mathbf{E_{n}}}^{th})),$
where $S(\rho||\rho^{\prime})=\text{Tr}\,\rho\,\ln\,\rho-\text{Tr}\,\rho\,\ln\,\rho^{\prime}$ describes the quantum relative entropy, $\rho_{\mathcal{S}_{n}}^{\prime}$ and $\rho_{\mathbf{E_{n}}}^{\prime}$ denotes the marginal states of the system and the environment respectively after the collision. $I(\rho_{\mathcal{S}\mathbf{E_{n}}})$ characterizes the mutual information that quantifies the system-environment correlation based on the intra-collision strength. The second term in Eq. (16) (followed from Eq. (2)) describes the entropy flow from the environment to the system. The LB for the NM process is
$\beta_{\mathbf{E_{n}}}\Delta\tilde{Q_{n}}$ $=$ $\Delta\tilde{S}_{n}+S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\,\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-I(\rho_{\mathcal{S}\mathbf{E_{n}}})$ (17)
-- $S(\rho_{\mathbf{E_{n}}}||\,\rho_{\mathbf{E_{n}}}^{th})),$
where $\Delta\tilde{Q_{n}}=-\Delta Q_{n}$ and $\Delta\tilde{S}_{n}=-\Delta S_{n}$. When $I(\rho_{\mathcal{S}\mathbf{E_{n}}})\leq S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\rho_{\mathbf{E_{n}}}^{th}))$, the LP holds as long as the established system-environment correlations are smaller than the upper bound. So, the condition that allows the system to violate the LP is
$I(\rho_{\mathcal{S}\mathbf{E_{n}}})>S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\,\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\,\rho_{\mathbf{E_{n}}}^{th})).$ (18)
It can be generalized, where the system is coupled to a composite environment with $M$ different environments. The LP violation condition for the multiple NM environments is found to be equivalent to that of Eq. (18).
A solution to the violation of the LP has been put forward in the work Zhang et al. (2021). Here, the authors considered a different scenario, the system ($\mathcal{S}$) interacts with $\mathcal{S}_{1}$ (ancillary system) which in turn is coupled to a Markovian environment $\mathbf{E}$. This composite environment induces NM dynamics (Fig. 4). The memory in this model is $\mathcal{S}_{1}$. In the Markovian limit, the conventional LB holds. The modified version of the LB for the heat dissipation that holds for the NM as well as the Markovian regime is
$\beta\dot{Q}_{\mathcal{S}}(t)\geq\dot{\tilde{S}}(\rho_{\mathcal{S}}(t))+\dot{I}(\rho_{\mathcal{S^{\prime}}}(t))+\dot{S}(\rho_{\mathcal{S}_{1}}(t)||\rho_{\mathcal{S}_{1}}^{th}),$ (19)
where $\dot{Q}_{\mathcal{S}}(t)$ describes the heat flux from the system to the environment, $\dot{\tilde{S}}(\rho_{\mathcal{S}}(t))$ is the entropy flux from $\mathcal{S}$ to $\mathbf{E}$ ($\tilde{S}(\rho_{\mathcal{S}}(t))=-S(\rho_{\mathcal{S}}(t))$), $\dot{I}(\rho_{\mathcal{S^{\prime}}}(t))$ describes the rate of mutual information that characterizes the correlation between the system and auxiliary system and $\dot{S}(\rho_{\mathcal{S}_{1}}(t)||\rho_{\mathcal{S}_{1}}^{th})$ describes the rate of quantum relative entropy. The generalization of the model for multiple environments has also been reported in this work. The modified LP for the multiple environment case is also found to be valid in both regimes and is a direct consequence of Eq. (2).
Another aspect that was encountered for violation of the LP in the NM domain is the information back-flow. In the work Hu et al. (2022), the authors evaluated the connection between the information back-flow in the NM dynamics and LP. For the analysis, a qubit is considered to be coupled with the environment. If the system is initially in the thermal state, it is inferred from the analysis that there is a one-to-one correspondence between the violation of the LE and the information backflow. Whereas the correspondence does not hold if the initial state of the system has coherence.
IV Landauer limit in Computing
A quantum computer Nielsen and Chuang (2002); Aharonov (1999); DiVincenzo (2000) harnesses quantum mechanical phenomena such as superposition and entanglement Mermin (1990); Linden et al. (2006); Wineland (2013) to perform computations. Qubits, serve as the fundamental units of information in quantum computers and leverage the principles of quantum superposition enabling quantum computers to execute certain types of calculations much more efficiently than classical computers. However, regardless of the computational task, it ultimately needs to be implemented on a physical system, implying that computations are subject to the constraints of the laws of physics. For instance, the LP places a minimum limit on the heat generated during bit erasure, and the quantum speed limit Golub and Ortega (2014); Caneva et al. (2009); Okuyama and Ohzeki (2018); Jones and Kok (2010); Deffner and Campbell (2017); del Campo et al. (2013); Deffner and Lutz (2013) dictates a limitation on how quickly a fundamental logical operation can be carried out. The modern computer uses a binary logic system, and LB for such binary computers is known to be $k_{B}T\ln 2$.
IV.1 Landauer’s Limit For N-based logical computer
A computer using the binary system can be exemplified as a single particle Szilard engine Szilard (1929) where the certainty of the particle’s presence in a chamber corresponds to the recording of 1 bit of information, and the uncertainty in the particle’s location corresponds to the erasure of 1 bit of information. However, computers are not limited to binary logic systems. In principle, they can be based on many-valued logic. For instance, a ternary logic-based computer (“trit”) is founded on the concept of a ternary symmetrical number system and ternary memory element (“flip-flap-flop”) Glusker et al. (2005); Brousentsov (1965); Stakhov (2002); Frieder et al. (1973); Knuth (1998). Recently, computers employing many-valued logic have garnered significance due to fundamental aspects and numerous applications Gottwald and Gottwald (2001); Chang (1958).
The exemplification of the LP for binary logic computers can be supplied by the Brownian particle in a double potential well as shown in Fig. 5. For a symmetrical well and a random bit $a_{0}=a_{1}=\frac{1}{2}$ (where $a_{0}$, $a_{1}$ denotes the probability) the LB is recovered by implementing Eq. (2). Now, whether LL holds for N-based logic is explored in Bormashenko (2019b). In the work, the authors have considered ternary logic, i.e., trit computing element for the analysis which is further generalized for N-based logic. Similar to binary logic computers, the LP is illustrated for ternary logic computers using a Brownian particle in a symmetric triple-well potential. For a random bit, $a_{0}=a_{1}=a_{2}=\frac{1}{3}$ (where $a_{0}$, $a_{1}$, $a_{2}$ denotes the probability), the LB is reached for the ternary logic by implementing Eq. (2), and this result has been further extended to N-based logic.
Figure 5: Schematic of the qubit model with a Brownian particle $M$ placed in a double potential well. The qubit can be trapped in either of the sides of the well. The corresponding information states are $m=0$ and $m=1$.
As computational algorithms become more complex, parallel computing becomes essential for the efficient execution of protocols. Parallel computing, as discussed in Kumar et al. (1994); Barney et al. (2010); Melhem (1992); Golub and Ortega (2014), involves the simultaneous utilization of multiple processors to solve algorithms. The primary objective is to distribute the workload among several processors to solve problems more swiftly or handle larger problems within the same timeframe. In the study Konopik et al. (2021), the authors examined the energy cost of finite-time irreversible computing using non-equilibrium thermodynamics principles. It has shown that the energy cost for computational tasks in a parallel computer, within a given finite time, closely adheres to the LL. This cost remains bounded even as the computational problem size increases.
IV.2 Landauer’s bound in the presence of time-symmetric protocol
Investigations into energy dissipation with time-symmetric protocols, demonstrates a fundamental trade-off between computational accuracy and energy cost Riechers et al. (2020). It shows that reducing logical error requires increasing energy dissipation, particularly under nonreciprocal operations, and thus time-symmetric protocols prevent computation from reaching the LB. In a subsequent work Wimsatt et al. (2021), authors utilized thermodynamic analysis of time-symmetric procedures (as prescribed in Riechers et al. (2020)) to thoroughly examine the trade-offs between accuracy and dissipation during information erasure. The authors employed nonequilibrium information thermodynamics to calculate the minimum energy dissipation needed for reliable erasure under time-symmetric control protocols. The energy costs associated with reliable erasure were found to be higher than those implied by the LB on information erasure. Moreover, these costs diverge in the limit of perfect computation. Therefore, the creation of time-asymmetric protocols is deemed necessary for effective and precise thermodynamic computation. Consequently, it is concluded that time asymmetry serves as a crucial design principle for thermodynamically efficient computing, warranting further investigation.
V Experimental Validation of Landauer’s Principle
With our ever-increasing control over the miniaturized scale and the advancement of quantum technology, the validation of the LB is becoming more and more plausible day-by-day. Here in this section, we will focus on the recent experimental advancement to test LP in different technological platforms both in the stochastic and quantum realm.
V.1 Optics Based Technology
Optical Tweezers:
The test of LP was possible (that remained untested over five decades) after the two basic advancements. One of them was to develop a method to investigate the work done on the particle as well as the heat dissipation by the particle based on the information on the trajectory of the particle and its potential. It was proposed and tested in the seminal paper by Sekimoto Sekimoto (1997, 2010). The second advancement is the development of methods to impose user-defined potential on small particles. For e.g., the usage of the localized potential force. This potential is created by optical tweezers and is used to explain LP under partial erasure Bérut et al. (2012). In this work, the authors have considered an overdamped colloidal particle inside the double potential well (Fig. 6) which is created by focusing a laser alternatively at two different positions with a high switching rate. The form of the potential well is determined by the intensity of the laser and the distance between the two focal points. If the particle is on the left side of the well, the state of the system is denoted by ‘0’, whereas if the particle is on the right side, it is ‘1’. The experimental process for this method can be summarized as follows:
$\bullet$ Initially the bead is considered to be trapped in one of the wells with a definite state. The central barrier is kept high so that the jumping time is very high.
$\bullet$ Now the intensity of the laser is reduced so that the barrier is low enough and the bead can jump from one to another.
$\bullet$ Finally, after the bead ends in the required well independent of the initial state (this causes the memory erasure of the process when it is set to 1), the barrier is raised to its previous stage.
Figure 6: The erasure protocol that is considered in the experiment.
Following the same methodology, the authors in their work Jun et al. (2014) have embraced a more flexible approach with a feedback loop to create virtual potential. An anti-Brownian electrokinetic (ABEL) feedback trap is used for testing LP. This model provides the advantage of measuring the work with high precision for testing LP. Berut et. al. Bérut et al. (2012) in their work were not able to acquire the complete erasure as they did not have full control over the potential shape. It was reported that in the asymptotic limit, LB is $\approx(0.13-0.49)k_{B}T$ whereas the full-erasure limit is $0.69k_{B}T$. Jun et. al. Jun et al. (2014), in their work, reports the complete erasure and shows that their approach results in reaching LB. A complete and detailed analysis of the various contribution that causes heat dissipation in the system is reported Bérut et al. (2015) in verifying LP.
In Gavrilov and Bechhoefer (2016), the authors have explored the erasure principle for the symmetry-breaking case by analyzing an asymmetric double-well potential. The analysis of this process, following the methodology of Sagawa and Ueda Sagawa (2014), conveyed that to erase a bit of information the average work can be less than $k_{B}T\ln 2$ provided that the volume of the phase space for each state is different. The memory cell in the experiment consists of an overdamped silica bead trapped in the double-well potential foist by ABEL trap. It is encountered in the experiment that the work for this asymmetric bit erasure can be less than that of $k_{B}T\ln 2$.
Interferometer:
To explore the LP for the underdamped and overdamped systems Dago et al. (2021); Dago and Bellon (2022), differential interferometer Paolino et al. (2013) have been considered as the platform for the analysis. The working system is a micromechanical oscillator (the role played by a conductive cantilever in the experiment) confined in a double potential well. Here, the cantilever is the memory cell. Within the framework of stochastic thermodynamics, it has been shown in the work Dago et al. (2021) that one can reach the LB with high precision within a time scale of 100 ms which was previously recorded to be 30 s. The work was further extended by the authors to explore the overdamped and underdamped case within the stochastic thermodynamic framework with fast operation Dago and Bellon (2022). There they encountered a transient temperature rise, so the mean work to erase the information increases but is still bounded by LP.
Magneto-Optical Kerr effect:
Nano-magnetic switches are one of the prime components that will play an extensive role in electronic applications like storage media. The information in such devices is encoded by electron spin. It is a bistable switch that comprises elongated ferromagnetic dots of various sizes and shapes. The authors in Martini et al. (2016) have investigated the energy cost, i.e., the LL of resetting the magnetic binary switches composed of elliptical or rectangular ferromagnetic dots of different sizes and shapes. The dissipated energy in this process is measured by the vectorial magneto-optical Kerr effect (MOKE) experiment in Permalloy (Ni80 Fe20Martini et al. (2016). The logic states ‘0’ and ‘1’ are described by the two orientations of magnetization. The experimental process for this method can be summarized as follows: (a) The process starts with the equilibrium magnetization. The system is in either of the states, and (b) several magnetic fields are applied to reduce the effect of the barrier as well as help the system jump from one to another to reach the final logic state.
In this model, the authors encountered that there is a deviation from the theoretical limit of up to three orders for magnetic dots with dimensions of several hundred nanometers. The morphological imperfection as well as the inhomogeneity of the magnetization are the primary causes of the deviation. Whereas if one reduces the dot size, it is shown to approach the theoretical LL.
Following the same technology, the authors in Hong et al. (2016) have explored the intrinsic energy dissipation for a single-bit operation. They have used a nano-scale digital magnetic memory as their working system. The MOKE experimental setup is considered for the analysis of the energy dissipation during the execution of the process. In this process the nanomagnet plays the role of the memory bit, and magnetic anisotropy is utilized to create the easy axis along which the net magnetization aligns to minimize magnetostatic energy.
So far the experimental validation has been restricted in the classical domain. The LB in the quantum regime was experimentally analyzed in the work Gaudenzi et al. (2018), where a crystal of molecular nanomagnet as the spin memory is considered for the analysis. In this model, the crystal of Fe8 molecular magnet Gatteschi et al. (2006) plays the role of quantum spin memory which measures the energy dissipation during the execution of the erasure process. The process is equivalent to the method used in Martini et al. (2016) to develop the double potential well and reset the system to the final state ‘1’ by applying a magnetic field. The erasure of the memory is still governed by the LP. Surprisingly, it was encountered that maximum energy efficiency is achieved while preserving the fast operation process unlikely to the classical system.
V.2 Trapped Ions
In a quantum regime where the information is encoded in a qubit, one needs to reconstruct the model that is considered in the classical regime to be applicable in the quantum realm. The work Yan et al. (2018) has explored the quantum LP based on trapped ion $Ca^{+}$. The ion is trapped in a linear Paul trap. Here, the LB is evaluated by the analysis of the system-reservoir correlation and the change in entropy during the execution of the erasure process. Trapped ions are considered as one of the perfect platforms for the exploration of quantum thermodynamics with high accuracy An et al. (2015); Huber et al. (2008); Roßnagel et al. (2014). The two internal levels of the system ion are considered as the qubit system and the vibrational modes of the ion as the finite temperature bath. The LP is analyzed in the system by observing the phonon number in the variation of the ion. The authors have confirmed that the LP holds in the quantum regime experimentally.
V.3 Nuclear Magnetic Resonance Technology
The process to measure the heat dissipation in quantum logic gates using the nuclear magnetic resonance (NMR) setup is proposed in Peterson et al. (2016). For the analysis, a three-qubit system is considered (the working system, environment, and the ancilla) to evaluate the heat dissipation of the process. In the first step of the process, the interferometric technique is considered for the reconstruction of the dissipated heat using the ancilla. In the second phase of the process, the change in the entropy of the system is measured through quantum state tomography Cramer et al. (2010); Christandl and Renner (2012); Lvovsky and Raymer (2009); Gross et al. (2010); Stricker et al. (2022); Liu et al. (2012). The system is developed by dissolving trifluoroiodoethylene (C2F3I) molecule in D6 (97%). The three ${}^{19}F$ nuclear spin forms the three-qubit system for the analysis. The extracted average heat during the execution of the process is found to be bounded by the LP.
V.4 Superconducting Technology
An experiment has been performed on a hardware platform namely, superconducting flux logic for analyzing the quantum LP Saira et al. (2020). The double well potential in this process arises due to the Josephson effect and the flux quantization. The bit erasure process is explored in the parametric regime where the approximation of metastable state is valid. It is observed that the process is bounded by LB.
VI Reversible Computation Model and Thermodynamic Interpretation
In this section, we first provide a brief overview of the reversible model of computation Keyes and Landauer (1970); Likharev (1982); Bennett (1982, 1989); Landauer (1961); Feynman (2018); Richard (1986); Zurek (1989b); Raussendorf and Briegel (2001); Bennett and Landauer (1985): Ballistic computer and Brownian computer mainly, followed by their thermodynamic interpretation. We first discuss the Ballistic computer (BLC) proposed by Fredkin and Toffoli Fredkin and Toffoli (1982) and its limitations. Subsequently, we discuss the Brownian computer (BWC) which utilizes thermal fluctuation to perform a computational process.
Before delving into reversible computation and its theromdynamic interpretation it is worthwhile to briefly overview the two key concepts: thermodynamic reversibility and logical reversibility. Thermodynamic reversibility Jarzynski (1997c); Crooks (1998, 1999); Jarzynski (2000); Seifert (2005); Kawai et al. (2007); Crooks (2011); Wolpert et al. (2024) can be defined as follows: A physical process is considered thermodynamically reversible if and only if the time evolution of the probability distribution in the process can be reversed. This reversal should include the time reversal of changes in external parameters, along with the inversion of the signs of both work and heat quantities. On the other hand, logical reversibility Landauer (1996); Bennett (1973, 2003); Sagawa (2014) is defined as follows: A computational process is logically reversible if and only if it is an bijection. In other words, for any output logical state, there is a unique input logical state. Otherwise, it is considered logically irreversible. According to Landauer Landauer (1961), a positive amount of heat emission is inevitable while a logically irreversible process occurs where information is erased or thrown away. These two fundamental concepts are crucial in the analysis of thermodynamic computation processes, underscore the pivotal role of thermodynamics in computational theory. How thermodynamic reversibility influences the heat generation in an logical irreversible process, for example the information erasure is summarized in Table 1.
Quasi-static Finite velocity
Thermodynamically Reversible Irreversible
Heat emission $=1/\beta\ln 2$ $>1/\beta\ln 2$
Table 1: Role of thermodynamic reversibility in logical irreversible processes.
VI.1 Ballistic Computer
The principle of the ballistic computation model Fredkin and Toffoli (1982) is based on elastic collisions. This model consists of hard spheres that collide between themselves elastically and with fixed reflective barriers. From the input side of the model, as shown in Fig. 7 (starting line here), a huge number of hard spheres (balls) are fired with equal velocity. If ‘1’ is there in the input, a ball is considered in the starting line else no ball for ‘0’. Due to the collision process inside, the ball changes its direction and collides with the other balls. The balls, after a finite number of collisions, reach their finishing point. This signifies the output of the computer. The mirror of this computer is equivalent to the logic gates of our digital computers, and the balls are equivalent to the signals.
Figure 7: A schematic of a BLC proposed by Toffoli. The key condition for this model is that the number of 1’s in the input must be equal to the number of 1’s in the output, ensuring that the Boolean function is both conservative and reversible in nature.
Any bijective function is computable in this model, but it will be unable to compute non-conservative (non-bijective) Boolean functions Øgaard (2021). Though it conveys to decrease in the amount of cost in energy, we encounter some drawbacks of this setup. Two main drawbacks of this system are its sensitivity to small perturbative changes, and the second one is related to the collision of the balls. It is quite difficult to make each collision elastic. Furthermore, these collisions result in the thermal randomness in the system.
To address the collision problem, one approach is to correct the instability in the velocity and position of the ball after each collision process. While this provides a solution, it renders the system thermodynamically irreversible. Another method to mitigate this effect is to use square balls instead of spherical ones. This approach eliminates the exponential growth of errors as the square balls remain parallel to the wall and each other. However, it is worth noting that the use of square balls is unnatural due to the non-existence of square atoms in nature.
Quantum effects can stabilize the system from this problem, but it will bring some new instability Benioff (1982). The wave-packet spreading causes instability in the system in the quantum realm. Benioff Benioff (1982) in his work has discussed a quantum version of the BLC, where he has proposed a way to culminate the effect of the noise due to the wave packet spreading by utilizing a time-independent Hamiltonian.
VI.2 Brownian computer
As thermal randomness is inevitable, the strategy of Brownian computers Bennett (1982) is to exploit it. In this model, the trajectory of the dynamical part of the system is influenced by thermal randomization in such a way that it attains Maxwell velocity distribution and is equivalent to a random walk. Despite its chaotic nature, the BWC is able to execute valuable computations.
The state transition for the BWC happens due to the random thermal movement of the part that carries the information. Due to its random nature, the transition can backtrace (move backward) in the computational process, undoing the transition executed recently, albeit the transition is slightly biased towards the forward direction. In the macro regime, the execution of computation using a BWC seems counterintuitive, but this is an obvious situation in the micro regime.
Bennett has proposed Bennett (1982) that one can execute a Turing machine (TM (see Sec. VII for TM)) using this thermal randomness. It is made up of clockwork, which is frictionless and rigid in its form. The parts of the clockwork TM should be interlocked so that they have the freedom to jiggle around locally, but are restricted from moving an appreciable amount for the execution of a logical transition. Bennett presumed that a driving force (some energy gradient) to execute the computation in less time and a trap for the stability of halting state are required, and they can be done with arbitrarily small entropy generation which maintains the effective reversibility of the model. However, in  Norton (2013), the authors argued that it has an entropic cost which renders the model irreversible. In the context of computational complexity, a comparable model of BWC was analyzed by Reif Reif (1979) to explore the relationship between P and PSPACE (P = PSPACE ) Arora and Barak (2009).
VI.3 Brownian computer: Thermodynamic interpretation
The thermodynamic analysis of the Brownian motion of particles, which are integral to the BWC, has been approached through various processes Norton (2013); Nicolis and De Decker (2017); Pal and Deffner (2020); Meerson et al. (2022); Lee and Peper (2010); Peper et al. (2013); Lee et al. (2016); Utsumi et al. (2022). In this context, we will specifically examine the thermodynamic properties of the BWC using a simplistic model proposed in Norton (2013). The discussion begins with a concise overview of the expansion of a single-molecule gas. Subsequently, Brownian computers with different constraints are explored within this expansion model. The analysis leads to the inference that Bennett’s assertion regarding the thermodynamic reversibility for the operation of BWCs is not tenable.
Single molecule gas expansion: Let’s contemplate an ideal single gas molecule situated at a specific temperature $T$ within a spacious chamber divided into $n$ parts by partitions, each with a volume $V$. In the initial phase, the gas molecule resides in the first cell with a volume $V$, as illustrated in Fig. 8(a). Subsequently, the partitions are removed, allowing the single gas molecule to expand to a larger volume throughout the chamber.
The system Hamiltonian is $H=\mathbb{L}(p)$, where $p$ is the momentum of the molecule and $\mathbb{L}$ is a quadratic function of $p$. So, the entropy for the system is evaluated as $S=\frac{\partial}{\partial T}(k_{B}T\,\,\ln Z)=k_{B}\,\ln(nV)+C_{p}(T)$, where the contribution from the momentum perspective is included in the constant $C_{p}(T)$, $n$ denotes the total number of chambers, and $Z$ is the partition function of the system.
Brownian computer: From a thermodynamic perspective, a BWC can be likened to a single-molecule gas expansion. In our discussion, we will differentiate between driven BWCs (where an external force propels the system) and undriven BWCs. Additionally, the introduction of a trap (a slight energy gradient to confine the molecule), as depicted in Fig. 8(b), enhances the entropic force driving the system.
In the case of the undriven BWC, it mirrors the single-molecule expansion, but its drawback lies in its lack of computational utility. The final equilibrium state in this scenario is uniformly distributed across all computation stages. Conversely, introducing a trap causes the molecule to be confined by the trap potential, resulting in a non-uniform final state. This alteration in the system increases its computational utility.
To accelerate the computational process, external energy (drive) needs to be supplied to the system. Among the considered configurations, the driven BWC with the trap is currently the most resource-intensive in terms of thermodynamic (irreversible) entropy required to propel the system.
Figure 8: A schematic of the BWC: a) without energy trap, b) with energy trap.
In summary, from the thermodynamic analysis of the Brownian computer, it can be deduced that the $n$-chamber BWC is fundamentally a thermodynamically irreversible process, characterized by a minimum entropy amount of $k\,\,\ln n$. Introducing various parameters, such as an energy trap and external driving, to the system results in entropy production that renders the model irreversible.
Bennett’s misidentification of BWC as a reversible thermodynamic process can be attributed to the focus on tracking internal energy rather than thermodynamic entropy. Analyzing thermodynamic reversibility solely based on the tracking of internal energy is misleading. The essential condition for verifying whether a process is reversible lies in tracking the total entropy of the system. If the total entropy ($S_{system}+S_{environment}$) remains constant throughout the process, it signifies a thermodynamically reversible process. Bennett and Landauer’s oversight of not tracking the BWC’s total entropy resulted in the misidentification that BWC model is a thermodynamically reversible process.
It has been indicated in recent studies Utsumi et al. (2022, 2023) that the thermodynamic cost of executing a computational process becomes less significant when employing a token-based Brownian circuit for computational cycles. This stands in contrast to the case of logically reversible Brownian TM, where entropy production is directly proportional to the logarithmic function of the state space.
VII Thermodynamics of computational models
The foundations of computer science are based on algorithms, data structures, and computation theory. In computer science, models of computation serve as mathematically precise frameworks for describing automated processes of symbolic reasoning. These models are diverse rather than singular, encompassing various approaches. A foundational concept in theoretical computer science is the existence of a class of models termed ‘Turing complete’ or ‘universal’. These models exhibit two key properties: (i) mutual equivalence and (ii) broader generality compared to non-equivalent models. Equivalence implies that any computation representable within one model can be translated seamlessly into another, and vice versa. Turing-complete models possess the capability to execute computations from non-Turing-complete models, but the reverse is not necessarily true.
Here we focus on the two primary aspects of computation: (a) The finite automata/finite state machine (FA/FSM), which belongs to the class of non-universal models, and (b) the Turing machine (TM) which belongs to the class of universal models. In the following, we briefly describe them. Subsequently, in the latter half, the thermodynamic aspects of these two computational models will be discussed.
VII.1 Mathematical foundations
The primary thermodynamic aspect of computation is the energetic cost of computation Maroney (2009); Faist et al. (2015); Parrondo et al. (2015); Kolchinsky and Wolpert (2017); Boyd et al. (2016); Ouldridge and Ten Wolde (2017); Boyd et al. (2018); Wolpert (2019); Wolpert and Kolchinsky (2020); Riechers and Gu (2021a, b); Kolchinsky and Wolpert (2021); Kardeş and Wolpert (2022). Delineated by the enormous energetic cost of computation, the urge of considering a computational metric of success involving the resource cost of the computation has found renewed attention of late Auffeves (2022). The estimation of the thermodynamic cost of computation is based on the following axioms Li and Vitányi (1992):
Axiom 1: No thermodynamic cost for a reversible computation process.
Axiom 2: Any irreversible process (irreversibly bit provided or deleted) that occurs in a computation process has a thermodynamic cost.
Axiom 3: For a reversible computational process, where the input set $\eta$ is replaced by the output $\zeta$, the set $\eta$ ($\zeta$) is not provided (deleted) irreversibly.
Axiom 4: All physical computations are considered to be effective (i.e., it boils down to the formal notion of TM computation).
Based on the axioms stated above the thermodynamic cost Li and Vitányi (1992) has been computed in terms of the computational complexity, called the Kolmogorov complexity (KC) Li et al. (2008); Vitányi (2013) of the bit string, which quantifies the shortest possible description (or program) that can generate a given string using a TM. The thermodynamic cost of a computation is determined by counting the number of bits that are irreversibly provided or erased. This measurement accounts for information compression to ensure an optimal representation of the computational records.
The KC of computing a bit string $\zeta$ from the initial bit string $\eta$ is expressed as
$K{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\zeta|\eta)}=min\{|p_{c}|+\,|W_{i}|:\xi_{i}(\bar{p}_{c},\bar{\zeta})=\eta,\,p_{c}\in\mathbb{S}^{\star}\}.$ (20)
Here $p_{c}$ denotes the program which is a finite sequence of symbols (or a bit string) belonging to the set $\{0,1\}^{*}$, $W_{i}$ is the enumeration of the TM, and ||$\bullet$|| ∙ | is the cardinality of the bit string. The cardinality of the program $p_{c}$ is the length of the bit string representing the program $p_{c}$. The Turing machine indexed by $W_{i}$ computes $\zeta_{i}$. The enumeration of the partial recursive function defined as $\xi_{i}(\bar{p}_{c},\bar{\zeta})=\eta$ is an effective invertible bijection from $N\times N$ to $N$, which effectively maps inputs (including the program and input bit string) to outputs (see Appendix A). This formulation quantifies the minimal computational effort required to transform $\zeta$ into $\eta$, taking into account both program length and machine description length.
Theorem 1: The thermodynamic cost $E(\eta,\zeta)$ of computing $\zeta$ from $\eta$ is given by
$E(\eta,\zeta)\approx K(\zeta|\eta)+K(\eta|\zeta).$ (21)
Outline of the proof of this theorem is provided in Appendix A.
The axioms introduced in Li and Vitányi (1992) were designed to establish a framework for analyzing the thermodynamic cost of computational machines using KC. KC is a purely mathematical measure that does not consider the physical processes involved in executing a computation. Since it focuses solely on the minimal program length required to generate a specific output, it does not account for the energy cost associated with resetting a computational machine. As a result, the process of resetting has been deliberately excluded from the axiomatic formulation developed for this purpose.
Furthermore, Zurek has shown Zurek (1989b) that the KC provides the energetic bound of individual computations. A generalized version of Zurek’s bound has also been established Kolchinsky (2023), which is applicable to all quantum as well as classical computations, including both stochastic and deterministic ones. The bound of the thermodynamic cost $E$ of computing $\zeta$ from $\eta$ reads:
$\frac{\beta}{\ln 2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E}\geq K{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\zeta|\eta)}-\log_{2}\frac{1}{p(\zeta|\eta)}-K(\mathbf{P})+\gamma_{\mathcal{C}},$ (22)
where $\log_{2}\frac{1}{p(\zeta|\eta)}$ is the noise associated with the computation, $K(\mathbf{P})$ denotes the KC of the protocol $P$, and $\gamma_{\mathcal{C}}$ is the additive constant which depends on the universal computer $\mathcal{C}$ (independent of $P$). A practical physical setup, inspired by the widely used “two-point measurement” schemes in quantum thermodynamics, is considered. In this framework, a computational subsystem $A$ executes the transformation $\eta\rightarrow\zeta$, while the subsystem $B$ serves as a bath. The protocol $P=(V,U,\beta,\epsilon)$ encompasses the chosen product basis $V$ for the two subsystems, the unitary $U$, the inverse temperature $\beta$, and the bath’s energy function $\epsilon$. The Kolmogorov complexity is calculated using the definition in Eq. (20).
In Zurek (1989b), Zurek conveyed that the loss of algorithmic information can be quantified in terms of KC of the shortest protocol possible. It can be considered as a “algorithmic fluctuation theorem” relating the second law of thermodynamics and the Physical Church-Turing thesis Kolchinsky (2023).
VII.2 Finite State Machine
VII.2.1 Finite Automata: Basic Aspect
First, consider a natural example of an automaton. Imagine a toll gate controlled by a computer. Assume the gate remains closed until the required amount, say $25$ bucks, is paid. Moreover, assume that there are three sets of coins only: $5$, $10$, and $25$ bucks. Now, let us consider a situation where the driver of the vehicle inserts 25 bucks in the sequence (5, 5, 10, 5). The state of the machines evolves as follows:
$q_{0}\rightarrow q_{1}\rightarrow q_{2}\rightarrow q_{4}\rightarrow q_{5}.$
The state diagram with all possible combinations is shown in Fig. 9. The gate opens or the computation ends if and only if the accepted (or halt) state (here $q_{5}$) is reached.
Figure 9: Transition diagram of the FA-controlled toll gate. $q_{0}$ is the start state. The state $q_{5}$ shown in two circles represents the final state (or the halt state). The insertion of the coins shown by arrows represents the alphabet.
Let us explain the steps of this controlled toll gate with a mathematical definition of FA below:
Definition 1. A FA is a 5-tuple $M=(\mathcal{Q},\,\Sigma,\,\delta,\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{0}},\,F)$. Here,
1. $\mathcal{Q}$ represents a finite set whose elements are the states of the system.
2. $\Sigma$ also represents a finite set whose elements are called the alphabets (a finite set of symbols).
3. $\delta:\,\,\mathcal{Q}\times\Sigma\rightarrow\mathcal{Q}$ represents the transition function.
4. $q_{0}\in\mathcal{Q}$ represents the start state.
5. $F\subseteq\mathcal{Q}$ describes the set of accepting (or halt) states.
The transition function $\delta(q_{i},b)$ takes an input $q_{i}\in Q$ and alphabet $b\in\Sigma$ (in Fig. (9) the sets of coins constitute the alphabet set $\Sigma$) and determines the output state $q_{j}\in Q$. The machine has to record the state at any instant in time to determine the next step, i.e., to perform another transition or halt. A sequence of alphabets leading to halt states is called a word. The collection of the words forms a language of the FA, which happens to be a regular one (any finite language is called a regular language) for FA.
DFA means the current state and the current symbol uniquely determine the next state. NFA means the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution. PFA means their is a conditional probability distribution over all possible next states given the current state and the current symbol.
FA have different forms like deterministic (DFA: the current state and the current symbol uniquely determine the next state), non-deterministic (NFA: the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution), and probabilistic finite automaton (PFA: there is a conditional probability distribution over all possible next states given the current state and the current symbol). In the stochastic (or probablistic) automaton, the single-valued transition function $\delta$ will be replaced by a conditional probability distribution. One can also observe multiple accept states described in the literature as ‘terminal states’ for the system Lawson (2003). One can even encounter multiple start states in the process.
FA has a wide range of applications in computer science, like designing hardware, designing compilers, network protocols, and computation Lawson (2003). Furthermore, FA also has great impacts across different fields, including biology, mathematics, logic, linguistics, engineering, and even philosophy Bird and Ellison (1994); Baer and Martinez (1974); Straubing (2012). However, here we will focus on the computational aspects of FA and the role of thermodynamics in it.
VII.2.2 FA: Thermodynamic Aspect
Though the reversible models of computation demonstrated by Bennett have vanishing thermodynamic costs following Axiom 1, they have major drawbacks for practical applications. All the reversible computation models take either infinite time to run or return a result with a very high probability of error. Recently, there has been exploration on the thermodynamic cost analysis of computational models in the quasi-static limit Wolpert (2015); Strasberg et al. (2015); Wolpert et al. (2023); Gopalakrishnan (2023) that are inadequate to estimate the energetic cost of computations in real scenarios due to long execution time and huge error probability. Therefore, the thermodynamic cost analysis of practical models of computation is of utmost interest. However, only very recently have there been a few explorations in this direction Chu and Spinney (2018); Ouldridge and Wolpert (2022); Manzano et al. (2024); Ouldridge and Wolpert (2023).
Universal computation machine (aka TM, see details below) requires infinite tape, i.e., infinite memory at its disposal, which often makes the model unsuitable for practical applications. On the other hand, finite state machines (FSM), aka FA, albeit non-universal, that use finite resources, are an alternative model as real-world computers have limited resources. The construction of a thermodynamically efficient model Chu and Spinney (2018); Ouldridge and Wolpert (2022); Kardeş and Wolpert (2022) of FSMs has been in the limelight in recent times to estimate the energetic cost of such models.
The appeal for studying FSMs offers an intriguing perspective: each state transition can be seen as a fundamental unit of computation, termed an elementary cycle. Understanding the thermodynamics of these elementary units enables one to grasp the thermodynamics of any FSM, as any FSM can be viewed as a sequence of these cycles. Moreover, these elementary cycles can be dissected further into more basic computational actions. Remarkably, only two types of basic computational steps are necessary for implementing any FSM: namely, the generalized versions of bit flips and bit sets. To grasp the thermodynamics of FSM, in Chu and Spinney (2018) the authors constructed a thermodynamically consistent FSM model by designing FSM as a time-inhomogeneous Markov chain Wolpert et al. (2019).
While investigating the energy consumption and the probability of accurate computation with the designed FSM model, Chu et. al. Chu and Spinney (2018) were able to infer that in the high accuracy regime of the FSM, the probability of error scales polynomially, while the implementation cost, measured in terms of the work required for the cycle, increases only logarithmically. When expressed in terms of energy differences between states of FSM, it is observed that the average work scales linearly with $\Delta E$ (the energy difference), as expected, while the error decreases exponentially. Essentially, it indicates that the model can achieve perfect accuracy, albeit at the cost of infinite energy dissipation. However, quasi-deterministic computation with practically negligible error probabilities can be accomplished at a finite, even modest energy expense. Nevertheless, particularly for high accuracies, it is evident that the proposed model dissipates energy well beyond the theoretical limit.
Intriguingly, in the high accuracy limit, the size of the input alphabet and the size of the machine cease to significantly impact the cost of computation. One might speculate that a larger tape alphabet enables more information processing per computational step with only marginal increases in energy costs compared to smaller alphabets. This suggests that it may be more efficient to operate Markovian computers with larger alphabets rather than smaller ones. However, it’s noteworthy that the error probability, in the worst-case scenario, can be observed to depend on both the size of the machine and the size of the alphabet.
This observation conveys that for any algorithm, multiple FSM implementations exist, some favorable in accuracy and energy efficiency within an error tolerance, others less so, thus presenting an intriguing possibility of trade-offs between performance, energy consumption, and accuracy, primarily determined by implementation rather than solely by the physical framework and the computation to be performed.
Following this direction, it has also been addressed Ouldridge and Wolpert (2022) that the cost for the computational characterization of the DFA divides regular language into two classes. One is the invertible local map and the other is the non-invertible local map. In the former case, zero minimal cost is encountered, whereas in the latter case, high cost is encountered.
An alternative approach other than the Markov chain model is addressed in Kardeş and Wolpert (2022), where they have developed a thermodynamic framework to define logical computers like DFA without specifying any extraneous parameters (like rate matrices, Hamiltonians, etc.) of the process that is considered to implement the computer. This framework doesn’t require the entropy production to be zero and is derived from an exchange fluctuation theorem Crooks (1999); Jarzynski (2000); Peliti and Pigolotti (2021); Esposito and Van den Broeck (2010). In particular, they use the Myhill-Nerode theorem Lewis and Papadimitriou (1998); Hopcroft et al. (2001) to prove that out of all DFAs which recognize the same language, the “minimal complexity DFA” is the one with minimal entropy production for all dynamics and iterations.
VII.3 Turing Machine
VII.3.1 Turing Machine: Basic Aspect
In 1936, Alan Turing proposed an abstract computation device Church (1937), later coined as the Turing Machine (TM), that can investigate the extents and limitations of all computable functions Hopcroft and Motwani (2000); Savage (1998). Church-Turing thesis Copeland (1997) states that “A function on the natural numbers is computable by a human being following an algorithm, ignoring resource limitations, if and only if it is computable by a Turing machine.” In the Physical Church-Turing thesis Piccinini (2011); Cotogno (2003), it has been further modified on the physical ground that the set of functions which one can compute by utilizing the mechanical algorithmic methods and abides by the laws of physics Pour-El and Richards (1982); Moore (1990); Wolpert (2019); Arrighi (2019); Wüthrich (2015), are also computable with the help of TM [^1] As all computational devices are physical, it has been argued in some works Baaz et al. (2011); Aaronson (2005) that one might bring some restrictions to the foundation of physics by utilizing the properties of the TM.. Various forms of definitions of the TM exist in the literature, which are computationally equivalent to each other. The formal definition of the TM is
Definition 2. A TM is defined by 7-tuple ($\mathcal{Q}$, $\Lambda$, $\Sigma$, $\delta$, $q_{0}$, $F$, $q_{r}$). Here,
1. $\mathcal{Q}$ is a finite set that describes the non-empty set of states.
2. $\Lambda$ is a finite set depicting the input alphabets.
3. $\Sigma$ represents a finite set of tape alphabet and $\Lambda\,\subseteq\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Sigma}$.
4. $\delta:\mathcal{Q}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Sigma}\,\rightarrow\,\mathcal{Q}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Sigma}\times\{L,R,S\}$ is called the transition function. Here, $\{L,R,S\}$ describes the direction of the movement of the head of the tape. Based on the command, the head moves left, right, or stays in the same position on the tape.
5. $q_{0}$ ($q_{0}\in\mathcal{Q}$) represents the start state of the Turing machine.
6. $F$ is called the accepted state or the halting state.
7. $q_{r}$ is called the rejected state.
In other definitions of the TM, one can encounter multiple sets of halting states.
Figure 10: A schematic representation of a TM includes an infinite tape, where the current state is indicated by $q$. This tape is segmented into equally spaced square boxes, each containing symbols from the tape alphabet. The TM scans the tape using its tape head, which can move either left or right along the tape.
At each step of the computation, the state of the TM reads the alphabet in the square where the tape head is placed and subsequently moves on to a new state $q^{\prime}$ (see Fig. 10). It writes a new alphabet ($a^{\prime}$) on the tape, and then moves its tape head either to the left or to the right. This process is repeated until the system attains the accepted state. Mathematically, this map can be expressed as $(q,a)\rightarrow\,(q^{\prime},a^{\prime},d)$, where $d$ denotes the right or left movement of the tape head. For a given TM, the arguments of the transition states are called instantaneous descriptions (IDs) of the TM. One can also encounter TM’s with no halts.
Earlier models like FA and push-down automata  [^2] The main difference of push-down automata with FA is that it has access to the top of the stack to decide which transition to take. are not accurate in recognizing the language Arora and Barak (2009). On the contrary, the TM is considered to be the most accurate model.
TM has a great impact on the analysis of computational complexity Hopcroft and Motwani (2000); Moore and Mertens (2011); Arora and Barak (2009); Sipser (1996); Li et al. (2008) and even in philosophy Copeland et al. (2013). One of the most important open problems in computer science that remains to be explored by the TM is whether $P=\,NP$ Lipton and Regan (2013); Razborov and Rudich (1994); FORTNOW (2003). Limitations of mathematics, like Gödel’s incompleteness theorem Gödel (1931) still remain to be one of the main challenges to solve by TM.
VII.3.2 TM: Thermodynamic Aspect
TMs hold a central position in computation theory as a complete model of computation, unlike FSMs/FAs, which are non-universal models. This significance has prompted researchers to focus on the thermodynamic analysis of TMs in order to design computationally efficient models from a thermodynamic perspective.
For the thermodynamic analysis of TM, one needs to design a thermodynamically efficient model of TM. Now, if one considers a single-tape TM, where the input tape is overwritten with the output tape, the computation becomes irreversible. Thus, to achieve reversibility in a TM, it’s necessary to have at least two tapes—one for input and the other for output. In a reversible TM, it should be possible to retrace the computational path and retrieve the initial state of the TM. This requirement underscores the importance of maintaining the original information intact throughout the computation process.
A logically reversible TM was proposed by Bennett Bennett (1982), where he showed that a reversible TM needs at least four times the number of steps that are needed for the execution of a computation in an irreversible TM. Further generalization to Bennett’s approach has been considered in Strasberg et al. (2015). While Bennett considered a single input ($\mathcal{S}_{inp}\rightarrow\mathcal{S}_{out}$), the authors of Strasberg et al. (2015) developed a TM which processes a continuous stream of input string on an infinite tape such as: ($\dots$, $a$, $\mathcal{S}_{inp}^{\prime}$, $a$, $\dots$, $a$, $\mathcal{S}_{inp}$, $a$, $\dots$). The input string denoted by $\mathcal{S}_{inp}^{\prime}$, $\mathcal{S}_{inp},\dots$ are separated by blank space symbols $a$, which denote the beginning and the end of the input string. Thus, the output string can be described as ($\dots$, $a$, $\mathcal{S}_{out}^{\prime}=\mathbb{U}(\mathcal{S}_{inp}^{\prime})$, $a$, $\dots$, $a$, $\mathcal{S}_{out}=\mathbb{U}(\mathcal{S}_{inp})$, $a$, $\dots$), where $\mathbb{U}:\mathcal{S}_{inp}\rightarrow\mathcal{S}_{out}$.
The multiple-tapes TM model proposed in Strasberg et al. (2015) consists of four tapes: input, output, working, and history tape, respectively; and a computational cycle with five stages. The working and the history tape comprise the TM, while the input and the output are provided externally. The five stages of the computational cycle are: a) copy the input into the working tape, b) computation, c) copy the output into the output tape, d) retract the working tape to retrieve the input via the history tape, and finally e) erase the working tape.
The dynamics of this logically reversible TM is modeled by a continuous Markov process that corresponds to a set of computational steps $\eta$ with probability $p_{\eta}$ that changes according to the $1^{\text{st}}$ order Markovian master equation Breuer et al. (2002); Rivas and Huelga (2012); Rotter and Bird (2015)
$\frac{d}{dt}p_{\eta}(t)$ $=$ $-\left(\mathcal{W}_{\eta+1,\eta}+\mathcal{W}_{\eta-1,\eta}\right)p_{\eta}(t)$ (23)
$+$ $\mathcal{W}_{\eta,\eta+1}p_{\eta+1}(t)+\mathcal{W}_{\eta,\eta-1}p_{\eta-1}(t),$
where $\eta\in\mathbb{Z}$, $\mathcal{W}$ is the rate matrix satisfying $\sum_{\eta^{\prime}}\mathcal{W}_{\eta,\eta^{\prime}}=0\quad\forall\eta$, and $\{\mathcal{W}_{\eta,\eta+1},\mathcal{W}_{\eta,\eta-1}\}$ describe the forward and the reverse rate respectively, which obeys the detailed balance condition: $\frac{\mathcal{W}_{\eta,\eta^{\prime}}}{\mathcal{W}_{\eta^{\prime},\eta}}=e^{-\beta(E_{\eta}-E_{\eta^{\prime}})}$. The rate matrix $\mathcal{W}$ is decomposed into blocks for each input during the computation. The transition between different blocks of the rate matrix is prohibited during the computation process.
For the thermodynamic cost analysis of this model, the physical system is associated with an energy landscape along the computational path Strasberg et al. (2015), say, the logical and the successor states differ in energy by an amount of $\epsilon_{c}>0$. The rate of entropy production for the computation is
$\dot{S}(t)$ $=$ $\frac{d}{dt}\mathcal{H}(t)+2\Gamma_{c}\,\beta\epsilon_{c}\,\sinh\left(\frac{\beta\epsilon_{c}}{2}\right)\geq 0,$
where $\Gamma_{c}$ denotes the rate setting for the overall time scale of the problem, $\mathcal{H}(t)=1/2\ln(4\pi\Gamma_{c}\epsilon_{c}\cosh{\beta\epsilon_{c}/2})$ is the Shannon entropy of the distribution. In the limit, $\epsilon_{c}\rightarrow 0$, the rate of entropy production tends to zero, which confirms that this TM model works in a thermodynamically reversible manner in the steady-state regime.
However, the small entropy production rate doesn’t confirm that the overall entropy production is zero. Like in the computational cycle, an unavoidable cost is encountered while resetting the TM. This is due to the increase in the Shannon entropy during the computation. This cost is dependent on the number of computational steps, and this verifies Norton’s notion Norton (2013, 2014) of thermodynamic irreversibility of computation.
Following various driving schemes Parrondo et al. (2015); Van den Broeck et al. (2013); Esposito and Van den Broeck (2010), a model to analyze the thermodynamic cost of the TM is proposed in Kolchinsky and Wolpert (2020). They have considered stochastic thermodynamics Wolpert (2019) for the analysis of the dynamics of these physical processes. Interested readers can go through the review article Wolpert (2019), which provides a detailed analysis of the stochastic thermodynamics in different aspects of computation.
In Kolchinsky and Wolpert (2020), the authors propose a different approach to analyze the thermodynamics of TM by combining techniques from algorithmic information theory and stochastic thermodynamics. A discrete state system (equivalent to the input and output of the TM) is coupled to a reservoir at temperature $T$ and it evolves under the influence of the driving protocol. Three kinds of thermodynamic costs are encountered for this TM model:
(1) The heat generated during the execution of the realization of TM will be processed for each input $z$. It is denoted as $Q(z)$.
(2) The heat generation for the entire computation that maps the input $z$ to the output $y$. This cost is referred to as the thermodynamic complexity of $y$.
(3) The average heat $\langle Q\rangle$ produced by a TM realization, computed over the input distribution that minimizes entropy production.
Two physical processes are considered for the realization of the TM. The first physical process that is considered for the analysis is the coin-flipping process for the universal Turing machine (UTM). This physical model is a thermodynamically reversible model, where the input is samples of the ‘coin-flipping’ distribution $p(z)\propto 2^{-l(z)}$, where $l(z)$ depicts the string length. The heat generated in this physical process is proportional to the computation program to execute the input $z$.
Being motivated by the physical Church-Turing thesis, a semi-computable process (coined as domination realization) is considered as the second physical process. It is shown that the second physical process is ‘optimal’ in the sense that the heat generated by this process for any input $z$ is smaller than or equal to any other computable realization of TM on $z$.
The methods discussed thus far provide valuable insights into the thermodynamic aspect of physically realizing TMs. Considering the centrality of TMs to both physics and computer science, there exists a need for additional exploration to develop a more feasible and realistic model for the physical implementation of TMs.
VII.4 Quantum Computation: Energy cost
Thus far, we have explored the energy costs associated with traditional computational models. In today’s rapidly evolving digital economy, computing processes are consuming energy at an accelerating pace Liu et al. (2023a); Lannelongue et al. (2021). The widespread adoption of machine learning algorithms, large-scale language models, and data-intensive operations has driven an unprecedented surge in energy demand Patterson et al. (2021); Arora and Kumar (2024); Scholten et al. (2024); An et al. (2023). This escalating trend highlights the urgent need for energy-efficient alternatives to conventional computation. With the increasing deployment of large-scale AI models, energy consumption has become a critical concern in the modern computing industry.
Given these challenges, the question arises: Could quantum computing provide a viable and energy-efficient alternative to classical computing? Quantum computers, which leverage the principles of superposition and entanglement, have the potential to perform certain computational tasks exponentially faster than their classical counterparts. If harnessed effectively, quantum computing could significantly reduce the energy footprint of complex computations, offering a path toward a more sustainable and efficient computing paradigm, and is a subject of ongoing research Preskill (2018).
Quantum computation is widely expected to outperform classical computation across various computational resources. However, establishing a clear and definitive advantage in energy consumption remains an intricate challenge. This difficulty stems from the lack of a robust theoretical framework that directly correlates the physical concept of energy with the computational complexity of quantum algorithms. In classical computing, energy dissipation is inherently linked to irreversible operations, governed by LP. In contrast, quantum computing is fundamentally grounded in unitary evolution and reversible computation, making direct comparisons between the two paradigms highly nontrivial.
Despite these challenges, recent advancements have made notable strides in bridging this gap. Researchers are actively pursuing both theoretical and experimental avenues to uncover the energy-efficiency benefits of quantum computing Meier and Yamasaki (2023); Góis et al. (2024); Green et al. (2022); Pandit et al. (2022); Ikonen et al. (2017); Martin et al. (2022); Paler and Basmadjian (2022). On the theoretical front, efforts are focused on formulating precise energy-complexity relationships for quantum algorithms, shedding light on the fundamental trade-offs between computational power and energy cost. It has been shown that quantum computing can offer a substantial amount of energy savings over classical methods for specific problems like– Simon’s problem Meier and Yamasaki (2023), and the Fourier transform algorithm Góis et al. (2024). Meanwhile, experimental investigations leverage state-of-the-art quantum processors, such as IBM’s quantum hardware Desdentado Fernández et al. (2021), to empirically assess energy consumption in practical quantum computations. These studies seek to provide compelling evidence that, for specific computational tasks, quantum computers exhibit superior energy efficiency compared to their classical counterparts—bolstering the case for quantum supremacy in the realm of energy-efficient computation.
VIII Thermodynamics of Error Correction
During the communication or storage, bits are prone to noise, which tampers them. Therefore, the primary challenge in communication and storing is to detect these errors and reduce their influence on the information sent or stored. Here, in this section, we explore the process to nullify the errors that occur during the communication or storage process, both in classical as well as quantum regimes.
In the quantum regime, the initial research was primarily focused on developing quantum codes Steane (1996a, b); Knill and Laflamme (1997); Gottesman (1998a, b); Bennett et al. (1996); Knill and Laflamme (1997); Lieb et al. (1961) that provided a rigorous framework for error correction Bennett et al. (1996); Knill and Laflamme (1997); Calderbank et al. (1998). Now, advanced concepts like fault-tolerant quantum computation Shor (1996); DiVincenzo and Shor (1996); Gottesman (1998a) provides the route-map to the threshold theorem for error correction in the quantum regime Knill et al. (1996); Aharonov and Ben-Or (1997).
VIII.1 Classical Error Correction
In a communication process, the data is transmitted from the sender to the receiver end through a channel susceptible to noise, commonly referred to as a noisy channel. The data string belongs to the set $\mathbb{S}$. The communication string undergoes encoding with the addition of extra bits (redundant bits). Upon reaching the receiver, the original message is reconstructed by processing the potentially corrupted message (due to a bit flip). This reconstruction process is known as decoding.
In the late 40’s of ${}^{\text{th}}$ century, the seminal work of Shannon Shannon (1948) led to the foundation of this field and was extended by Hamming in his work Hamming (1950). Since then, this field has gained importance for developing better communication protocols. The extent to which error correction (EC) of the missing bits is possible depends on the design of the error-correcting code (ECC). Generally, there exist two types of ECC, they are block code Jafarkhani (2005); Adler et al. (1983); Feltstrom et al. (2009) and convolutional code Dholakia (1994); Forney (1970); Alfarano et al. (2023) as depicted in Fig. 11. Here we focus on a subfield of block code: Linear code. There are other models of error correction codes that are not covered here, interested readers can go through Pless (1978); Hoffman et al. (1991) for further information.
Figure 11: A schematic representation of the domain of the classical EC.
The formal definition of the ECC is:
Definition 3. The ECC is defined as an injective map from $n$ symbols (messages bits) to $m$ symbols (code bits):
$Enc:\Lambda_{n}\rightarrow\Lambda_{m},$
where $\Lambda$ represents the set of symbols.
$\bullet$ The domain of the set, i.e., $\Lambda_{n}$ represents the message space, and $\Lambda_{m}$ represents the encoded message (codeword). Here $n$ denotes the message length.
$\bullet$ Block length: It refers to the length of the codeword which is mapped to $m$-bit strings.
$\bullet$ Code: A set of codewords produced by encoding messages. In general, $m\geq n$.
$\bullet$ Rate: It is defined as the ratio of $n$ over $m$. It quantifies the efficiency of the code.
A linear code is generally called a $[m,n,d]$ code, where $m$ describes the length of the codeword, $n$ denotes the length of the message string, and $d$ describes the minimum hamming distance. The hamming distance between two given vectors is given by the number of positions the corresponding vectors differ.
VIII.2 Quantum Error Correction
Classical EC is a well-developed theory based on the demand for better communication systems. One-to-one mapping from classical to quantum error correction (QEC) is not possible as the quantum world has some constraints of its own, like qubits are governed by the no-cloning principle Nielsen and Chuang (2002). As a consequence even the simple repetition code [^3] In the repetition code, the encoding is done as, for example, $00\rightarrow 0$ 000 and $11\rightarrow 1$ 111. It will correct a corrupted state, say $010$ to its majority value, $000$ in this case., which belongs to the linear block code class, does not work in the quantum domain. Additionally, the phenomenon of wavefunction collapse upon measurement makes quantum unique from classical. In the seminal work of Peter Shor Shor (1995) the first QEC protocol was proposed. Shor in his work has demonstrated that quantum information can be encoded by exploiting the idea of entanglement of qubits. Works in this direction Calderbank and Shor (1996); Preskill (1998); Kitaev (1997); Knill et al. (1998); Gottesman (1998a) have demonstrated that one can suppress the error rate in the quantum regime provided the qubits meet some physical conditions.
A general qubit can be represented
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (24)
where $\alpha$ and $\beta$ represents complex number satisfying $|\alpha|^{2}+|\beta|^{2}=1$. Thus, a qubit has the power to encode information in the infinite number of possible superposition of the computational basis states, which are denoted by $|0\rangle$ and $|1\rangle$. Therefore, the qubits are subjected to an infinite number of errors. However, due to the digitization of the errors using the Pauli operator, the error counts are reduced to two fundamental errors Nielsen and Chuang (2002). One is the $X$-type error, which is the bit flip error (similar to classical), and the other is the $Z_{pau}$-type error, which is the phase error.
Figure 12: A schematic of the $[[m,n,d]]$ stabilizer code. $|\psi\rangle_{In}$ represents the quantum data register, and $|0\rangle_{m-n}$ represents the redundancy qubits. $H$ is the Hadamard operator.
Similar to the classical linear code, the stabilizer code in quantum Steane (1996c); Calderbank et al. (1997, 1998); Gottesman (1996, 1997) is represented as $[[m,n,d]]$ (as shown in Fig. 12). Here $m$ represents the total count of physical qubits, the count of the logical qubits is given by $n$, and $d$ describes the code distance which determines the number of correctable errors. A stabilizer code encodes $n$ logical qubits into $m$ physical qubits. The stabilizer represents an abelian subgroup of the $m$-fold Pauli group. The notation of the quantum codes is in double brackets to differentiate it from the classical code, which is shown by a single bracket.
Here, we described the basic intuition of QEC that is required in the latter half to understand the thermodynamic interpretation of EC. Interested readers who want to explore more about QEC can go through the reviews Gottesman (2010); Devitt et al. (2013); Lidar and Brun (2013); Terhal (2015) in this direction, which covers QEC and its subfields.
VIII.3 Thermodynamic Interpretation
From a thermodynamic perspective, EC is analogous to a refrigeration process. In the seminal work Vedral (2000), Vedral has performed a thermodynamic analysis of EC both in classical and quantum domains, incorporating an MD-based model of EC. In Cafaro and van Loock (2014), the authors, building upon Vedral’s result, have extended it for approximate QEC when the observation on the system is imperfect, implying sub-optimal information gain. An alternative formalism is explored in Korepin and Terilla (2002) to establish the conditions of quantum codes and investigate QEC conditions from a thermodynamic standpoint.
In Vedral (2000), the ``$0"$ and ``$1"$ states are represented by whether a single molecule of an ideal gas is in the LHS or RHS of a box, respectively, like in Bennet’s MD setup. Say, initially the molecule is in the LHS or 00 state. If it expands isothermally, it works $\Delta W=k_{B}T\ln 2$ at the expense of the same amount of free energy stored in it. This increases its entropy by $\Delta S=k_{B}\,\ln 2$.
If the molecule now jumps to the RHS, with probability $\frac{1}{2}$ say, then we say an error has occurred, which means the molecule has lost the power to perform work. To restore the work power we need to compress the molecule at one side of box and for the same we need to do $\Delta W=k_{B}T\ln 2$ amount of work.
Let us restore the work capacity of the molecule or correct the error with the help of another molecule (as shown in Fig. 13)
(1) Let’s consider that initially, the molecules are on the LHS and RHS of the respective boxes.
(2) Now consider that some error occurs to the particle in the box $\mathcal{A}_{\alpha}$.
(3) Now $\mathcal{B}_{\alpha}$ correlates itself (by some means whose details are not important) with $\mathcal{A}_{\alpha}$ by observing it, such that the molecules of both systems either occupy the LHS or the RHS of their respective boxes.
(4) Based on the state of the system $\mathcal{B}_{\alpha}$ one will move the system $\mathcal{A}_{\alpha}$ to its respective side. This leaves the system $\mathcal{B}_{\alpha}$ in a randomized state. Effectively we correct $A_{\alpha}$ by transferring the error to $B_{\alpha}$.
(5) In the last step, the system $\mathcal{B}_{\alpha}$ is brought back to its initial state by isothermal compression, and it requires work to be done.
In summary, to correct $A_{\alpha}$ or reduce its entropy $B_{\alpha}$’s free energy is wasted just like how a refrigerator works. Now to restore $B_{\alpha}$’s free energy its entropy ($\Delta S=k_{B}\,\ln 2$) has to be dumped into the environment and it needs at least $\Delta W=k_{B}T\ln 2$ amount of work to be done which makes it exactly similar to the LP, thereby consistent with the second law of thermodynamics.
Figure 13: A schematic of the reversible cycle of the classical error correction protocol.
A similar protocol of EC for pure states in the quantum domain is:
(1) The initial joint-state is given by $|\psi_{co}\rangle|m\rangle|env\rangle$, where $|\psi_{co}\rangle$ describes the encoded state, $|m\rangle$ represents the state of the measurement device, and $|env\rangle$ is the environmental state.
(2) The combined-state after the introduction of the error by its environment is described as $\sum_{j}Err_{j}|\psi_{co}\rangle|m\rangle|env_{j}\rangle$, where $\{Err\}$ is the error operator and $|env\rangle_{j}$ is the orthogonal environmental state.
(3) Next, the environment is traced out leaving the combined state as $\sum_{j}Err_{j}|\psi_{co}\rangle\langle\psi_{co}|Err_{j}^{\dagger}\otimes|m\rangle\langle m|$.
(4) Next, when the system is observed, correlation is generated between the measurement apparatus and the system resulting the comnibed state $\sum_{j}Err_{j}|\psi_{co}\rangle\langle\psi_{co}|Err_{j}^{\dagger}\otimes|m_{j}\rangle\langle m_{j}|$, where $\langle m_{j}|m_{k}\rangle=\delta(j,k)$. It means the observation is perfect, otherwise, perfect recovery of the system would not have been possible.
(5) In this step, the correction of the error is executed. The state of the combined-system is $|\psi_{co}\rangle\langle\psi_{co}|\otimes\sum_{j}|m_{j}\rangle\langle m_{j}|$. Note that the state is not equivalent to the initial state, as a resetting operation of the measuring device is required.
(6) To reset it, a garbage system is included giving the joint state $|\psi_{co}\rangle\langle\psi_{co}|\otimes\sum_{j}|m_{j}\rangle\langle m_{j}|\otimes|m\rangle\langle m|$. Now, by swapping the garbage and the measurement device, we get $|\psi_{co}\rangle\langle\psi_{co}|\otimes|m\rangle\langle m|\otimes\sum_{j}|m_{j}\rangle\langle m_{j}|$. Thus, the setup is reset and ready for another cycle of QEC.
Note that the same entropy analysis as the classical EC is applicable here. The system, measuring device, and the garbage system play the role of system $A_{\alpha}$, $B_{\alpha}$, and the environment that sinks the entropy of $B$ respectively. This thermodynamic analysis of QEC can be extended even for a mixed state Vedral (2000).
If one considers imperfect measurement in step 4 of the above protocol for the execution of QEC, then only approximate recovery of the corrupted state is possible. This problem is studied in detail in Cafaro and van Loock (2014), where the authors have introduced some ancilla qubits to keep track of the errors. They have argued that it is like an refrigeration process, where the error (entropy) is transferred from the data-qubits to the ancillary-qubits, thats purify (cools down) the data-qubits.
Recent studies have found a similarity between the quantum heat engine and QEC codes Landi et al. (2020); Danageozian et al. (2022). The authors have fortified their intuition by conducting a comprehensive analysis of the thermodynamic properties of the quantum engine-based ECC.
IX Miscellaneous
In this section, some complementary topics in the LP and thermodynamic aspects of computation are collated here.
IX.1 Landauer bound of Electronic Circuit
In the context of electrical circuits, LP pertains to the process of resetting a bit in a digital memory element, such as a flip-flop or a register Lambson et al. (2011b); Keyes (1988); Hanggi and Jung (1988). To overcome the thermal fluctuations that exist at finite temperatures, energy must be wasted to reset a bit to a reference state. This energy loss due to resetting leads to an increase in the circuit’s temperature, potentially causing additional losses due to leakage current and other factors. The LB of electronic circuits poses a significant constraint on the efficiency of digital computation and thus plays a crucial role in the design of low-power electrical circuits and the advancement of energy-efficient computing systems. Various methods Freitas et al. (2021); Gopal et al. (2022); Freitas et al. (2022) to minimize energy dissipation in electronic circuits have been explored lately.
In Freitas et al. (2021), a comprehensive theory is proposed for nonlinear electronic circuits affected by thermal noise. These circuits encompass devices with arbitrary I-V (current-voltage) curves but are subject to shot noise Sivre et al. (2019); Djukic and Van Ruitenbeek (2006). This proposed theory incorporates a large class of electronic circuits, namely tunnel junctions, diodes, and metal-oxide semiconductor (MOS) transistors in sub-threshold operation. By considering the stochastic nonequilibrium thermodynamics of these circuits, the authors of Freitas et al. (2021) formulate the thermodynamics of computing. The irreversible entropy production in such circuits is expressed in terms of thermodynamic potentials and forces. Specifically, the authors analyze a stochastic model of a subthreshold complementary metal-oxide-semiconductor (CMOS) inverter or NOT gate and derive an analytical solution for the steady state in the Markovian limit. In a nutshell, the investigation in Freitas et al. (2021) delves into how nonequilibrium thermal fluctuations impact the transfer function of the gate, utilizing the solution derived from the master equation.
IX.2 Landauer bound in switch protocols
Conventional CMOS technology faces the challenge of generating excessive heat during computation, far exceeding the theoretical bound. This limitation poses a barrier to the advancement of switches. In pursuit of next-generation switches for advancing computer technology, the scientific community has turned its attention to mechanical switches Jang et al. (2005); Cha et al. (2005); Fujita et al. (2007); Jang et al. (2008a, b). In Neri et al. (2015), the authors utilized molecular dynamics simulations to explore the minimum energy needed for reset and switch protocols in a bit encoded by compressed clamped-clamped graphene buckled ribbon.
Another alternative technology, bistable nanomagnetic switches, offers the ability to store information with low heat dissipation, where each logic state corresponds to a distinct equilibrium orientation of magnetization. In Madami et al. (2014), the authors conducted virtual experiments based on quasistatic micromagnetic simulations at a fixed temperature in practical nanomagnetic switches Cowburn and Welland (2000); Imre et al. (2006); Csaba et al. (2002) to explore minimal energy consumption during a reset operation. Their findings confirm that the LB is accurately achieved for elliptical switches composed of elongated nanomagnets with lateral sizes below 100 nm, provided that the erasure technique employed is slow and occurs over an appropriate time interval.
IX.3 Computer as a heat engine
The study of the thermal machine is one of the primary perspectives of thermodynamics in classical Carnot (1872); Martini (1983); Walker et al. (1985a, b); Van Wylen and Sonntag (1985); Reed (1898); Barton (2019); Scovil and Schulz-DuBois (1959); Szilard (1929) and quantum regime Kim et al. (2011); Kosloff (2013); Roßnagel et al. (2016); Martínez et al. (2016); Chattopadhyay and Paul (2019); Uzdin et al. (2015); Chattopadhyay et al. (2021a); Chattopadhyay (2020); Mohan et al. (2024); Santos and Chattopadhyay (2023); Mukhopadhyay et al. (2018); Sur et al. (2024); Das et al. (2019); Naseem et al. (2020); Chattopadhyay et al. (2021b); Pandit et al. (2021); Singh et al. (2020, 2023). In a simple sense, one can convey that the Carnot engine extracts an amount of heat $Q_{hot}$ from the hot reservoir at temperature $T_{hot}$ and transfers an amount of heat $Q_{cold}$ to the sink at temperature $T_{cold}$. The work done to execute this process is $W_{hot\rightarrow cold}=Q_{hot}-Q_{cold}$. Optimal efficiency is observed when $\mathcal{L}_{hot}-\mathcal{L}_{cold}=0$, where $\mathcal{L}_{hot}=-Q_{hot}/T_{hot}$ is the negentropy that is imprinted in the engine. So the loss that occurs in the execution of the process is just throwing away the negentropy of the amount $Q_{cold}/T_{cold}$. The concept of a computer being equivalent to a Carnot cycle has been explored Carnot (1872); Costa de Beauregard (1989); Brillouin (1962); Prigogine and Nicolis (1985) and it is inferred from the results of Brillouin (1962); Costa de Beauregard (1989) that an ideal computer encounters a zero work balance, while the information delivered during a process is $\ln 2\,dl=d\mathcal{L}_{hot}-d\mathcal{L}_{cold}$ (where $dl$ denotes the entropic cost of processing information). In contrast, while the ideal computer’s operation suggests a loss of negentropy, it also emphasizes the potential for harnessing information to perform work, challenging traditional views of thermodynamic limitations. This duality underscores the complex relationship between information and energy in physical systems.
A physical model has recently emerged, which focuses on autonomous quantum thermal machines for computational analysis Lipka-Bartosik et al. (2024). These machines are comprised of interacting bits that are connected to baths with distinct temperatures, and are referred as “thermodynamic neurons.” In this setup, the machine undergoes evolution to a non-equilibrium steady state, and the computation output is determined by the temperature of an auxiliary finite-size reservoir. This model exhibits versatility and can be utilized to implement various linearly separable functions, like NOT and NOR gates.
IX.4 Nonergodic systems and it’s thermodynamics
For information erasure, Landauer argued that the system would confront a decrease in entropy while estimating the minimal dissipation by introducing an operation restore-to-one (RTO), whereas in Ishioka and Fuchikami (2001), the authors have demonstrated that there is no change in the thermodynamic entropy even after the RTO operation.
To support this assertion, the authors devised a thought experiment involving a particle confined in a bistable-monostable potential well interacting with a heat reservoir. This model, termed quantum flux parametron, was introduced by Goto et. al. Shimizu et al. (1989); GOTO et al. (1996). In the analysis, the state of the system will be described as ‘1’ when the particle is found on the RHS of the potential well and ‘0’ when one finds the particle on the left. The schematic representation of the thought experiment is depicted in Fig. 14.
Figure 14: A schematic of the thought experiment is shown. The red solid sphere is the single atom of the system.
The first four steps of the thought experiment are called the erasing process, and the final three steps represent the writing process. If RTO is applied to the system, and the system is in the state zero, the same configuration will be observed subjected to a condition that the state is known before the execution of this experiment. Landauer argued that the system will observe a decrease in entropy after RTO operation, but it is inferred in Ishioka and Fuchikami (2001) through the lens of Clausius’s definition of thermodynamic entropy that the erasure process characterized by a transition from a nonergodic to an ergodic state is fundamentally irreversible and entails the production of entropy. In contrast, heat generation predominantly occurs during the writing process. Remarkably, the reverse of erasure, namely, a transition from an ergodic to a nonergodic state, can be interpreted as a form of spontaneous symmetry breaking, which is associated with a decrease in thermodynamic entropy. Thus, the thermodynamic entropy remains invariant during RTO operation.
IX.5 Thermodynamics of algorithm
It is widely believed that the emergence of quantum computers will aid in solving longstanding problems in number theory Borevich and Shafarevich (1986); Hua (2012), combinatorial search Aigner (1988); Katona (1973), and even P and NP-classified problems Neukart (2023). To gain a deeper understanding of quantum speedups, it is essential to examine a realistic model of computation that takes into account factors such as time complexity Sipser (1996) and time-space tradeoffs. Several works Banegas and Bernstein (2018); Beals et al. (2013); Bernstein (2009); Fluhrer (2017) have delved into these directions. A recent study Perlner and Liu (2017) in this direction has investigated quantum speedups from a thermodynamic perspective. For the analysis of algorithmic cost (in both classical and quantum regimes) from a thermodynamic viewpoint, the Brownian model of computation is employed. In Perlner and Liu (2017), the authors consider the collision-finding algorithm and preimage search for analysis in their thermodynamic interpretation of algorithms.
The parallel collision search algorithm proposed by Van Oorshot and Wiener Van Oorschot and Wiener (1999) stands as the leading classical collision finding algorithm. This algorithm can detect a collision in an expected serial depth of $\mathcal{O}(\sqrt{G}/G_{\alpha})$, where $G$ denotes the range of the function and $G_{\alpha}$ depicts the parallel processes with memory $\mathcal{O}(1)$. Brassard, Hoyer, and Tapp (BHT) extended this algorithm in the quantum realm Brassard et al. (1998). The operations for this algorithm is $\mathcal{O}(G^{\frac{1}{3}})$ with memory size $G^{\frac{1}{3}}$.
Giovanetti et. al. in their recent study Giovannetti et al. (2008), address the memory cost by proposing a quantum random access memory (RAM) model, where the authors have conveyed that the memory access operation can be executed at logarithmic energy cost despite high gate complexity. So the question remains whether one can propose a realistic model where one encounters improvement in the complexity of the quantum algorithm over the classical one. The Brownian model is considered for the analysis, and it was inferred that the quantum algorithm has no advantage over the classical one.
Applying a similar analysis to that used for the collision-finding algorithm can be envisioned for Claw Finding problem Tani (2009); Belovs and Reichardt (2012); Jaques and Schanck (2019); Liu et al. (2023b); Brassard et al. (1998), where the objective of the Claw Finding problem is to find the collision between two functions with different domain sizes. The quantum version of this algorithm was investigated in Tani (2007), revealing that the energy cost for finding a collision is lower than the classical counterpart.
The on-the-go erasure protocol Meier and del Rio (2022) for the period finding algorithm has been addressed lately. Comparing Grover’s algorithm Grover (1996) with classical search algorithms rather than quantum versus classical collision search, reveals that Grover’s algorithm is more efficient in terms of query complexity than its classical counterpart but both regime exhibit similar asymptotic energy consumption (where powered and un-powered Brownian computation model is considered for the analysis). Grover’s algorithm plays a substantial role in cryptography Grassl et al. (2016); Lavor et al. (2003); Hsu (2003); Rahman and Paul (2021, 2020), and thereby prompts us to explore the thermodynamic analysis of cryptographic protocols.
X Conclusion and Future Direction
The analysis of the cost (like thermodynamic cost, operational cost) while executing a process has been one of the central points of attraction for physicists as well as computer scientists Aifer et al. (2023); Coles (2023). In physics, it is the thermodynamic cost of the process that plays a critical role, whereas in computer science, it is the computational cost. The analysis of the thermodynamic computational cost for a process was initiated from the physical Church-Turing thesis Kolchinsky (2023), where they conveyed that every computational process is physical. Various approaches have been considered for the analysis of the different computational processes. With the advent of modern statistical theory, the research in this area got boosted, yet it still remains a challenging task to explore thermodynamic computational cost (aka Landauer principle) for OQS, many-body physics, and at the quantum phase transition (QPT) Sachdev (1999); Vojta (2003); Osborne and Nielsen (2002); Heyl (2018); Sen et al. (2005); Prabhu et al. (2011) point. QPT takes place in a quantum many-body system Bandyopadhyay et al. (2021); Sur and Ghosh (2020); Fetter and Walecka (2012); Tasaki (2020); De Chiara and Sanpera (2018); Mukherjee et al. (2007); Bose (2003); Dutta (2015); Gómez et al. (1996); Ganahl et al. (2012); Fukuhara et al. (2013); Subrahmanyam (2004); Iyoda and Sagawa (2018); Tian et al. (2013) which happens at absolute zero temperature as a result of quantum fluctuations. The analysis of the minimum cost of heat dissipation of the erasure process in QPT is an open domain to explore.
In recent times, the applications of these principles are not just restricted to the field of physics or computer science but are also prevalent in other fields like chemical networks Chen et al. (2014); Soloveichik et al. (2008); Murphy et al. (2018); Qian and Winfree (2011), molecular biology Prohaska et al. (2010); Benenson (2012), and even in neurology Laughlin (2001); Balasubramanian et al. (2001).
To get a better understanding of the bond between thermodynamics and the computational process, further investigation in this direction is still required. For example, in the case of a finite automaton, one can investigate the maximum thermodynamic cost that is required to accept a language for automata. Also, one can calculate the minimal cost for any deterministic finite automaton. One can also work on developing a theory to analyze the non-deterministic finite automata in terms of thermodynamics. Models to describe the complex Turing machine, and also network theory from the thermodynamic viewpoint, are an open area of research.
It is known that different systems have different heat signatures. One can utilize this property of the system for various purposes, such as for security in cryptographic protocols. So one can explore the communications protocols and crypto-systems from a thermodynamic viewpoint. Algorithms in the form of a search algorithm from a thermodynamic viewpoint have already been analyzed. Further exploration in this direction is an open area of research. Thermodynamic analysis of quantum computations needs a rigorous investigation for a better understanding of quantum computers and to develop hardware with lower-cost functions.
Correction of errors during a computational process or communication is crucial. The analysis of the error correction protocols from a thermodynamic viewpoint is at its baby stage. Modeling of the EC models by the physical system to explain it thermodynamically needs further investigation. So the thermodynamic approach to explaining EC is an open book to read.
XI Acknowledgement
P.C., A.M. would like to thank Nilakanta Meher and Saikat Sur of the Weizmann Institute of Science for their valuable inputs and suggestions.
Appendix A Theorem of thermodynamic computation
Theorem 1: $E(\eta,\zeta)=K(\eta|\zeta)+K(\zeta|\eta)$ upto logarithmic term.
Proof 1: The upper and lower bounds for the thermodynamic cost are proposed.
Claim: $E(\eta,\zeta)\leq K(\eta|\zeta)+K(\zeta|\eta)+2[K(K(\zeta|\eta)|\zeta)+K(K(\eta|\zeta)|\eta)]$.
Proof: The computation is divided in three parts. In the first part of the program it computes $\zeta$ from $\eta$, and is depicted by $p_{c}$ and $|p_{c}|=K(\zeta|\eta)$. In the second part of the program $q_{c}$ computes $K(\eta|\zeta)$ from $\eta$ and $|q_{c}|=K(K(\eta|\zeta)|\eta)$, and in the final part of the program $r_{c}$ computes $K(\zeta|\eta)$ from $\zeta$, and $|r_{c}|=K(K(\zeta|\eta)|\zeta)$. Let’s now analyze the computation process step by step.
  • In the first step of the computation process $p_{c}$ computes $\zeta$ from $\eta$ and leaves behind garbage bits $g_{a}(\eta,\zeta)$.
  • $\zeta$ is copied, and then use one of its copies, along with the garbage bits, reverses the computation process to get $\eta$ and $p_{c}$.
  • Now $\eta$ is copied, and then use one of its copies along with $q_{c}$ to compute $K(\eta|\zeta)$ along with the garbage bits.
  • The shortest program is executed, which is depicted as $s_{c}$ to compute $\eta$ from $\zeta$ with the help of $\eta$, $\zeta$, $K(\eta,\zeta)$. In this process, extra garbage bits are produced.
  • Now, at this stage, $s_{c}$ is copied, and repeat the process shown in the third and fourth bullet. This helps to cancel the extra garbage bits. So we have $p_{c}$, $q_{c}$, $r_{c}$, $s_{c}$, $\eta$, $\zeta$.
  • $\zeta$ is copied again, and similarly, use one of its copies to compute $K(\zeta|\eta)$. It again results in garbage bits.
  • The shortest program is executed but for $p_{c}$ to compute $\zeta$ from $\eta$ with the help of $\eta$, $\zeta$, $K(\zeta,\eta)$. In this process, some extra garbage bits are produced.
  • Now, a copy of $p_{c}$ is deleted, and the process as shown in the sixth and seventh bullet is repeated. This helps to cancel the extra garbage bits. So we have $\zeta$, $r_{c}$, $s_{c}$, $q_{c}$.
  • $\eta$ is computed from $s_{c}$ and $\zeta$ and then reduce a copy of $\eta$ by canceling it. Now we are left with $\zeta$, $r_{c}$, $s_{c}$, $q_{c}$.
  • In the final step $s_{c}$, $r_{c}$, $q_{c}$ are erased.
$s_{c}$, $q_{c}$, and $r_{c}$ are thermodynamically erased in this computational process, leaving behind the output $\zeta$. This provides proof of the claim. Now the second claim, i.e., the upper bound of the thermodynamic cost of the computational process, where $\zeta$ is computed from $\eta$.
Claim: $E(\eta,\zeta)\geq K(\zeta|\eta)+K(\eta|\zeta)$.
Proof: The length of the shortest program to compute $\zeta$ from $\eta$ is defined as $K(\zeta|\eta)$. During the computation process, garbage bits $g_{c}(\eta,\zeta)$ are produced. By definition Zurek (1989b), the cardinality of the garbage bits is greater than or equal to the shortest program. So, to compute $\zeta$ from $\eta$, the garbage bits need to be erased, which is equivalent to at least $K(\eta|\zeta)$ bits. This proves the second claim.
So the claims prove Theorem 1.

### v2

@@
-Landauer Principle and Thermodynamics of Computation
+thanks: Contributed equallythanks: Contributed equally
 Abstract
@@
 Recently, there have been a few experimental investigations to reach this bound both in the classical and quantum domains.
-There has also been a spate of activities to enquire about this LB in finite time, with finite-size heat baths, non-Markovian and nonequilibrium environment in the quantum regime where the effects of fluctuations and correlation of the systems with the bath can no longer be ignored.
+There has also been a spate of activities to enquire about this LB in finite time, with finite-size heat baths, non-Markovian and nonequilibrium environments in the quantum regime where the effects of fluctuations and correlation of the systems with the bath can no longer be ignored.
 This article provides a comprehensive review of the recent progress on the Landauer bound, which serves as a fundamental principle in the thermodynamics of computation.
 We also provide a perspective for future endeavors in these directions.
-Furthermore, we review the recent exploration toward establishing energetic bounds of a computational process.
-We also review the thermodynamic aspects of error correction, which is an indispensable part of information processing and computations.
+Furthermore, we review the recent explorations toward establishing energetic bounds of a computational process.
+We also discuss the thermodynamic aspects of error correction, which is an indispensable part of information processing and computations.
 In doing so, we briefly discuss the basics of these fields to provide a complete picture.
@@
 $\bullet$ LL $\rightarrow$ Landauer Limit
-$\bullet$ LP $\rightarrow$ Landauer Principle
+$\bullet$ LP $\rightarrow$ Landauer’s Principle
 $\bullet$ LE $\rightarrow$ Landauer Erasure
@@
 I Introduction
-John von Neumann, in his 1948 lectures, posed a fundamental question regarding whether a computer operating at temperature $T$ must necessarily dissipate heat.
-The intricate details of this concept were later formalized in his book Theory of Self-Reproducing Automata, which was completed in 1966 by Arthur W.
+John von Neumann, in his 1948 lectures, posed a fundamental question on whether a computer operating at temperature $T$ must necessarily dissipate heat.
+The details of this lectures were later formalized in his book Theory of Self-Reproducing Automata, completed in 1966 by Arthur W.
 Burks Von Neumann et al.
 (1966).
-Building upon this idea, Landauer famously asserted that “real-world computation involves thermodynamic costs” and emphasized the substantial implications of this principle Landauer (1961, 1991).
-Further support for this notion came from Brillouin’s thought experiment Brillouin (1962), which validated the connection between computation and thermodynamics, albeit with some error probability.
+In 1961, Landauer asserted that “real-world computation involves thermodynamic costs” and emphasized the implications of the same Landauer (1961, 1991).
 The man-made computers and even all naturally occurring processes like biological computers have thermodynamic costs.
@@
 The connection between thermodynamics and a logically irreversible process gets most prominent in the context of Maxwell’s demon (MD) Maxwell (1871); Maruyama et al.
+(2009); Junior et al.
 (2025).
-Landauer (later Bennett) argued that one must pay an entropic cost that gets dumped in the environment as heat while performing a logically irreversible process that erases or throws away information.
+Landauer (later Bennett) argued that one must pay an entropic cost that gets dumped into the environment as heat while performing a logically irreversible process that erases or throws away information.
 This argument played the pivotal role in exorcising the MD and thereby saving the second law of thermodynamics.
 It also established that information is physical.
-There have been several explorations to model MD in physical systems and study the Landauer principle (LP).
+There have been several explorations to model MD in physical systems and study Landauer’s principle (LP).
 However, they mostly belong to equilibrium statistical physics.
-But when we approach to the miniaturized domain, the systems are, in general, highly non-equilibrium in their form.
+But when we approach to the miniaturized domain, the systems are, in general, out of equilibrium.
 The major breakthroughs in non-equilibrium statistical physics Zwanzig (2001); Kubo et al.
 (2012); Jarzynski (1997a, b) allow us to analyze the thermodynamic behavior of systems that are arbitrarily far from their equilibrium, and even for systems that undergo arbitrary external driving.
-With the advent of non-equilibrium statistical physics, the researchers have shown a keen interest to analyze the thermodynamic cost for the erasure process in this domain which includes finite time, finite size, open quantum systems, and so on Goold et al.
+With the advent of non-equilibrium statistical physics, the researchers have shown a keen interest to analyze the thermodynamic cost for the erasure process in this domain which includes finite time, finite size, non-Markovian (NM) open quantum dynamical effects, and so on Goold et al.
 (2015); Khoudiri et al.
@@
 (2013); Deffner and Jarzynski (2013); Barato and Seifert (2014); Strasberg et al.
+(2014); Ma et al.
+(2022); Mondal et al.
 (2023).
@@
 It can be quantified by counting the wrongly copied bits while execution.
-Though the error can be reduced at the macro level, at the molecular level, perfect copying is not achievable due to thermal fluctuations.
+Though the error can be reduced at the macro level, but at the molecular level, perfect copying is not achievable due to thermal fluctuations.
 It is the primary source of error at the molecular level.
 The replication process is limited by the thermal noise, so it must be interpreted in terms of thermodynamics as proposed by von Neumann Von Neumann (1956).
-This raises an important question: whether one can develop a connection between the errors that occur when copying remains.
+This raises an important question: Can one establish a fundamental connection between the thermodynamic cost and the errors incurred during the information copying process?
 Generally, while executing a copying process, it has to undergo various intermediate steps to control the accuracy and speed of the process.
 This is true for artificial as well as natural scenarios Johnson (1993).
-If we try to explain the errors from thermodynamic laws (second law), one must account for the fact that the copying process is repeated cyclically rather than occurring as a single-shot operation Bennett (1982).
-The works Sartori and Pigolotti (2015); Korepin and Terilla (2002) forge a profound link between thermodynamics and informational errors, demonstrating that errors arising in copying protocols are intrinsically connected to thermodynamic observables, which characterize the errors.Another aspect to connect thermodynamics and information theory is built on the basis of using entanglement in quantum systems Horodecki et al.
-(2001); Popescu and Rohrlich (1997); Rohrlich (2001).
-In this article, we review the recent developments in Landauer erasing in various contexts, including the most recent experimental demonstrations, thermodynamic aspects, and cost analysis of computation and error correction.
+If we try to explain the errors from thermodynamic laws (the second law), one must account for the fact that the copying process is repeated cyclically rather than occurring as a single-shot operation Bennett (1982).
+The works Sartori and Pigolotti (2015); Korepin and Terilla (2002) forge a profound link between thermodynamics and informational errors, demonstrating that errors arising in copying protocols are intrinsically connected to thermodynamic observables, which characterize the errors.
+Understanding the thermodynamic cost of computation is essential in both classical and quantum computation, as it reveals how the physical laws fundamentally constrain computational efficiency.
+Core models such as reversible computation, Turing machines, and finite automata offer deep theoretical insights into the thermodynamic cost of computational models.
+Reversible computing frameworks, in principle, allow for vanishingly low energy dissipation by preserving information throughout the process, yet practical limitations inevitably lead to non-negligible thermodynamic costs.
+Similarly, reversible Turing machines can asymptotically approach this ideal limit, though operations such as memory resetting and tape erasure introduce unavoidable energy losses.
+Finite automata, despite their relative simplicity, capture the essence of how logical state transitions and memory utilization contribute to entropy production.
+Collectively, these models provide a rigorous foundation for exploring the intrinsic trade-offs among energy consumption, computational speed, and information preservation in physically realizable computing systems.
+In this article, we review the recent developments on LP in various contexts, including the most recent experimental demonstrations, thermodynamic aspects, and cost analysis of computational models and error correction.
 Scope of the review
-Aiming to highlight recent advancements and fundamental insights into Landauer’s principle (LP) and the thermodynamics of computation, this review addresses several key aspects spanning both theoretical and experimental fronts.
+Aiming to highlight recent advancements and fundamental insights into LP and the thermodynamics of computation, this review addresses several key aspects spanning both theoretical and experimental fronts.
 We begin in Sec.
@@
 Sec.
-III reviews significant recent developments concerning LP, including scenarios involving finite time, finite-size systems, and non-Markovian reservoirs.
+III reviews significant recent developments on LP, including scenarios involving finite time dynamics, finite-size and non-Markovian reservoirs.
 In Sec.
-IV, we explore the role of Landauer’s limit (LL) in computational processes, followed by Sec.
+IV, we explore the role of Landauer limit (LL) in computational processes, followed by Sec.
 V, which presents an overview of recent experimental efforts aimed at realizing and validating LL in practical systems.
@@
 In Sec.
-VIII, we critically examine the thermodynamic consistency of error-correcting codes.
+VIII, we discuss the thermodynamic consistency of error-correcting codes.
 Prior to concluding and outlining potential future research directions in Sec.
-X, we briefly address a range of complementary topics under the umbrella of miscellaneous discussions in Sec.
-IX, such as viewing the computer as a heat engine, the thermodynamics of algorithms, and Landauer bound in switching protocols.
+X, we briefly address a range of complementary topics such as viewing the computer as a heat engine, the thermodynamics of algorithms, and the Landauer bound (LB) in switching protocols, under the umbrella of miscellaneous discussions in Sec.
+IX.
 Areas not covered in this review
-Though the review covers an extensive amount of works in the context of LP and thermodynamics of computation, there are some related topics and works, which are not discussed, that are worth mentioning.
-For exmaple, while most of the aspects of the LP have been reviewed here, some of the left-out aspects in this domain are like the role of LP in gravity Bormashenko (2019a); Haranas et al.
+Though the review covers an extensive amount of works in the context of LP and thermodynamics of computation, there are some related topics and work, that are not discussed here, but are worth mentioning.
+For example, while most of the aspects of the LP have been reviewed here, some of the left-out aspects in this domain are like the role of LP in gravity Bormashenko (2019a); Haranas et al.
 (2021); Daffertshofer and Plastino (2007), relativity Herrera (2020), quantum field theory Xu et al.
 (2022), many-body phenomena Bonança (2023); Parrondo (2001), material Zivieri (2022).
-In this review, we have considered mainly two fundamental computational models: Turing machine and Finite state machine, namely, for detailed thermodynamic analysis.
+In this review, we have considered mainly two fundamental computational models: the Turing machine and the Finite state machine, namely, for detailed thermodynamic analysis.
 Some aspects, like LB in analog computers and algorithmic thermodynamics, have not been discussed here.
 Interested readers can go through Diamantini et al.
-(2016) for the LB analysis in the analog system and Baez and Stay (2012) for algorithmic thermodynamics, which in turn allows one to apply the laws and techniques of thermodynamics for studying algorithmic information theory; in contrast, we have focused on the thermodynamic cost of algorithms in Sec.
+(2016) for the LB analysis in the analog system and Baez and Stay (2012) for algorithmic thermodynamics, which in turn allows one to apply the laws and techniques of thermodynamics for studying algorithmic information theory; however, we have touched upon the thermodynamic cost of algorithms briefly in Sec.
 IX.
@@
 Similarly, the thermodynamics of controlled systems is not covered.
-Please go through Touchette and Lloyd (2004, 2000); Barato and Seifert (2017); Sagawa and Ueda (2008, 2012); Wilming et al.
+Interested readers can go through Touchette and Lloyd (2004, 2000); Barato and Seifert (2017); Sagawa and Ueda (2008, 2012); Wilming et al.
 (2016); Large and Large (2021); Gingrich et al.
@@
 The search for the development of more impactful computing circuits leads to the question: What would be the physical limitation of this process?
-Rolf Landauer, in the year 1961, in his seminal work Landauer (1961) proposed an important limit to the conjecture surfaced by von Neumann Von Neumann et al.
-(1966), which is coined as “Landauer bound” (LB).
-This is a physical principle that provides the lower bound on the consumption of a computation process, i.e., an irreversible process in the computer dissipates a minimum amount of heat per bit Bennett (2003), which is expressed as
-$\Delta Q\geq k_{B}T\ln 2,$ (1)
-where $k_{B}$ is the Boltzman constant and $T$ is the temperature at which the computation process operates.
-The bound at room temperature is approximately 0.018 $eV$ (2.9 x $10^{-21}$ J), whereas the modern world computers use a million times more energy per operation Lambson et al.
+Rolf Landauer, in the year 1961, in his seminal work Landauer (1961) proposed an important limit to the posed question by von Neumann Von Neumann et al.
+(1966).
+Landauer stated that an irreversible process, where information is either deleted or thrown away, dissipates a minimum amount of energy per bit Bennett (2003), which is dumped as heat into the environment, increasing its entropy.
+This minimum amount of heat dissipated to erase one bit of information, known as “Landauer bound” (LB), is given by
+$Q\geq k_{B}T\ln 2,$ (1)
+where $k_{B}$ is the Boltzman constant and $Q$ is the heat dumped into the environment at temperature $T$.
+The bound at room temperature is approximately 0.018 $eV$ (2.9 x $10^{-21}$ J), whereas the modern world computers use a billion times more energy Lambson et al.
 (2011a); Moore (2012).
-Landauer Principle (LP) though widely accepted has faced various challenges and objections Earman and Norton (1998, 1999); Shenker (1998); Maroney (2005); Norton (2005, 2011); Bennett (2003).
-The prime objections to LP were put forth by Earman and Norton Earman and Norton (1999), where they have argued that LP is dependent on the second law of thermodynamics, it can be considered to be either unnecessary or insufficient for “exorcism of Maxwell’s Demon”.
-The other objections over LP are generally of three kinds Bennett (2003): (a) The principle fails, as thermodynamic quantities such as heat and work are fundamentally unrelated to mathematical constructs like logical reversibility.
-As a result, drawing direct parallels between them lacks any meaningful basis (as it was thought of to be a memory erasure process at that time).
+In 1867, Maxwell proposed a thought experiment, which he later elaborated on in his book “Theory of heat” Maxwell (1871), where a demon has a chamber of ideal gas in equilibrium at some fixed temperature.
+The demon creates a partition between the chamber and allows the molecules to cross the partition from, say, left to right if their velocity is greater than some threshold value.
+The molecules in the right can only go to the left if their velocities are less than the threshold value.
+Thus, the demon can create a temperature difference, which can further be exploited to extract work in principle, but without performing any work.
+This contradicts the second law of thermodynamics, as one can extract work from a single temperature source.
+Szilard proposed a simplified version of this thought experiment with a single gas molecule known as Szilard’s engine in 1929.
+This thought experiment puzzled the scientific communities for almost a century till the discovery of LP in 1961 Landauer (1961).
+The exorcism of MD can be done by virtue of the fact that the demon needs at least the amount of energy it has extracted earlier, on average, to erase their memory, before it can work in the next cycle.
+Brillouin’s argument to reconcile MD with the second law was that the demon must pay an entropic cost to acquire information, which he dubbed as negentropy Brillouin (1962).
+It was Bennet who showed in 1982 Bennett (1982) that the information gain and storage can be done reversibly in principle, and confirmed that only the erasure step must bear the entropic cost.
+Thus, the LP successfully exorcises the MD.
+Though widely accepted for reconciling MD with the second law Ladyman et al.
+(2007); Bennett (2003), Landauer’s Principle (LP) has faced various criticisms and objections Earman and Norton (1998, 1999); Shenker (1998); Maroney (2005); Norton (2005, 2011); Bennett (2003).
+The prime objections to LP were put forth by Earman and Norton Earman and Norton (1999).
+They argued that LP is dependent on the second law of thermodynamics; it can be considered to be either unnecessary or insufficient for “exorcism of Maxwell’s Demon”.
+In Sagawa and Ueda (2008); Cao and Feito (2009) it has also been argued recently that LP is a consequence of the second law.
+The other objections over LP are generally of three kinds Bennett (2003): (a) According to many, at that time at least, thermodynamic quantities such as heat and work are fundamentally unrelated to mathematical concepts like logical irreversibility.
+Therefore, drawing direct parallels between them seems unfounded.
 (b) In all cases of data-processing operations, there is a dissipation of at least $k_{B}T\ln 2$ amount of energy irrespective of the condition whether it is logically reversible or not.
-(c) Another reason for objection is that, in principle, logically irreversible operations can be engineered in a thermodynamically reversible way.
-Attempts have been made to demonstrate LP with precision.
-Piechocinska Piechocinska (2000) offered proof of LP with the help of statistical mechanics.
-The proof had a particular physical model assumption.
-The notable advancement was provided in the work Turgut (2009); Ladyman et al.
-(2007, 2008); Cao and Feito (2009); Leff and Rex (2002), where it provides a generalized version of the LP without a particular physical model assumption.
-The generalization of the LP surfaced in Vaccaro and Barnett (2011), where it has been shown that the information erasure will cause an increase in the entropy of the environment with no energy cost but the cost can be attributed to angular momentum of a spin-reservoir.
-In fact, angular momentum is exploited rather than energy from a spin-reservoir to erase the memory and it is consistent with the second law of thermodynamics.
-In the low-temperature limit, quantum effects play an important role in the analysis of the thermodynamic aspect of the erasure principle.
-The validity of the erasure principle in this limit has been challenged Allahverdyan and Nieuwenhuizen (2000); Nieuwenhuizen and Allahverdyan (2002); Hörhammer and Büttner (2005); Hilt and Lutz (2009).
-It has been claimed that due to the presence of entanglement between the system and the environment LP is broken, which implies information is erased while heat is absorbed Allahverdyan and Nieuwenhuizen (2001); Hörhammer and Büttner (2008); Cápek and Sheehan (2005); Maruyama et al.
-(2009) which counters the LP.
-Lutz et al.
+In a realistic physical scenario, it is even much more.
+(c) Another reason for objection is that, in principle, even logically irreversible operations can be engineered in a thermodynamically reversible way.
+In Bennett (2003), the author discussed in detail and refuted all these criticisms.
+Attempts have been made to establish LP rigorously.
+Piechocinska Piechocinska (2000) offered a proof of LP with the help of statistical mechanics.
+The proof assumes a particular physical model.
+The notable advancements have been done in Turgut (2009); Ladyman et al.
+(2007, 2008); Cao and Feito (2009); Leff and Rex (2002), where LP has been generalized without assuming a particular physical model.
+A notable generalization of the LP surfaced in Vaccaro and Barnett (2011), where it has been shown that the information erasure will cause an increase in the entropy of the environment with no energy cost, but the cost can be attributed to the angular momentum of a spin-reservoir.
+In fact, angular momentum is exploited rather than energy from a spin-reservoir to erase the memory, and it is consistent with the second law of thermodynamics.
+With rapid technological advancements and our ever increasing grasp on the miniaturized scale, probing and controlling the small scale systems that exhibits quantum traits are becoming more and more possible.
+In this regime, where we can no longer ignore the fluctuations, quantum correlations between the system and the environment, the quantum effects bring fundamental challenges in the analysis of memory erasure.
+The validity of the erasure principle in this limit has been questioned Allahverdyan and Nieuwenhuizen (2000); Nieuwenhuizen and Allahverdyan (2002); Hörhammer and Büttner (2005); Hilt and Lutz (2009).
+It has been claimed that due to the presence of entanglement between the system and the environment LP is violated, which implies information is erased while heat is absorbed Allahverdyan and Nieuwenhuizen (2001); Hörhammer and Büttner (2008); Cápek and Sheehan (2005); Maruyama et al.
+(2009).
 Hilt et al.
-(2011), tackled this quantum conundrum and demonstrated that LP remains valid regardless of the specific nature of the interaction between the system and its environment.
-Their findings provided significant evidence supporting the applicability of LP in quantum settings.
-To address the need for a theory with minimal assumptions in both regimes, approaches are considered that derive the LP without relying explicitly on the second law of thermodynamics Shizume (1995); Piechocinska (2000); Sagawa and Ueda (2009, 2011).
-Reeb’s work Reeb and Wolf (2014) is on this line of thought.
-The assumptions that are considered for the analysis are minimal based on the benchmark works.
-The assumptions are: (a) In the process, the system and the reservoir belong to a Hilbert space, (b) the reservoir is initially in a thermal state, (c) the system and the reservoir will be initially uncorrelated, and (d) the evolution process is unitary.
-With this minimum condition, a sharpened equality version of LP has been derived for a system $\mathcal{S}$ and a reservoir $R$ as
-$\beta\Delta Q=\Delta S+I(\mathcal{S}^{\prime}:R^{\prime})+S(\rho_{R}^{\prime}||\rho_{R})\geq\Delta S.$ (2)
-Here $\prime$ denotes the final states, $\Delta S$ is the change of von Neumann entropy of the system, $I(\mathcal{S}^{\prime}:R^{\prime})$ describes the mutual information and quantifies the correlation of the system with the bath, and $\rho_{R}$ denotes the reservoir state.
+Hilt et al.
+(2011), have tackled this quantum conundrum and demonstrated that LP remains valid regardless of the specific nature of the interaction between the system and its environment.
+Their findings have provided significant evidence supporting the applicability of LP in the quantum settings.
+To address the need for a theory with minimal assumptions, approaches have been considered that derive the LP without relying explicitly on the second law of thermodynamics in both classical and quantum regimes Shizume (1995); Piechocinska (2000); Sagawa and Ueda (2009, 2011).
+Reeb and Wolf’s work Reeb and Wolf (2014) is the most significant one along this direction in quantum setups.
+They have considered the following minimal set of assumptions: (a) In the process, the system and the reservoir belong to corresponding Hilbert spaces.
+(b) The reservoir is initially in a thermal state.
+(c) The system and the reservoir are initially uncorrelated.
+(d) The evolution process is unitary on the joint system-resrvoir state.
+With this minimal setup, a sharpened equality version of LP has been derived for a system $\mathcal{S}$ and a reservoir $R$, for the dissipated heat $Q$ to the environment as
+$\beta Q=\Delta S+I(\mathcal{S}^{\prime}:R^{\prime})+S(\rho_{R}^{\prime}||\rho_{R})\geq\Delta S.$ (2)
+Here $\prime$ denotes the final states, $\Delta S$ is the decrease in von Neumann entropy of the system, $I(\mathcal{S}^{\prime}:R^{\prime})$ describes the mutual information which quantifies the correlation of the system with the bath Nielsen and Chuang (2002), and $\rho_{R}$ denotes the reservoir state.
 $\beta=1/k_{B}T$ denotes the inverse temperature of the bath.
-$S(\rho_{R}^{\prime}||\rho_{R})$ quantifies the free energy increase in the bath, where $S(\rho||\rho^{\prime})=\text{Tr}\,\rho\,\ln\,\rho-\text{Tr}\,\rho\,\ln\,\rho^{\prime}$ is the quantum relative entropy between $\rho,\,\rho^{\prime}$.
-The setup mentioned above can be generalized by considering an initial correlation in the process.
-So the assumption gets modified as: (a) In the process, the system, reservoir, and memory are in a joint quantum state initially, (b) the reservoir is initially in a thermal state, (c) the heat exchange and the entropy of the system are expounded on the marginal states, and (d) the evolution is processed by an unital positive trace-preserving map.
+$S(\rho_{R}^{\prime}||\rho_{R})$ quantifies the free energy increase in the bath, where $S(\rho||\rho^{\prime})=\text{Tr}\,[\rho\,\ln\,\rho]-\text{Tr}\,[\rho\,\ln\,\rho^{\prime}]$ is the quantum relative entropy between $\rho,\,\rho^{\prime}$ Nielsen and Chuang (2002).
+They further generalized the setup Reeb and Wolf (2014) by considering an initial correlation in the process, which modifies the assumption as follows: (a) In the process, the system, reservoir, and memory are in a joint quantum state initially, (b) the marginal state of the reservoir is initially thermal, (c) the heat and entropy exchange are expounded on the marginal states, and (d) the evolution is processed by an unital positive trace-preserving map.
 In this generalized setup, where the system, reservoir, and memory are initially correlated, the standard Landauer bound can be modified significantly.
@@
 When the evolution is governed by a unital, trace-preserving map, which does not reduce entropy on its own, the role of initial correlations becomes even more prominent.
-In the present-day protocols where the LP plays a vital role, Reeb’s version of the erasure process is considered for the analysis.
-Even with the minimal assumption of Reeb, it is clear that one requires dissipative dynamics to explore the LP.
-However, it is a difficult task in pure Hamiltonian dynamics where the evolution is for the system only.
+Even with the minimal assumption of Reeb and Wolf, it is considered that the system is governed by non-unitary dynamics.
+However, it is a difficult task to explore erasing in pure Hamiltonian dynamics for the system.
 It has been conveyed in Holtzman et al.
-(2021) that it is possible to implement an erasure bit with no thermodynamic cost using Hamiltonian dynamics if one has the information of the system with infinite accuracy.
+(2021) that it is possible to implement a bit-erasure with no thermodynamic cost using Hamiltonian dynamics if one has the information of the system with infinite accuracy.
 In this case, it corresponds to the energy of the system being known with infinite accuracy.
+This requires diverging computational resources as we will discuss later in Sec.
+III B.
 So far, we have discussed LB for a non-zero finite temperature $T$ of the environment.
@@
 (2020).
-There, the authors have assumed that initially the environment is in a thermal state, as in the original case of LEP.
+There, the authors have assumed that initially the environment is in a thermal state, as in the original case of LP.
 The state of the system or the type of system-environment interaction is considered to be general.
@@
 III Landauer Principle considering the open-system dynamics
-The study of the dynamics of a quantum system that interacts with other system or systems (environment) is described as an open quantum dynamics (OQD) Breuer et al.
+The study of the dynamics of a quantum system that interacts with other system(s) (environment) is described as an open quantum dynamics Breuer et al.
 (2002); Rivas and Huelga (2012); Rotter and Bird (2015); Viola et al.
@@
 (2002).
-Till now, we have discussed the LB considering the entropic bounds sans the dynamics during erasing.
+Till now, we have discussed the LB considering the entropic bounds sans the dynamic detail or fluctuations during the erasing processes.
 The dynamics of erasing bear significant importance in practical scenarios.
-For example, the LB is reached in infinite time.
+For example, LB is reached in infinite time.
 But the practical uses demand finite-time erasure.
-Also, during the erasing procedure, the environment can get far from its equilibrium (due to finite size), and the fluctuations due to it are crucial for LB.
+Also, during the erasing procedure, the environment can get far from equilibrium (due to finite size), and the fluctuations due to it are crucial for LB.
 What do the NM features add to the LP?
@@
 III.1 Finite Time Landauer’s Principle
-The heat dissipated (LL) in the environment to erase one bit of information is achieved under the assumption of a quasistatic transformation.
-In practical scenarios, the information erasure occurs in finite time.
-Therefore, finite-time analysis Andresen (2011) bears much importance both in the classical and quantum regimes.
+The heat dissipated in the environment to erase one bit of information is achieved under the assumption of a quasistatic transformation.
+In practical scenarios, the information erasure occurs in a finite time.
+Therefore, the finite-time analysis Andresen (2011) bears much importance both in the classical and quantum regimes.
 There has been an upsurge of research on finite-time erasure of bits both in the quantum and stochastic thermodynamics Seifert (2012); Van den Broeck and Esposito (2015).
@@
 Building on this approach, the authors in their work Proesmans et al.
-(2020a) propose a framework that also provides full control over the potential landscape but removes the constraint that the final state must be a best-fitted state.
+(2020a) proposed a framework that enables full control over the system’s potential without requiring a pre-specified final state, allowing more flexible and efficient erasure protocols.
+Using this, they derive tight lower and upper bounds on the dissipated work during optimal finite-time bit erasure.
 Extensions Proesmans et al.
-(2020b) and alternative approaches Zhen et al.
+(2020b) in the direction of Proesmans et al.
+(2020a) and alternative approaches Zhen et al.
 (2021); Dago et al.
@@
 Additionally, the driving of the control Hamiltonian is performed slowly in the time interval $t\in[0,\tau]$, relative to the relaxation timescale of the dynamics which implies that the system remains close to the equilibrium state at all times, i.e., $\rho_{t}=\pi_{t}+\tau^{-1}\delta\rho_{t}$ where corrections of higher orders $O(\tau^{-n})$ can be neglected.
-In this quasistatic limit, the dissipation approaches the LE bound.
+In this quasistatic limit, the dissipation approaches the LB.
 The LP is performed by taking the initial Hamiltonian $H_{0}\simeq 0$ and then slowly increasing the energy gap of the Hamiltonian until it reaches far beyond $k_{B}T$.
@@
 $\dot{\theta}_{t}=0$ describes the classical bit whereas $\dot{\theta}_{t}\neq 0$ describes the non-commuting quantum double-well case.
-Building on the insights of Miller et al.
-(2020), recent work Taranto et al.
-(2023) explored the thermodynamic limits of quantum cooling, juxtaposing LP with Nernst’s unattainability principle.
-This study introduced a Carnot-Landauer limit, demonstrating that perfect cooling—achieving absolute zero temperature—demands either infinite energy or unbounded control complexity.
-The cooling bound is given by:
+It is important to mention that perfectly cooling a system to its desired final state and perfect erasing of information are fundamentally equivalent Taranto et al.
+(2023), as both correspond to reducing the rank of the system to 111.
+However, Nearst’s unattainability principle, often referred to as the third law of thermodynamics, states that to perfectly cool a physical system, one needs diverging resources, either in terms of energy or time.
+The authors of Taranto et al.
+(2023) showed that it is possible to cool perfectly with finite energy and time provided one is equipped with coherent control whose complexity diverges.
+In essence, under coherent control, perfect cooling is possible if either time, energy, or control complexity diverges.
+They further showed that with incoherent control (or cooling aided by a heat engine), perfect cooling remains fundamentally unattainable within finite time and finite control complexity, regardless of the amount of energy extracted from the auxiliary engine.
+Consequently, they established that for cooling under incoherent control with finite control complexity, it is possible to cool a system arbitrarily close to a desired temperature within accuracy $\epsilon_{\beta}>0$, with infinite number of operations and thereby in infinite time, with a energy, as measured by the amount of heat drawn from a hot bath at inverse temperature $\beta_{h}$, which is arbitrarily close to the bound called Carnot-Landauer limit, that is
 $\left|\Delta E_{H}-\frac{1}{\eta_{c}}\Delta F_{S}^{(\beta)}\right|<\epsilon_{E},$ (5)
-where $\eta_{c}=1-\frac{\beta_{H}}{\beta}$ is the Carnot efficiency, $\beta_{h}$ and $\beta$ being the temperatures of the two thermal baths, $\Delta E_{H}$ quantifies the energy transferred to the subsystem from the hot bath, and $\Delta F_{S}^{(\beta)}$ represents the free energy difference between the initial and final states of the system, and $\epsilon_{E}\geq 0$ denotes the error in the cooling protocol.
-The role of coherence in finite-time processes becomes even more pronounced when considering many-body effects and collective erasure.
+where $\eta_{c}=1-\frac{\beta_{H}}{\beta}$ is the Carnot efficiency, $\Delta E_{H}$ quantifies the energy drawn from the hot bath, and $\Delta F_{S}^{(\beta)}$ represents the free energy difference between the initial and final states of the system at inverse temperature $\beta$, and $\epsilon_{E}>0$.
+The Carnot-Landauer bound $\frac{1}{\eta_{c}}\Delta F_{S}^{(\beta)}$ reduces to the LB when $\beta_{H}=0$.
+Both these bounds can be reached asymptotically with diverging control complexity or time.
+The role of quantum features like coherence in finite-time processes becomes even more pronounced when considering many-body effects and collective erasure.
 The collective effects of many-body systems in finite-time erasure Rolandi et al.
@@
 Unlike independent qubit erasure, where excess work scales linearly with system size, many-body protocols exhibit sublinear scaling: $W_{\text{diss}}\propto N^{x},\quad x<1$, indicating an accelerated convergence to LB and an enhancement in the efficiency of information erasure.
-Building on Miller’s framework, Van Vu and Saito Van Vu and Saito (2022) further examined the interplay between coherence and erasure speed, establishing that finite-time erasure introduces an additional distance cost (distance error to the ground state) beyond the LL.
+In this direction, Van Vu and Saito Van Vu and Saito (2022) further examined the interplay between coherence and erasure speed, establishing that finite-time erasure introduces additional cost dependent on the distance error to the desired final state and inversely proportional to the total time taken by the process, beyond the standard LB.
 Their findings reinforced the idea that quantum coherence consistently amplifies dissipation, regardless of the control protocol or driving speed.
@@
 These algorithms would encode a complete understanding of the initial conditions and the exact transformations required, bypassing the usual thermodynamic cost that arises from ignorance or randomness in the input.
-However, this does not violate Landauer’s principle; rather, it circumvents the physical cost by assuming unbounded computational resources, shifting the “cost” from thermodynamic to computational complexity  Zurek (1989a).
+However, this does not violate LP; rather, it circumvents the physical cost by assuming unbounded computational resources, shifting the “cost” from thermodynamic to computational complexity  Zurek (1989a).
 In practice, limitations arise due to thermodynamic speed limits Deffner and Campbell (2017) and coherence-induced excess dissipation Faist et al.
 (2015), finite computational resources.
-Recent studies Van Vu and Saito (2022) further quantify the finite-time cost, revealing that even with optimal control, a residual energy cost remains due to quantum fluctuations and irreversibility.
-Thus, while the Landauer limit can be theoretically approached, the interplay of coherence, control complexity, and finite-time constraints fundamentally prevents its exact realization in realistic settings.
-III.2 Non-equilibrium Landauer Process
-The miniaturization in modern technology has led to the development of small systems that are out-of-equilibrium in classical Jarzynski (2011); Seifert (2012) as well as in the quantum regime Esposito et al.
+While the LB can be theoretically approached, the interplay of coherence, control complexity, and finite-time constraints fundamentally prevents its exact realization in realistic settings.
+III.2 Non-equilibrium Landauer Erasure
+The miniaturization in modern technology has led to the control of small systems that are out-of-equilibrium in classical Jarzynski (2011); Seifert (2012) as well as in the quantum regime Esposito et al.
 (2009); Campisi et al.
@@
 The fluctuation relation Jarzynski (2011, 1997c, 2004); Crooks (1999); Tasaki (2000); Kurchan (2000); Mukamel (2003); Campisi et al.
-(2011b) plays a promising role in understanding the thermodynamics of these small systems that operate under the non-equilibrium condition where the thermal and quantum fluctuations cannot be neglected.
-In this review article, we specifically focus on nonequilibrium erasure protocols within the quantum regime.
+(2011b) plays the central role in understanding the thermodynamics of these small systems that are far from the equilibrium.
+Here, we mainly focus on nonequilibrium erasure protocols within the quantum regime.
 A key challenge in non-equilibrium thermodynamics is characterizing the dissipation associated with quantum erasure.
-Reeb and Wolf Reeb and Wolf (2014) established a fundamental lower bound for heat dissipation in equilibrium conditions, highlighting the role of quantum coherence in modifying LP.
-However, real-world erasure processes often operate in non-equilibrium regimes, where such equilibrium-based bounds may not be directly applicable.
-To bridge this gap, their framework has been recast from a fluctuation relation perspective, which provides a generalized thermodynamic bound for erasure beyond equilibrium.
+Reeb and Wolf Reeb and Wolf (2014) have established a fundamental lower bound for heat dissipation, using the tools of quantum information theory and quantum statistical mechanics, that includes corrections arising from finite-size reservoir and system-bath correlation.
+Though Reeb and Wolf have not incorporated the fluctuations arising due to the finite-size effects in their derivation, one might be tempted to think that the correction due to the finite-size effects is somewhat related to the nonequilibrium fluctuations, as we are far away from the thermodynamic limit.
+Indeed, an equivalent form Eq.
+(2), has been derived independently in Esposito et al.
+(2010) to express irreversible entropy production of a finite system out of equilibrium.
+This indicates the necessity of LB-like bound in far from equilibrium conditions, considering the fluctuations explicitly in the setup.
+To bridge this gap, researchers have sought to establish generalized bounds for information erasure beyond equilibrium, employing the concepts of fluctuation relations Jarzynski (2011, 1997c, 2004).
 In Goold et al.
-(2015), an erasure protocol that involves a finite-size environment that interacts with the system is proposed.
-The system is a single qubit system that is coupled to a finite-dimensional environment.
-A tantalizing fact is that the nonunitality of the open system dynamics leads to a tighter bound for the heat dissipation for the erasure process.
-This opened the door for the analysis of the cost of computation in non-equilibrium circumstances.
-Following the same methodology, a comparative analysis of the performance of LB with that of the bound proposed in Goold et al.
-(2015) for the non-equilibrium condition is addressed Campbell et al.
-(2017); Zhang et al.
+(2015), an erasure protocol that involves a finite-size environment is studied, and a modified bound considering the fluctuations during the protocol is proposed.
+This has opened the door for the analysis of the cost of computation in non-equilibrium circumstances.
+Under the same four assumptions that have been used in Reeb and Wolf (2014) to derive Eq.
+(2), the authors of Goold et al.
+(2015) have first derived the following heat fluctuation-like relation of the environment for the erasure protocol
+[…]
+Here $P(Q)$ denotes the heat distribution of the environment, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{l=jk}}=\sqrt{\lambda_{j}}\langle s_{k}|U|s_{j}\rangle$, with $\lambda_{j}$ being the eigenvalue and $|s_{j}\rangle$ being the eigenstate of $\rho_{\mathcal{S}}$, $U$ is the joint unitary operator acting on the system and environment and $\textbf{A}=\sum_{l}A_{l}A^{\dagger}_{l}$.
+The above equation is in the same spirit of Jarzynski equality which connects the average exponentiated work to the equilibrium free energy exploiting the work distribution Jarzynski (1997c).
+If the operator A is expanded in terms of the initial state of both $\mathcal{S}$ and $\mathbf{E}$ under the action of the global unitary evolution, the heat fluctuation relation becomes
+$\langle e^{-\beta Q}\rangle=\text{tr}[\textbf{M}\rho_{\mathcal{S}}],$ (7)
+where $\textbf{M}=\text{tr}_{\mathbf{E}}[U^{\dagger}\mathbf{I}_{\mathcal{S}}\otimes\rho_{\mathbf{E}}U$].
+The desired heat dissipation during the erasure process satisfies111 In this nonequilibrium LP section, we denote the averaged heat inside bra-ket to distinguish it from the heat exchange in a single trajectory.
+In Eq.
+(1), Eq.
+(2), and all other sections of the review, $Q$ (“heat”) denotes heat dissipated on average.
+$\beta\langle Q\rangle\geq\mathbf{B}_{Q},$ (8)
+where $\mathbf{B}_{Q}=-\,\ln(\text{tr}[\textbf{A}\rho_{\mathbf{E}}])=-\,\ln(\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}])$, termed as the thermodynamic bound in Campbell et al.
 (2017).
-Reeb and Wolf Reeb and Wolf (2014) in their work provided the minimum framework that is required for the Landauer process in the equilibrium condition.
-The minimum framework for the execution of the Landauer process in the non-equilibrium condition has been addressed in Taranto et al.
-(2018), where the system and the environment are equivalent to the models in Campbell et al.
-(2017); Zhang et al.
-(2023).
-Although this approach offers valuable insights, it remains limited to particular models and does not yet provide a universally applicable methodology.
-To rigorously understand LP in the context of quantum non-equilibrium dynamics, further investigation is essential to develop a minimal yet broadly generalizable framework that captures the complexities unique to quantum systems.
-The prerequisites for the erasure protocol are (i) A system $\mathcal{S}$ with $H_{\mathcal{S}}$ as the free Hamiltonian of the system.
-(ii) An environment $\mathbf{E}$ initially uncorrelated with the system, i.e., $\rho_{\mathcal{S}\mathbf{E}}=\rho_{\mathcal{S}}\otimes\rho_{\mathbf{E}}$.
-(iii) The initial state of the environment will be the Gibbs state.
-(iv) The interaction of the environment and the system is unitary.
-The heat fluctuation relation of the environment for the erasure protocol is:
-[…]
-which follows from the equality proposed by Jarzynski Jarzynski (1997c) using the work distribution.
-$P(Q)$ denotes the heat distribution of the environment.
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}_{l=jk}}=\sqrt{\lambda_{j}}\langle s_{k}|\hat{U}|s_{j}\rangle$, with $\lambda_{j}$ being the eigenvalue and $|s_{j}\rangle$ being the eigenstate of $\rho_{\mathcal{S}}$.
-The process is unital iff $\hat{\textbf{A}}=\mathbf{I}_{\mathbf{E}}$.
-If the operator $\hat{\textbf{A}}$ is expanded in terms of the initial state of both $\mathcal{S}$ and $\mathbf{E}$ under the action of the global unitary evolution, the heat fluctuation relation becomes
-$\langle e^{-\beta Q}\rangle=\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}],$ (7)
-where $\hat{\textbf{M}}=\text{tr}_{\mathbf{E}}[\hat{U}^{\dagger}\mathbf{I}_{\mathcal{S}}\otimes\rho_{\mathbf{E}}\hat{U}$]. The desired heat dissipation during the erasure process is
-$\beta\langle Q\rangle\geq\mathbf{B}_{Q},$ (8)
-where $\mathbf{B}_{Q}=-\,\ln(\text{tr}[\hat{\textbf{A}}\rho_{\mathbf{E}}])=-\,\ln(\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}])$ is termed the thermodynamic bound.
-The operator $\hat{\textbf{A}}$ is dependent on the choice of the state of the system.
-So, for the execution of the erasure process of a state of choice, one needs to compute $\hat{\textbf{A}}$ for every instance.
-On the other hand, $\hat{\textbf{M}}$ as defined in Eq.
-(7) can be easily evaluated just by implementing the unitary interaction.
-This proposed bound is tested on a physical system where the system is considered to be a single qubit and the environment comprises an interacting spin chain (as shown in Fig.
+The process is unital iff $\textbf{A}=\mathbf{I}_{\mathbf{E}}$, and therefore the bound can be insightful to compare the heat fluctuations of the environment under unital and nonunital channels.
+The operator A is dependent on the choice of the state of the system.
+So, for the execution of the erasure process of a state of choice, one needs to compute A for every instance.
+On the other hand, M as defined in Eq.
+(7) can be evaluated just by implementing the unitary interaction.
+This proposed bound is tested Goold et al.
+(2015) on a physical system where the system is considered to be a single qubit and the environment comprises an interacting spin chain (as shown in Fig.
 1).
@@
 $\mathbf{B}_{Q}$ $=$ $-\ln\left[2(1-\phi^{2})(\alpha^{2}+p+2p\alpha^{2})+\phi^{2}\right].$ (10)
-Here $\phi=\cos(2Jt)$ at time $t$, $p=[1+\tanh(\beta B)]/2$ and $\alpha\in\mathbb{R}$.
+Here $\phi=\cos(2Jt)$ at time $t$, $p=[1+\tanh(\beta B)]/2$ and $\alpha(\in\mathbb{R})$ is an initial state parameter of the system.
 For $\alpha\approx 1$ and for a particular temperature it is observed that $\mathbf{B}_{Q}\geq\mathbf{B}_{rw}$, where $\mathbf{B}_{rw}$ denotes the bound proposed in Reeb and Wolf (2014).
+It is important to mention that the tightness of the bound $\mathbf{B}_{Q}$ over $\mathbf{B}_{rw}$ is parameter dependent.
+There can be regions in the parameter space where the bound $\mathbf{B}_{rw}$ is tighter than the thermodynamic one.
 Figure 1: The schematic of the process where the system $\mathcal{S}$ interacts with the environment $\mathbf{E}$.
-A comparison of different forms of the LBs reveals a subtle dependence on the initial state of the system and the environment’s temperature.
-In the parameter space, sharp boundaries emerged when assessing the relative effectiveness of these bounds Campbell et al.
+In Campbell et al.
+(2017), the authors have performed a detailed comparative study of the thermodynamic bound $\textbf{B}_{Q}$ and the bound $\beta\langle Q\rangle\geq\Delta S$ (see Eq.
+(2)), dubbed as the entropic bound.
+They have explored the role of the initial state parameters on the tightness of the two bounds with respect to the dissipated heat.
+It has also been shown that the thermodynamic bound is not sensitive to the initial state coherence, in contrast to the entropic bound.
+Moreover, they have shown that there are regions where both bounds are ineffective, being less informative than the Clausius statement of the second law.
+In Taranto et al.
+(2018), the authors have reported a very important aspect of heat fluctuations in erasure protocol.
+Building upon the framework of Goold et al.
+(2015), they have shown using generalized concentration of measure that fluctuations of heat from its mean are exponentially damped with the size of the larger subsystem (system or environment), and this suppression scales linearly with increasing inverse temperature of the environment, regardless of the system’s microscopic dynamics.
+Though Goold et al.
+(2015) has paved the way to incorporate the fluctuations-like relations of heat to get an improved bound on the dissipated heat, there remains plenty of scope for further explorations.
+For example, the quest for a bound that considers fluctuations of heat, which is more informative than the Clausius inequality in general, remains.
+Generalization of these bounds to an initially correlated system-environment would be challenging and intriguing.
+The full counting statistic Esposito et al.
+(2009) formalism has been applied to analyze the bounds (lower and upper) on the average heat dissipation in an erasure process in Guarnieri et al.
 (2017).
-There were scenarios where the bounds were negative in both frameworks, yet the process’s heat dissipation remained positive.
-In such cases, the bounds proved to be weaker than the Clausius statement of the second law.
-The full counting statistic Esposito et al.
-(2009) formalism, when applied to analyze the bounds (lower and upper) on the average heat dissipation in an erasure process, a single-parameter bound is developed that can be made arbitrarily tight and is independent of the map being used for the execution of the process Guarnieri et al.
+A single-parameter family of bounds is developed that can be made arbitrarily tight and is independent of the underlying map used for the execution of the process Guarnieri et al.
 (2017).
-This inherently marks the difference of the lower bound of this process over the previous bound Goold et al.
+This inherently marks the difference of the lower bound of the nonequilibrium erasure process over the previous bound Goold et al.
 (2015).
-For this formalism, the minimal set of assumptions is considered as follows in Reeb and Wolf (2014), which validates LP.
-The full counting statistics of heat dissipation are defined as the change of energy in the environment that characterizes the mean value of heat dissipation.
-The lower bound for the heat dissipation using the CGF Rockafellar (1970) is
+For this formalism, the minimal set of assumptions is considered as in Reeb and Wolf (2014).
+In this full counting statistics of dissipated heat formalism, the lower bounds for the heat dissipation can be given using the (CGF) Rockafellar (1970) as
 $\beta\langle Q\rangle_{t}\geq-\frac{\beta}{\eta}\Theta(\eta,\beta,t)\equiv\mathbf{B}_{Q}^{\eta}(t)\quad(\eta>0),$ (11)
-with $\eta$ being the counting parameter.
-Using the large deviation function (useful for studying dynamical phase transitions), which is a powerful tool to study statistical properties for a long time scale Garrahan and Lesanovsky (2010); Lesanovsky et al.
+$\eta$ being the counting parameter. The Landauer-like bounds here are proposed with one parameter on a two-time measurement protocol. For $\eta=\beta$ the result of Goold et al. (2015) is recovered.
+Furthermore, using the large deviation function, which is a powerful tool to study statistical properties for a long time scale Garrahan and Lesanovsky (2010); Lesanovsky et al.
 (2013); Pigeon et al.
-(2015), the upper bound for heat dissipation is proposed as
+(2015), the upper bounds of heat dissipation are given as
 $\beta\langle Q\rangle_{t}\leq\frac{\beta}{|\eta|}\Theta(\eta,\beta,t)\equiv\mathbf{\tilde{B}}_{Q}^{\eta}(t)\quad(\eta<0).$ (12)
-The Landauer-like bound here is proposed with one parameter on a two-time measurement protocol.
-For $\eta=\beta$ the derived form of the bound is equivalent to that of the bound proposed in Goold et al.
-(2015).
-This bound is tested on a physical system where the system is considered to be a three-level V-system and the environment is modeled by a two-level system (as in Fig.
+These bounds are tested on a physical system where the system is considered to be a three-level V-system and the environment is modeled by a two-level system (as in Fig.
 2).
@@
 Figure 2: The schematic of the process with a three-level system and a two-level environment.
-This physical model highlights the tightness of the proposed bound for heat dissipation during the execution of the erasure protocol.
-III.3 Landauer Bound in Non-Markovian Process
-The Markovian approximation is the most convenient way to express the dynamics of OQS.
-In this process, the evolution timescale is considered to be larger than the correlation time of the environment.
+This physical model highlights the tightness of the proposed bounds.
+In Zhang et al.
+(2023), the authors have studied thermodynamics of nonequilibrium reservoirs with the help of Landauer-like bound.
+They have shown that a bound in terms of equality, called the Landauer-like bound, always holds even for nonequilibrium reservoirs, but an inequality obtained from the equality by excluding a few contributions might be violated.
+Violation of this inequality, dubbed as Landauer-like inequality, indicates the possibility of spontaneous heat transfer from a cold bath to a hot bath, beating the Carnot bound and extracting work from a single nonequilibrium reservoir.
+III.3 Landauer Bound under Non-Markovian Dynamics
+The Markovian approximation is the most convenient way to express the dynamics of OQS Rau (1963); Alicki and Lendi (2007); Scarani et al.
+(2002); Ziman and Bužek (2005); Gennaro et al.
+(2009).
+In this process, the evolution timescale of the system is considered to be larger than the correlation time of the environment.
 In other words, in this approximation, the memory effect (or information backflow) is neglected.
-However, it plays a non-minuscule role in the analysis of the dynamics of the system and thermodynamic cost.
-There comes the role of non-Markovian (NM) dynamics to explore the dynamics of the system De Vega and Alonso (2017); Breuer et al.
-(2016).
-Various methods and approaches have surfaced for defining as well as quantifying non-Makovianity Breuer et al.
+However, it plays a non-minuscule role in the analysis of the dynamics of the system and thermodynamic cost of erasure.
+This backflow of information or non-Markovian (NM) De Vega and Alonso (2017); Breuer et al.
+(2016) dynamics has been studied following various methods and approaches Breuer et al.
 (2009); Rivas et al.
 (2010); Chruściński and Maniscalco (2014).
-Out of the various methods, one is the applicability of the collision model for NM dynamics Rau (1963); Alicki and Lendi (2007); Scarani et al.
+Under NM dynamics, where coherence endures and information retrogrades from the environment, a coherence-dependent correction to LB Reeb and Wolf (2014) becomes imperative.
+Unlike Markovian erasure, where dissipation follows a strict lower bound, NM effects can modify these limits, potentially reducing or amplifying heat dissipation.
+Here, we discuss LP under NM dynamics, exploring how memory effects influence erasure costs and coherence-driven corrections.
+This provides a more comprehensive view of quantum erasure beyond the standard Markovian framework.
+Out of the various methods for exploring NM dynamics, one of the prominent approaches is adoption of the collision model for OQS Rau (1963); Alicki and Lendi (2007); Scarani et al.
 (2002); Ziman and Bužek (2005); Gennaro et al.
 (2009).
-In the collision model, the NM dynamics for a system can be achieved when the interaction of the system and the environment is intervened by degrees of freedom of the ancilla.
-Under NM dynamics, where coherence endures and information retrogrades from the environment, a coherence-dependent correction to LB Reeb and Wolf (2014) becomes imperative.
-Unlike Markovian erasure, where dissipation follows a strict lower bound, NM effects can modify these limits, potentially reducing or amplifying heat dissipation.
-Here, we extend Landauer’s principle to NM dynamics, exploring how memory effects influence erasure costs and coherence-driven corrections.
-This provides a more comprehensive view of quantum erasure beyond the standard Markovian framework.
-The formulation of a Ladauer-like principle for heat flux in an erasure process delineated by the collision model of OQS is first proposed in Lorenzo et al.
+The process of analyzing the information to energy conversion in the collision model provides the platform for the foundation of the LP under NM dynamics.
+The formulation of a Ladauer-like principle for heat flux in an erasure process delineated by the collision model has been first proposed in Lorenzo et al.
 (2015).
-The process of analyzing the information to energy conversion in the collision model provides the platform for the foundation of the LP in NM dynamics.
 Figure 3: The system $\mathcal{S}$ interacts with the particle of the environment $\mathbf{E}$ which is termed as the subenvironment.
@@
 Here, $H_{E_{n}}$ denotes the free Hamiltonian of the n-th subenvironment and $H_{\mathcal{S}}$ denotes the free Hamiltonian of the system.
-The system will interact with the environment, and the interaction will be a sequence of pairwise collisions with the subenvironments.
+The system interact with the environment, and the interaction is governed by a sequence of pairwise collisions with the subenvironments.
 The environment is of a large size to curtail the situation that the system interacts with the same subenvironment twice.
-Each collision of this process is described by unitary evolution $U=e^{-igH_{int}\tau}$, where $g$ is the interaction strength, $\tau$ is the collision time, and $H_{int}$ denotes the interaction Hamiltonian of the system and the environment.
+Each collision of this process is described by unitary evolution $U=e^{-igH_{int}\tau}$, where $g$ is the interaction strength, $\tau$ is the interaction time, and $H_{int}$ denotes the time-independent interaction Hamiltonian of the system and the environment.
 The information that is stored in the system gets diluted while interacting with the environment.
@@
 […]
-$\beta\dot{Q}(t)\geq\dot{S^{\prime}}(\rho),$ (15)
-where $S^{\prime}(\rho)=-S(\rho)$.
+$\beta\dot{Q}(t)\geq\dot{\tilde{S}}(\rho),$ (15)
+where $\tilde{S}(\rho)=-S(\rho)$.
 Eq.
-(15) provides the formulation of the LP in NM dynamics of OQS.
-The formalism proposed in the work Lorenzo et al.
-(2015) has twofold power.
-One, it provides a time-resolved analysis of the erasure by thermalization, which enabled the formulation of the LP for the NM dynamics.
-Secondly, it was possible to elucidate the role of correlation in the information-erasure processes.
-This formalism provides the linkage between the erasure process with the intrasystem correlations that arise in the open quantum dynamics for multipartite systems.
-The violation of the LB due to strong correlation (proposed in Lorenzo et al.
-(2015)) is encountered in the spin-1/2 particle system Pezzutto et al.
-(2016) when used as the framework for the analysis.
-The sequence of discrete-time collisions of the system with one of the environmental particles at a time occurs in this process, and the Hamiltonian is ruled by the Heisenberg Hamiltonian.
+(15) provides the LP in OQS formulation involving heat and entropy fluxes.
+This relation provides a time-resolved analysis of the erasure by thermalization.
+Moreover, it is possible to elucidate the role of correlation among the subsystems, contributing through the entropy, in the information-erasure processes.
+This in turn enable us to probe the effect of the NM dynamics in information-erasure.
+The local violation of the LB due to the system-environment correlation (proposed in Lorenzo et al.
+(2015)) is demonstrated in the spin-1/2 particle system Pezzutto et al.
+(2016) when interaction between the system and environment is governed by the Heisenberg interaction.
+The sequence of discrete-time collisions of the system with one of the environmental particles at a time occurs in this process.
 In the long time limit, the environment (for non-interacting particles) undergoes homogenization dynamics, i.e., in the long time limit which is equivalent to a large number of collisions of the system with the environment, the state of the system approaches asymptotically to the initial preparation of the environment Ziman et al.
@@
 (2002).
-If the environment is considered to be composed of interacting particles, the system undergoes an NM process.
-Here, the elements of the environment are in a thermal state, and in the asymptotic limit homogenization dynamics are encountered.
-This behavior occurs when the state of the elements of the environment has weak fluctuation.
-Due to the inter-environment interactions, there is a memory effect in the dynamics of the system, which features the non-Markovianity.
-It was observed that there is an instantaneous violation of the LB for the system, the cause of which is the strong system-environment correlation.
+If the environment is considered to be composed of interacting particles, the system undergoes an NM dynamics.
+The elements of the environment are in a thermal state, and in the asymptotic limit, homogenization dynamics are encountered.
+This behavior occurs when the state of the elements of the environment has weak fluctuations.
+However, due to the inter-environment interactions, there is a memory effect in the dynamics of the system.
+It has been observed that during the time evolution, the system-environment correlation grows strong enough to exhibit NM features leading to an instantaneous violation of the LB.
 Figure 4: The system $\mathcal{S}$ interacts with the auxiliary system $\mathcal{S}_{1}$, which then interacts with the environment $\mathbf{E}$ and is redirected to its initial value after each collision.
@@
 -- $I(\rho_{\mathcal{S}\mathbf{E_{n}}})-S(\rho_{\mathbf{E_{n}}}||\rho_{\mathbf{E_{n}}}^{th})),$
-where $S(\rho||\rho^{\prime})=\text{Tr}\,\rho\,\ln\,\rho-\text{Tr}\,\rho\,\ln\,\rho^{\prime}$ describes the quantum relative entropy, $\rho_{\mathcal{S}_{n}}^{\prime}$ and $\rho_{\mathbf{E_{n}}}^{\prime}$ denotes the marginal states of the system and the environment respectively after the collision.
+where $S(\rho||\rho^{\prime})=\text{Tr}\,[\rho\,\ln\,\rho]-\text{Tr}\,[\rho\,\ln\,\rho^{\prime}]$ describes the quantum relative entropy, $\rho_{\mathcal{S}_{n}}^{\prime}$ and $\rho_{\mathbf{E_{n}}}^{\prime}$ denotes the marginal states of the system and the environment respectively after the collision.
 $I(\rho_{\mathcal{S}\mathbf{E_{n}}})$ characterizes the mutual information that quantifies the system-environment correlation based on the intra-collision strength.
 The second term in Eq.
-(16) (followed from Eq.
-(2)) describes the entropy flow from the environment to the system.
-The LB for the NM process is
+(16) describes the entropy flow from the environment to the system.
+The modified LB for the NM process can be expressed as
 $\beta_{\mathbf{E_{n}}}\Delta\tilde{Q_{n}}$ $=$ $\Delta\tilde{S}_{n}+S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\,\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-I(\rho_{\mathcal{S}\mathbf{E_{n}}})$ (17)
 -- $S(\rho_{\mathbf{E_{n}}}||\,\rho_{\mathbf{E_{n}}}^{th})),$
-where $\Delta\tilde{Q_{n}}=-\Delta Q_{n}$ and $\Delta\tilde{S}_{n}=-\Delta S_{n}$.
-When $I(\rho_{\mathcal{S}\mathbf{E_{n}}})\leq S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\rho_{\mathbf{E_{n}}}^{th}))$, the LP holds as long as the established system-environment correlations are smaller than the upper bound.
-So, the condition that allows the system to violate the LP is
-$I(\rho_{\mathcal{S}\mathbf{E_{n}}})>S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\,\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\,\rho_{\mathbf{E_{n}}}^{th})).$ (18)
-It can be generalized, where the system is coupled to a composite environment with $M$ different environments.
-The LP violation condition for the multiple NM environments is found to be equivalent to that of Eq.
+where $\Delta\tilde{Q_{n}}=-\Delta Q_{n}$ is the heat dissipated to the environment and $\Delta\tilde{S}_{n}=-\Delta S_{n}$ entropy decrease of the system.
+When $I(\rho_{\mathcal{S}\mathbf{E_{n}}})\leq S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\rho_{\mathbf{E_{n}}}^{th}))$, the LP holds, or we can say that the established system-environment correlations are smaller than the upper bound.
+Therefore, the condition that allows to violate the LP is
+$I(\rho_{\mathcal{S}\mathbf{E_{n}}})>S(\rho_{\mathcal{S}\mathbf{E_{n}}}^{\prime}||\,\rho_{\mathcal{S}_{n}}^{\prime}\rho_{\mathbf{E_{n}}}^{th})-S(\rho_{\mathbf{E_{n}}}||\,\rho_{\mathbf{E_{n}}}^{th}).$ (18)
+The authors have demonstrated violation of LP by taking the system-environment interaction to be the Heisenberg interaction.
+Their findings further suggest that the complexity of the environment has little to do with the violation of LP under NM dynamics.
+They have further shown that their results can be generalized to where the system is coupled to multiple environments.
+The LP violation condition for the multiple NM environments is found to be a straightforward generalization of Eq.
 (18).
-A solution to the violation of the LP has been put forward in the work Zhang et al.
+A solution to the apparent violation of the LP in the NM domain has been put forward in the work Zhang et al.
 (2021).
@@
 This composite environment induces NM dynamics (Fig.
-4).
-The memory in this model is $\mathcal{S}_{1}$.
+4) through the memory effect induced by $\mathcal{S}_{1}$.
 In the Markovian limit, the conventional LB holds.
@@
 where $\dot{Q}_{\mathcal{S}}(t)$ describes the heat flux from the system to the environment, $\dot{\tilde{S}}(\rho_{\mathcal{S}}(t))$ is the entropy flux from $\mathcal{S}$ to $\mathbf{E}$ ($\tilde{S}(\rho_{\mathcal{S}}(t))=-S(\rho_{\mathcal{S}}(t))$), $\dot{I}(\rho_{\mathcal{S^{\prime}}}(t))$ describes the rate of mutual information that characterizes the correlation between the system and auxiliary system and $\dot{S}(\rho_{\mathcal{S}_{1}}(t)||\rho_{\mathcal{S}_{1}}^{th})$ describes the rate of quantum relative entropy.
+Their model allows them to study the Markovian limit, and it has been shown that in the Markovian limit the above equation reduces to (15).
 The generalization of the model for multiple environments has also been reported in this work.
-The modified LP for the multiple environment case is also found to be valid in both regimes and is a direct consequence of Eq.
+The modified LP for the multiple environment case is also found to be valid in both regimes.
+Importantly, the quantum information theoretic framework that has been adopted to derive the the modified LBs in the NM domain discussed here broadly belongs to the same framework used to derive Eq.
 (2).
-Another aspect that was encountered for violation of the LP in the NM domain is the information back-flow.
-In the work Hu et al.
-(2022), the authors evaluated the connection between the information back-flow in the NM dynamics and LP.
+In Hu et al.
+(2022), the authors have evaluated the connection between the information back-flow in the NM dynamics and the violation of LP.
 For the analysis, a qubit is considered to be coupled with the environment.
-If the system is initially in the thermal state, it is inferred from the analysis that there is a one-to-one correspondence between the violation of the LE and the information backflow.
-Whereas the correspondence does not hold if the initial state of the system has coherence.
+If the system is initially in the thermal state, it is inferred from the analysis that there is a one-to-one correspondence between the violation of the LP and the information backflow.
+Whereas, the correspondence does not hold if the initial state of the system has coherence.
 IV Landauer limit in Computing
-A quantum computer Nielsen and Chuang (2002); Aharonov (1999); DiVincenzo (2000) harnesses quantum mechanical phenomena such as superposition and entanglement Mermin (1990); Linden et al.
+A quantum computer Nielsen and Chuang (2002); Aharonov (1999); DiVincenzo (2000) harnesses quantum mechanical resources such as superposition and entanglement Mermin (1990); Linden et al.
 (2006); Wineland (2013) to perform computations.
-Qubits, serve as the fundamental units of information in quantum computers and leverage the principles of quantum superposition enabling quantum computers to execute certain types of calculations much more efficiently than classical computers.
+Qubits serve as the fundamental units of information in quantum computers and leverage the principles of quantum superposition enabling quantum computers to execute certain types of calculations much more efficiently than classical computers.
 However, regardless of the computational task, it ultimately needs to be implemented on a physical system, implying that computations are subject to the constraints of the laws of physics.
-For instance, the LP places a minimum limit on the heat generated during bit erasure, and the quantum speed limit Golub and Ortega (2014); Caneva et al.
+For instance, the LP places a minimum limit on the heat generated during a bit erasure, and the quantum speed limit Golub and Ortega (2014); Caneva et al.
 (2009); Okuyama and Ohzeki (2018); Jones and Kok (2010); Deffner and Campbell (2017); del Campo et al.
 (2013); Deffner and Lutz (2013) dictates a limitation on how quickly a fundamental logical operation can be carried out.
-The modern computer uses a binary logic system, and LB for such binary computers is known to be $k_{B}T\ln 2$.
-IV.1 Landauer’s Limit For N-based logical computer
-A computer using the binary system can be exemplified as a single particle Szilard engine Szilard (1929) where the certainty of the particle’s presence in a chamber corresponds to the recording of 1 bit of information, and the uncertainty in the particle’s location corresponds to the erasure of 1 bit of information.
+IV.1 Landauer Limit For N-based logical computer
+A computer using the binary logic can be exemplified as a single particle Szilard engine Szilard (1929) where the certainty of the particle’s presence in a chamber among the two corresponds to the recording of 1 bit of information, and the uncertainty in the particle’s location corresponds to the erasure of 1 bit of information.
 However, computers are not limited to binary logic systems.
@@
 (1973); Knuth (1998).
-Recently, computers employing many-valued logic have garnered significance due to fundamental aspects and numerous applications Gottwald and Gottwald (2001); Chang (1958).
+Recently, computers employing many-valued logic have garnered significance due to both fundamental aspects and numerous applications Gottwald and Gottwald (2001); Chang (1958).
 The exemplification of the LP for binary logic computers can be supplied by the Brownian particle in a double potential well as shown in Fig.
 5.
-For a symmetrical well and a random bit $a_{0}=a_{1}=\frac{1}{2}$ (where $a_{0}$, $a_{1}$ denotes the probability) the LB is recovered by implementing Eq.
-(2).
-Now, whether LL holds for N-based logic is explored in Bormashenko (2019b).
-In the work, the authors have considered ternary logic, i.e., trit computing element for the analysis which is further generalized for N-based logic.
+For a symmetrical well and a completely random bit with $a_{0}=a_{1}=\frac{1}{2}$, where $a_{0}$, $a_{1}$ denotes the probability of being trapped in the wells, the LB is reached.
+Now, whether LB holds for N-based logic is explored in Bormashenko (2019b), where the authors have considered ternary logic, i.e., trit computing element for the analysis which is further generalized for N-based logic.
 Similar to binary logic computers, the LP is illustrated for ternary logic computers using a Brownian particle in a symmetric triple-well potential.
-For a random bit, $a_{0}=a_{1}=a_{2}=\frac{1}{3}$ (where $a_{0}$, $a_{1}$, $a_{2}$ denotes the probability), the LB is reached for the ternary logic by implementing Eq.
-(2), and this result has been further extended to N-based logic.
+For a random bit, $a_{0}=a_{1}=a_{2}=\frac{1}{3}$, where $a_{0}$, $a_{1}$, $a_{2}$ denotes the probability of being trapped in the wells, the LB is reached for the ternary logic, and this result has been further extended to N-based logic.
 Figure 5: Schematic of the qubit model with a Brownian particle $M$ placed in a double potential well.
@@
 (2010); Melhem (1992); Golub and Ortega (2014), involves the simultaneous utilization of multiple processors to solve algorithms.
-The primary objective is to distribute the workload among several processors to solve problems more swiftly or handle larger problems within the same timeframe.
-In the study Konopik et al.
-(2021), the authors examined the energy cost of finite-time irreversible computing using non-equilibrium thermodynamics principles.
-It has shown that the energy cost for computational tasks in a parallel computer, within a given finite time, closely adheres to the LL.
+The primary objective is to distribute the workload among several processors to solve problems more swiftly or handle larger problems within the same time window.
+In Konopik et al.
+(2021), the authors examined the energy cost of finite-time computing using non-equilibrium thermodynamics principles.
+It has been shown that the energy cost for computational tasks in a parallel computer, within a given finite time, closely adheres to the LB.
 This cost remains bounded even as the computational problem size increases.
-IV.2 Landauer’s bound in the presence of time-symmetric protocol
+IV.2 Landauer Limit in Presence of Time-symmetric Protocol
 Investigations into energy dissipation with time-symmetric protocols, demonstrates a fundamental trade-off between computational accuracy and energy cost Riechers et al.
 (2020).
-It shows that reducing logical error requires increasing energy dissipation, particularly under nonreciprocal operations, and thus time-symmetric protocols prevent computation from reaching the LB.
+It shows that reducing logical error requires increasing energy dissipation, particularly under nonreciprocal operations, and therefore time-symmetric protocols prevent computation from reaching the LB.
 In a subsequent work Wimsatt et al.
-(2021), authors utilized thermodynamic analysis of time-symmetric procedures (as prescribed in Riechers et al.
+(2021), authors did thermodynamic analysis of time-symmetric procedures (as prescribed in Riechers et al.
 (2020)) to thoroughly examine the trade-offs between accuracy and dissipation during information erasure.
 The authors employed nonequilibrium information thermodynamics to calculate the minimum energy dissipation needed for reliable erasure under time-symmetric control protocols.
-The energy costs associated with reliable erasure were found to be higher than those implied by the LB on information erasure.
-Moreover, these costs diverge in the limit of perfect computation.
-Therefore, the creation of time-asymmetric protocols is deemed necessary for effective and precise thermodynamic computation.
+The energy costs associated with this reliable erasure were found to be higher than those implied by the LB.
+Moreover, these costs diverge in the limit of perfect erasing.
+Therefore, the creation of time-asymmetric protocols is deemed necessary for effective and precise thermodynamic analysis of computation.
 Consequently, it is concluded that time asymmetry serves as a crucial design principle for thermodynamically efficient computing, warranting further investigation.
@@
 With our ever-increasing control over the miniaturized scale and the advancement of quantum technology, the validation of the LB is becoming more and more plausible day-by-day.
-Here in this section, we will focus on the recent experimental advancement to test LP in different technological platforms both in the stochastic and quantum realm.
+Here in this section, we will focus on the recent experimental advancement to test LP in different technological platforms.
 V.1 Optics Based Technology
 Optical Tweezers:
-The test of LP was possible (that remained untested over five decades) after the two basic advancements.
-One of them was to develop a method to investigate the work done on the particle as well as the heat dissipation by the particle based on the information on the trajectory of the particle and its potential.
-It was proposed and tested in the seminal paper by Sekimoto Sekimoto (1997, 2010).
+The test of LP has become possible (that remained untested over five decades) after the two basic advancements.
+One of them is to develop a method to investigate the work done on the particle as well as the heat dissipation by the particle based on the information on the trajectory of the particle and its potential.
+It has been proposed and tested in the seminal paper by Sekimoto Sekimoto (1997, 2010).
 The second advancement is the development of methods to impose user-defined potential on small particles.
@@
 $\bullet$ Now the intensity of the laser is reduced so that the barrier is low enough and the bead can jump from one to another.
-$\bullet$ Finally, after the bead ends in the required well independent of the initial state (this causes the memory erasure of the process when it is set to 1), the barrier is raised to its previous stage.
+$\bullet$ Finally, after the bead ends in the required well independent of the initial state (this causes the memory erasure when it is set to 1), the barrier is raised to its previous stage.
 Figure 6: The erasure protocol that is considered in the experiment.
@@
 (2021); Dago and Bellon (2022), differential interferometer Paolino et al.
-(2013) have been considered as the platform for the analysis.
+(2013) has been considered as the platform for the analysis.
 The working system is a micromechanical oscillator (the role played by a conductive cantilever in the experiment) confined in a double potential well.
@@
 Within the framework of stochastic thermodynamics, it has been shown in the work Dago et al.
-(2021) that one can reach the LB with high precision within a time scale of 100 ms which was previously recorded to be 30 s.
+(2021) that one can reach the LB with high precision within a time scale of 100 ms, which was previously recorded to be 30 s.
 The work was further extended by the authors to explore the overdamped and underdamped case within the stochastic thermodynamic framework with fast operation Dago and Bellon (2022).
@@
 (2014).
-The two internal levels of the system ion are considered as the qubit system and the vibrational modes of the ion as the finite temperature bath.
+The two internal levels of the system ion are considered as the qubit system, and the vibrational modes of the ion as the finite temperature bath.
 The LP is analyzed in the system by observing the phonon number in the variation of the ion.
@@
 V.4 Superconducting Technology
-An experiment has been performed on a hardware platform namely, superconducting flux logic for analyzing the quantum LP Saira et al.
+An experiment has been performed on a hardware platform, namely, superconducting flux logic, for analyzing the quantum LP Saira et al.
 (2020).
 The double well potential in this process arises due to the Josephson effect and the flux quantization.
-The bit erasure process is explored in the parametric regime where the approximation of metastable state is valid.
+The bit erasure process is explored in the parametric regime where the approximation of the metastable state is valid.
 It is observed that the process is bounded by LB.
@@
 We first discuss the Ballistic computer (BLC) proposed by Fredkin and Toffoli Fredkin and Toffoli (1982) and its limitations.
-Subsequently, we discuss the Brownian computer (BWC) which utilizes thermal fluctuation to perform a computational process.
-Before delving into reversible computation and its theromdynamic interpretation it is worthwhile to briefly overview the two key concepts: thermodynamic reversibility and logical reversibility.
+Subsequently, we discuss the Brownian computer (BWC), which utilizes thermal fluctuation to perform a computational process.
+Before delving into reversible computation and its thermodynamic interpretation, it is worthwhile to briefly overview the two key concepts: thermodynamic reversibility and logical reversibility.
 Thermodynamic reversibility Jarzynski (1997c); Crooks (1998, 1999); Jarzynski (2000); Seifert (2005); Kawai et al.
 (2007); Crooks (2011); Wolpert et al.
-(2024) can be defined as follows: A physical process is considered thermodynamically reversible if and only if the time evolution of the probability distribution in the process can be reversed.
+(2024); Sagawa (2018) can be defined as follows: A physical process is considered thermodynamically reversible if and only if the time evolution of the probability distribution in the process can be reversed.
 This reversal should include the time reversal of changes in external parameters, along with the inversion of the signs of both work and heat quantities.
@@
 7 (starting line here), a huge number of hard spheres (balls) are fired with equal velocity.
-If ‘1’ is there in the input, a ball is considered in the starting line else no ball for ‘0’.
+If ‘1’ is there in the input, a ball is considered in the starting line, else no ball for ‘0’.
 Due to the collision process inside, the ball changes its direction and collides with the other balls.
@@
 Though it conveys to decrease in the amount of cost in energy, we encounter some drawbacks of this setup.
-Two main drawbacks of this system are its sensitivity to small perturbative changes, and the second one is related to the collision of the balls.
+The two main drawbacks of this system are its sensitivity to small perturbative changes, and the second one is related to the collision of the balls.
 It is quite difficult to make each collision elastic.
@@
 The wave-packet spreading causes instability in the system in the quantum realm.
-Benioff Benioff (1982) in his work has discussed a quantum version of the BLC, where he has proposed a way to culminate the effect of the noise due to the wave packet spreading by utilizing a time-independent Hamiltonian.
+Benioff Benioff (1982) in his work has discussed a quantum version of the BLC, where he has proposed a way to mitigate the effect of the noise due to the wave packet spreading by utilizing a time-independent Hamiltonian.
 VI.2 Brownian Computer
@@
 If the total entropy ($S_{system}+S_{environment}$) remains constant throughout the process, it signifies a thermodynamically reversible process.
-Bennett and Landauer’s oversight of not tracking the BWC’s total entropy resulted in the misidentification that BWC model is a thermodynamically reversible process.
+Bennett and Landauer’s oversight of not tracking the BWC’s total entropy resulted in the misidentification that the BWC model is a thermodynamically reversible process.
 It has been indicated in recent studies Utsumi et al.
@@
 Imagine a toll gate controlled by a computer.
-Assume the gate remains closed until the required amount, say $25$ bucks, is paid.
-Moreover, assume that there are three sets of coins only: $5$, $10$, and $25$ bucks.
+Assume the gate remains closed until the required amount, say 252525 bucks, is paid.
+Moreover, assume that there are three sets of coins only: 555, 101010, and 252525 bucks.
 Now, let us consider a situation where the driver of the vehicle inserts 25 bucks in the sequence (5, 5, 10, 5).
@@
 The collection of the words forms a language of the FA, which happens to be a regular one (any finite language is called a regular language) for FA.
-DFA means the current state and the current symbol uniquely determine the next state.
-NFA means the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution.
-PFA means their is a conditional probability distribution over all possible next states given the current state and the current symbol.
+DFA means that the current state and the current symbol uniquely determine the next state.
+NFA means that the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution.
+PFA means there is a conditional probability distribution over all possible next states given the current state and the current symbol.
 FA have different forms like deterministic (DFA: the current state and the current symbol uniquely determine the next state), non-deterministic (NFA: the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution), and probabilistic finite automaton (PFA: there is a conditional probability distribution over all possible next states given the current state and the current symbol).
-In the stochastic (or probablistic) automaton, the single-valued transition function $\delta$ will be replaced by a conditional probability distribution.
+In the stochastic (or probabilistic) automaton, the single-valued transition function $\delta$ will be replaced by a conditional probability distribution.
 One can also observe multiple accept states described in the literature as ‘terminal states’ for the system Lawson (2003).
@@
 Universal computation machine (aka TM, see details below) requires infinite tape, i.e., infinite memory at its disposal, which often makes the model unsuitable for practical applications.
-On the other hand, finite state machines (FSM), aka FA, albeit non-universal, that use finite resources, are an alternative model as real-world computers have limited resources.
+On the other hand, finite state machines (FSM), aka FA, albeit non-universal, that use finite resources, are an alternative model, as real-world computers have limited resources.
 The construction of a thermodynamically efficient model Chu and Spinney (2018); Ouldridge and Wolpert (2022); Kardeş and Wolpert (2022) of FSMs has been in the limelight in recent times to estimate the energetic cost of such models.
@@
 One is the invertible local map and the other is the non-invertible local map.
-In the former case, zero minimal cost is encountered, whereas in the latter case, high cost is encountered.
+In the former case, zero minimal cost is encountered, whereas in the latter case, a high cost is encountered.
 An alternative approach other than the Markov chain model is addressed in Kardeş and Wolpert (2022), where they have developed a thermodynamic framework to define logical computers like DFA without specifying any extraneous parameters (like rate matrices, Hamiltonians, etc.) of the process that is considered to implement the computer.
@@
 For a given TM, the arguments of the transition states are called instantaneous descriptions (IDs) of the TM.
-One can also encounter TM’s with no halts.
+One can also encounter TMs with no halts.
 Earlier models like FA and push-down automata  [^3] The main difference of push-down automata with FA is that it has access to the top of the stack to decide which transition to take.
@@
 (1998).
-Now, advanced concepts like fault-tolerant quantum computation Shor (1996); DiVincenzo and Shor (1996); Gottesman (1998a) provides the route-map to the threshold theorem for error correction in the quantum regime Knill et al.
+Now, advanced concepts like fault-tolerant quantum computation Shor (1996); DiVincenzo and Shor (1996); Gottesman (1998a) provide the route-map to the threshold theorem for error correction in the quantum regime Knill et al.
 (1996); Aharonov and Ben-Or (1997).
@@
 Here we focus on a subfield of block code: Linear code.
-There are other models of error correction codes that are not covered here, interested readers can go through Pless (1978); Hoffman et al.
+There are other models of error correction codes that are not covered here; interested readers can go through Pless (1978); Hoffman et al.
 (1991) for further information.
@@
 One-to-one mapping from classical to quantum error correction (QEC) is not possible as the quantum world has some constraints of its own, like qubits are governed by the no-cloning principle Nielsen and Chuang (2002).
-As a consequence even the simple repetition code [^3] In the repetition code, the encoding is done as, for example, $00\rightarrow 0$ 000 and $11\rightarrow 1$ 111.
-It will correct a corrupted state, say $010$ to its majority value, $000$ in this case., which belongs to the linear block code class, does not work in the quantum domain.
+As a consequence, even the simple repetition code [^4] In the repetition code, the encoding is done as, for example, $\rightarrow 00$ 000 and $\rightarrow 11$ 111.
+It will correct a corrupted state, say 010010010, to its majority value, 000000000 in this case.
+This belongs to the linear block code class and does not work in the quantum domain.
 Additionally, the phenomenon of wavefunction collapse upon measurement makes quantum unique from classical.
@@
 (1998); Gottesman (1998a) have demonstrated that one can suppress the error rate in the quantum regime provided the qubits meet some physical conditions.
-A general qubit can be represented
+A general state of a qubit is represented as
 $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (24)
 where $\alpha$ and $\beta$ represents complex number satisfying $|\alpha|^{2}+|\beta|^{2}=1$.
-Thus, a qubit has the power to encode information in the infinite number of possible superposition of the computational basis states, which are denoted by $|0\rangle$ and $|1\rangle$.
-Therefore, the qubits are subjected to an infinite number of errors.
+Thus, a qubit has the power to encode information in the infinite number of possible superposition of the computational basis states, which are denoted by |0|0$\rangle$| 0 ⟩ and |1|1$\rangle$| 1 ⟩.
 However, due to the digitization of the errors using the Pauli operator, the error counts are reduced to two fundamental errors Nielsen and Chuang (2002).
@@
 An alternative formalism is explored in Korepin and Terilla (2002) to establish the conditions of quantum codes and investigate QEC conditions from a thermodynamic standpoint.
-In Vedral (2000), the ``$0"$ and ``$1"$ states are represented by whether a single molecule of an ideal gas is in the LHS or RHS of a box, respectively, like in Bennet’s MD setup.
-Say, initially the molecule is in the LHS or 00 state.
+In Vedral (2000), the 0``0"‘ ‘ 0 ” and 1``1"‘ ‘ 1 ” states are represented by whether a single molecule of an ideal gas is in the LHS or RHS of a box, respectively, like Bennet’s MD setup.
+Say, initially the molecule is in the LHS or 0 state.
 If it expands isothermally, it works $\Delta W=k_{B}T\ln 2$ at the expense of the same amount of free energy stored in it.
@@
 Effectively we correct $A_{\alpha}$ by transferring the error to $B_{\alpha}$.
-(5) In the last step, the system $\mathcal{B}_{\alpha}$ is brought back to its initial state by isothermal compression, and it requires work to be done.
+(5) In the last step, the system $\mathcal{B}_{\alpha}$ is brought back to its initial state by isothermal compression, and it requires work.
 In summary, to correct $A_{\alpha}$ or reduce its entropy $B_{\alpha}$’s free energy is wasted just like how a refrigerator works.
-Now to restore $B_{\alpha}$’s free energy its entropy ($\Delta S=k_{B}\,\ln 2$) has to be dumped into the environment and it needs at least $\Delta W=k_{B}T\ln 2$ amount of work to be done which makes it exactly similar to the LP, thereby consistent with the second law of thermodynamics.
+Now to restore $B_{\alpha}$’s free energy its entropy ($\Delta S=k_{B}\,\ln 2$) has to be dumped into the environment and it needs at least $\Delta W=k_{B}T\ln 2$ amount of work, which makes it exactly similar to the LP, and thereby consistent with the second law of thermodynamics.
 Figure 13: A schematic of the reversible cycle of the classical error correction protocol.
@@
 (2011); Kosloff (2013); Roßnagel et al.
-(2016); Martínez et al.
+(2016); Mart´ınez et al.
 (2016); Chattopadhyay and Paul (2019); Uzdin et al.
@@
 (2021); Singh et al.
-(2020, 2023).
+(2020, 2023); Gelbwaser-Klimovsky et al.
+(2013, 2015); Misra et al.
+(2022).
 In a simple sense, one can convey that the Carnot engine extracts an amount of heat $Q_{hot}$ from the hot reservoir at temperature $T_{hot}$ and transfers an amount of heat $Q_{cold}$ to the sink at temperature $T_{cold}$.
@@
 X Conclusion and Future Direction
-The analysis of the cost (like thermodynamic cost, operational cost) while executing a process has been one of the central points of attraction for physicists as well as computer scientists Aifer et al.
+Both the thermodynamic and computational cost of a process have been of prime interests for physicists as well as computer scientists Aifer et al.
 (2023); Coles (2023).
-In physics, it is the thermodynamic cost of the process that plays a critical role, whereas in computer science, it is the computational cost.
-The analysis of the thermodynamic computational cost for a process was initiated from the physical Church-Turing thesis Kolchinsky (2023), where they conveyed that every computational process is physical.
-Various approaches have been considered for the analysis of the different computational processes.
-With the advent of modern statistical theory, the research in this area got boosted, yet it still remains a challenging task to explore thermodynamic computational cost (aka Landauer principle) for OQS, many-body physics, and at the quantum phase transition (QPT) Sachdev (1999); Vojta (2003); Osborne and Nielsen (2002); Heyl (2018); Sen et al.
-(2005); Prabhu et al.
-(2011) point.
-QPT takes place in a quantum many-body system Bandyopadhyay et al.
-(2021); Sur and Ghosh (2020); Fetter and Walecka (2012); Tasaki (2020); De Chiara and Sanpera (2018); Mukherjee et al.
-(2007); Bose (2003); Dutta (2015); Gómez et al.
-(1996); Ganahl et al.
-(2012); Fukuhara et al.
-(2013); Subrahmanyam (2004); Iyoda and Sagawa (2018); Tian et al.
-(2013) which happens at absolute zero temperature as a result of quantum fluctuations.
-The analysis of the minimum cost of heat dissipation of the erasure process in QPT is an open domain to explore.
+While the physicists are more concerned with the thermodynamic cost a, the computer scientists mostly care about computational cost.
+The analysis of the thermodynamic computational cost for a process surfaced from the physical Church-Turing thesis Kolchinsky (2023), where it was conveyed that every computational process is physical.
+Various approaches have been considered for the analysis of thermodynamic cost of the different computational processes.
+In this article we have reviewed the recent findings in these line in both the quantum and classical regime as well as provided an overview on the foundation and frameworks of thermodynamics of computation.
 In recent times, the applications of these principles are not just restricted to the field of physics or computer science but are also prevalent in other fields like chemical networks Chen et al.
@@
 (2001).
-To get a better understanding of the bond between thermodynamics and the computational process, further investigation in this direction is still required.
-For example, in the case of a finite automaton, one can investigate the maximum thermodynamic cost that is required to accept a language for automata.
+To get a better understanding of the bond between thermodynamics and the computational process, further investigations are still required.
+For example, in the case of a finite automaton, it is tantalizing to investigate the upper bound thermodynamic cost that is required to accept a language for automata.
 Also, one can calculate the minimal cost for any deterministic finite automaton.
-One can also work on developing a theory to analyze the non-deterministic finite automata in terms of thermodynamics.
-Models to describe the complex Turing machine, and also network theory from the thermodynamic viewpoint, are an open area of research.
+Developing a theory to analyze the non-deterministic finite automata with thermodynamic tools could be promising.
+Models to describe the complex Turing machine, and also network theory from the thermodynamic viewpoint, are open areas of research.
+Correction of errors during a computational process or communication is crucial.
+The analysis of the error correction protocols from a thermodynamic viewpoint is at its baby stage.
+Modeling of the EC models by the physical system to explain it thermodynamically needs further investigation.
+So the thermodynamic approach to explaining EC is an open book to read.
+While modern statistical and quantum information theory have advanced our understanding of the thermodynamic cost of computation, applying these insights to many-body physics, especially near quantum phase transitions (QPTs) Sachdev (1999); Vojta (2003); Osborne and Nielsen (2002); Heyl (2018); Sen et al.
+(2005); Prabhu et al.
+(2011) remains a significant challenge.
+QPTs in a quantum many-body system Bandyopadhyay et al.
+(2021); Sur and Ghosh (2020); Fetter and Walecka (2012); Tasaki (2020); De Chiara and Sanpera (2018); Mukherjee et al.
+(2007); Bose (2003); Dutta (2015); Gómez et al.
+(1996); Ganahl et al.
+(2012); Fukuhara et al.
+(2013); Subrahmanyam (2004); Iyoda and Sagawa (2018); Tian et al.
+(2013) occur at zero temperature due to quantum, rather than thermal fluctuations and are characterized by critical phenomena such as diverging correlation lengths and long-range entanglement.
+These features complicate the application of standard thermodynamic frameworks and raise open questions about how the minimal heat dissipation, governed by LP, behaves near criticality.
+Some compelling future direction involves relaxing the assumption of thermal reservoirs.
+These athermal characteristics could significantly modify the minimal energetic cost of erasure in general as well as near the QPT.
+Another fundamental challenge lies in understanding the behavior of LB as one approaches absolute zero, where the third law of thermodynamics and quantum effects become increasingly relevant.
+Additionally, pre-existing correlations between the system and a finite-sized reservoir may influence the erasure process in ways not captured by conventional formulations.
+Though there have been few explorations along these directions as we have reviewed, further explorations are essential for a deeper understanding of information thermodynamics in realistic, quantum-coherent environments.
 It is known that different systems have different heat signatures.
@@
 Thermodynamic analysis of quantum computations needs a rigorous investigation for a better understanding of quantum computers and to develop hardware with lower-cost functions.
-Correction of errors during a computational process or communication is crucial.
-The analysis of the error correction protocols from a thermodynamic viewpoint is at its baby stage.
-Modeling of the EC models by the physical system to explain it thermodynamically needs further investigation.
-So the thermodynamic approach to explaining EC is an open book to read.
 XI Acknowledgement
-P.C., A.M. would like to thank Nilakanta Meher and Saikat Sur of the Weizmann Institute of Science for their valuable inputs and suggestions.
+P.C., A.M.
+would like to thank Nilakanta Meher and Saikat Sur of the Weizmann Institute of Science for their valuable inputs and suggestions.
+A.M.
+acknowledges partial support from the Anusandhan National Research Foundation, Government of India, under grant number ANRF/ECRG/2024/003836/PMS.
 Appendix A Theorem of thermodynamic computation